### v1

# More for Keys, Less for Values: Adaptive KV Cache Quantization

Abstract
This paper introduces an information-aware quantization framework that adaptively compresses the key-value (KV) cache in large language models (LLMs). Although prior work has underscored the distinct roles of key and value cache during inference, our systematic analysis—examining singular value distributions, spectral norms, and Frobenius norms—reveals, for the first time, that key matrices consistently exhibit higher norm values and are more sensitive to quantization than value matrices. Furthermore, our theoretical analysis shows that matrices with higher spectral norms amplify quantization errors more significantly. Motivated by these insights, we propose a mixed-precision quantization strategy, KV-AdaQuant, which allocates more bit-width for keys and fewer for values since key matrices have higher norm values. With the same total KV bit budget, this approach effectively mitigates error propagation across transformer layers while achieving significant memory savings. Our extensive experiments on multiple LLMs (1B - 70B) demonstrate that our mixed-precision quantization scheme maintains high model accuracy even under aggressive compression. For instance, using 4-bit for Key and 2-bit for Value achieves an accuracy of 75.2%, whereas reversing the assignment—2-bit for Key and 4-bit for Value—yields only 54.7% accuracy. The code is available at https://tinyurl.com/kv-adaquant.
More for Keys, Less for Values: Adaptive KV Cache Quantization
Mohsen Hariri1, Lam Nguyen1, Sixu Chen1, Shaochen Zhong2, Qifan Wang3, Xia Hu2, Xiaotian Han1, Vipin Chaudhary1 1Case Western Reserve University; 2Rice University; 3Meta mohsen.hariri@case.edu
1 Introduction
Size K2V4 K4V2
1B 0.06 0.34
8B 0.55 0.75
14B 0.78 0.91
14B 0.76 0.87
Figure 1: Key cache needs more bits. (Left): Spectral norms of the key cache (blue) and value cache (orange) across layers in Llama3.3-70B show that key caches consistently exhibit higher norms. (Right): GSM8k accuracy for two schemes: $\text{K}_{2}\text{V}_{4}$, representing 2-bit allocation for the K cache and 4-bit allocation for the V cache and $\text{K}_{4}\text{V}_{2}$, representing 4-bit allocation for the K cache and 2-bit allocation for the V cache, demonstrates that allocating more bits to the key cache maintains strong performance, confirming the efficacy of norm-aware, mixed-precision quantization.
Large language models (LLMs) have advanced NLP by enabling parallel processing and capturing complex dependencies Vaswani et al. (2017); Sutskever et al. (2014), which have scaled from GPT-1’s 117M Radford et al. (2018), GPT-2’s 1.5B Radford et al. (2019) to GPT-3’s 175B Brown et al. (2020) and GPT-4’s 1.8T parameters OpenAI et al. (2024). Open-source models like Llama have grown to 405B Touvron et al. (2023), with Mistral Large (123B) Mistral-AI (2023) and DeepSeek V3 (671B) DeepSeek-AI (2023) also emerging. In parallel with increasing parameter counts, longer context windows are becoming essential for chain-of-thought reasoning, necessitating an expansion of the KV cache and contributing to memory bottlenecks Wei et al. (2022). New systems support context windows of up to 1M tokens DeepMind (2025), 200k tokens OpenAI (2024, 2025), and 130k tokens DeepSeek-AI et al. (2025). NVIDIA’s Blackwell architecture introduces low-precision formats like FP8, FP4, and INT4 NVIDIA (2025), emphasizing the need for precise quantization methods. The KV cache is a critical component in these models and its effective operation is paramount and significant.
To accommodate larger context windows, the KV cache can become a significant memory bottleneck. One effective solution is to quantize the key and value caches from high-precision formats (e.g., FP32 or BF16) to lower-precision ones (e.g., INT4 or INT2). However, we first discovered that the Key cache generally exhibits higher norms than the Value cache, making it more sensitive to uniform quantization. Since matrices with larger norms are more prone to quantization errors, bit-width allocation should be tailored to each matrix’s norm distribution. To confirm this, we conducted experiments with Llama3.3-70B (see Figures 1 (Left) and 2), as the Key cache (blue) consistently shows elevated norm values compared to the Value cache (orange). Furthermore, Figure 1 (Right) demonstrates that the K(4bit)V(2bit) scheme significantly outperforms the K(2bit)V(4bit) configuration on GSM8K dataset. These results underscore the need for a norm-aware, mixed-precision strategy that allocates higher precision to Keys and lower precision to Values.
Guided by these inspiring observations, we propose a mixed-precision quantization strategy, KV-AdaQuant, which allocates different bit-width to more bits for keys and fewer bits for values. since key matrices consistently have higher norm values. With the same total KV bit budget, this approach effectively mitigates error propagation across transformer layers while achieving significant memory savings. KV-AdaQuant uses singular values, spectral norms, and Frobenius norms to assign more bits to Keys and fewer to Values, balancing memory savings with preserving critical information. This work makes the following contributions: Guided by these insights, we propose a mixed-precision quantization strategy, KV-AdaQuant, which assigns a higher bit-width to key caches and a lower bit-width to value caches—reflecting the consistently higher norm values observed in key matrices. KV-AdaQuant leverages metrics such as singular values, spectral norms, and Frobenius norms to identify the critical information in KV cache. By maintaining the same overall KV bit budget, our approach effectively mitigates error propagation across transformer layers while offering significant memory savings. Our contributions include:
  • We systematically investigate the KV cache in transformers, uncovering structural patterns in how information is stored and retrieved.
  • We introduce KV-AdaQuant, an information-aware quantization method that leverages singular value distributions to allocate mixed precision—assigning higher bit-widths to keys and lower bit-widths to values.
  • We demonstrate the efficacy of our approach on the GSM8K benchmark, achieving substantial memory savings without compromising performance.
2 Key and Value are Different
This section investigates the differences between key and value caches to enhance quantization strategies.
2.1 Singular Value Distribution of the KV Cache
An extensive collection of models, ranging in size up to 70B parameters (Llama 3.3 70B Grattafiori et al. (2024)), is examined to explore the singular values of the KV cache. Random batches are sampled from the C4 Dodge et al. (2021), GSM8K Cobbe et al. (2021), and MMLU Hendrycks et al. (2021) datasets, and KV caches are extracted across all attention heads and layers. Figure 3 illustrates the distribution of singular values for the KV cache, starting from the 5th largest singular value onward, highlighting how keys and values diverge. Additionally, Figure 7 in Appendix C presents the full-range distribution, further revealing the discrepancy in larger singular values. Beyond individual singular values, the spectral norm ($\|M\|_{2}$) and the Frobenius norm ($\|M\|_{F}$) are computed, providing empirical evidence of the substantial gap in key and value matrix singular values and illustrating how this characteristic of the KV cache can inform the theoretical analysis of quantization error propagation in multi-layer attention networks. This discrepancy underscores the importance of careful precision allocation to mitigate performance degradation while saving more memory in the inference time.
2.2 Rank of KV Cache
The rank of a matrix corresponds to the number of nonzero singular values and determines the effective dimension of its column or row space. In multi-head self-attention, the key ($K$) and value ($V$) matrices possess one fixed dimension, $d_{\mathrm{head}}$, and one variable dimension corresponding to the sequence length, $\mathrm{seq\_len}$. Consequently, $K$ and $V$ can be analyzed in two distinct regimes: (a) when $\mathrm{seq\_len}<d_{\mathrm{head}}$, and (b) when $\mathrm{seq\_len}\geq d_{\mathrm{head}}$. In the first regime, the matrices are relatively short and thus offer limited opportunities for compression. Accordingly, many existing KV-cache compression methods either avoid compressing in this stage or employ specialized strategies that differ from those used later.
When $\mathrm{seq\_len}\geq d_{\mathrm{head}}$, the $K$ and $V$ matrices assume rectangular shapes with dimensions $\mathrm{seq\_len}\times d_{\mathrm{head}}$. Although the exact rank is data-dependent (since $K=XW_{K}$, where $X$ is drawn from a high-dimensional embedding space), empirical observations suggest that both key and value cache typically approach a rank close to $d_{\mathrm{head}}$. Therefore, in practical settings it is reasonable to assume that there are approximately $d_{\mathrm{head}}$ nonzero singular values for both $K$ and $V$. In the remainder of this section, we focus on this regime to analyze how spectral and Frobenius norm considerations guide our quantization strategy.
2.3 Stability of the Dequantized KV Cache: Error Propagation
The spectral norm of a matrix $M$, denoted $\|M\|_{2}$, quantifies the maximum stretching factor that $M$ can impart on any vector. In a multi-layer transformer (e.g., Llama 3.3 Grattafiori et al. (2024) with 80 layers or Phi-4 Abdin et al. (2024) with 16 layers), the sequential application of transformations implies that quantization errors introduced at one layer can propagate and amplify in subsequent layers.
To illustrate this effect, consider a simplified layer without normalization or additional scaling Elhage et al. (2021), where the hidden state at layer $(l+1)$ is given by
$h_{l+1}=h_{l}+W_{l}h_{l},$ (1)
Let $W_{l}$ denote the weight matrix at layer $l$. Define $\widetilde{W}_{l}$ as the result of quantizing the KV cache, and $W_{l}$ as the unquantized version. Then, the local error incurred at layer $l$ can be bounded as
$\begin{split}\|h_{l+1}-\widetilde{h}_{l+1}\|_{2}&=\|(W_{l}-\widetilde{W}_{l})h_{l}\|_{2}\\ &\leq\|W_{l}-\widetilde{W}_{l}\|_{2}\|h_{l}\|_{2}.\end{split}$ (2)
Propagating such errors through multiple layers yields a cumulative effect that is proportional to the product of the spectral norms at each layer. In other words, even small perturbations introduced by quantization can be significantly amplified if the weight matrices exhibit large singular values. Empirical observations indicate that key matrices consistently have higher spectral norms than value matrices. This suggests that errors in key matrices are more prone to propagate and magnify throughout the transformer stack. As a result, it is advisable to employ less aggressive quantization (e.g., 4-bit rather than 2-bit) for key matrices, while value matrices—characterized by lower spectral norms—can tolerate more aggressive quantization without incurring substantial distortion.
2.4 Total Energy of KV Cache (MSE Perspective)
Figure 2: Frobenius norm of key and value cache across layers for the Llama 3.3 70B model on the C4 dataset. The x-axis represents the layer index. The shaded region shows the min-max range across attention heads within each layer, and the solid line indicates the mean norm per layer. Key matrices exhibit significantly higher norms compared to value matrices, highlighting their stronger representational capacity.
The Frobenius norm of a matrix, denoted by $\|M\|_{F}$, is defined as the square root of the sum of the squared entries:
$\|M\|_{F}=\sqrt{\sum_{i,j}M_{i,j}^{2}}$
Equivalently, it can be viewed as the square root of the sum of the eigenvalues of $M^{\top}M$, meaning it captures the total “energy” of $M$. In the context of mean-squared error (MSE), the Frobenius norm provides a direct indicator of how much overall distortion might occur when quantizing the matrix.
Quantization Error and the Frobenius Norm.
When a matrix $M$ is quantized to a lower-precision representation $\widetilde{M}$, the quantization error in MSE terms is often measured by the squared Frobenius norm:
$\|M-\widetilde{M}\|_{F}^{2}=\sum_{i,j}(M_{i,j}-\widetilde{M}_{i,j})^{2}.$
A common assumption is that this error scales with both the range of the entries in $M$ and the granularity determined by the bit precision. Formally, one may write
$\mathbb{E}\bigl{[}\|M-\widetilde{M}\|_{F}^{2}\bigr{]}\;\propto\;\|M\|_{F}^{2}\times 2^{-2b},$
where $b$ denotes the number of bits used per entry. [^1] The proportionality depends on the quantizer design and the distribution of entries in $M$. For more details see Appendix A. For matrices with the same dimensions (e.g., key and value cache), the MSE and squared Frobenius norm are directly proportional by a factor of $1/mn$. Consequently, when analyzing quantization error for matrices of identical size:
  • Minimizing $\|M-\widetilde{M}\|_{F}^{2}$ is equivalent to minimizing $\text{MSE}(M,\widetilde{M})$, as the proportionality constant $1/mn$ does not affect optimization or relative comparisons.
  • The scaling relationship $\mathbb{E}\bigl{[}\|M-\widetilde{M}\|_{F}^{2}\bigr{]}\propto\|M\|_{F}^{2}\times 2^{-2b}$ holds equivalently for MSE, since both metrics share the same dependence on $b$ and $\|M\|_{F}^{2}$.
Thus, the Frobenius norm of the quantization and mean square of the quantization error can be used interchangeably for key and value cache, as they encode the same error structure up to a constant scaling factor.
Figure 3: Singular value distribution of key and value cache for the Llama 3.3 70B model on the C4 dataset. The x-axis represents the singular value indices, ordered from the 5th largest to the smallest, while the y-axis denotes the magnitude of the singular values. The shaded region indicates the minimum-to-maximum range across attention heads within each layer, and the solid lines represent the mean singular value magnitude at each index. This zoomed-in view reveals that, beyond the largest singular value (spectral norm), key matrices exhibit consistently higher singular values even in the lower ranges, underscoring their greater representational significance compared to value matrices. Consequently, key matrices should be quantized with less aggressive compression. See Figure 7 in Appendix C for the complete range.
Larger Frobenius Norms Require Higher Precision.
As we discussed in  2.1 empirical measurements across multiple transformer models reveal that key matrices have significantly larger Frobenius norms than value matrices. This disparity suggests that if key matrices are quantized too aggressively (e.g., with fewer bits), the MSE incurred could be disproportionately high. Formally, for two matrices $K$ (key) and $V$ (value), if
$\|K\|_{F}\;\gg\;\|V\|_{F},$
then allocating the same limited bit precision to both $K$ and $V$ would cause
$\|K-\widetilde{K}\|_{F}^{2}\;\gg\;\|V-\widetilde{V}\|_{F}^{2},$
since the error term for $K$ is amplified by its larger energy. Furthermore, errors introduced in earlier layers propagate through subsequent layers in a multi-layer transformer architecture, compounding the negative impact of coarse quantization.
3 Method
3.1 Adaptive KV Cache Quantization (KV-AdaQuant)
Building on the theoretical findings that matrices with higher spectral norms propagate quantization error more strongly and that larger Frobenius norms incur greater quantization error, along with empirical evidence indicating that key matrices exhibit higher norms than value matrices, a straightforward strategy emerges: quantize keys with higher precision and values with fewer bits. In particular, allocating 4-bit precision to keys and 2-bit precision to values effectively controls the overall mean squared error (Figure 4), reducing memory usage while mitigating error propagation through successive transformer layers.
(a) Mean4squared quantization error across layers for 2-bit quantization.
(b) Mean squared quantization error across layers for 4-bit quantization.
Figure 4: Quantization error (mean squared error) of key and value cache for the Llama 3.3 70B model on the C4 dataset across all transformer layers. The x-axis represents the layer index, and the y-axis indicates the mean squared quantization error. Key matrices exhibit consistently higher quantization error compared to value matrices, with larger discrepancies observed in lower-bit quantization.
3.2 Differences between KV-AdaQuant and Other Methods
The Hugging Face Transformers library implements a KV cache quantization technique that applies uniform precision to KV cache. Although straightforward, this approach can be suboptimal due to the distinct sensitivities of keys and values to quantization errors. Consequently, assigning the same bit-width to both key and value often introduces notable performance degradation. Furthermore, increasing the bit-width for value matrices offers minimal or no additional benefits, indicating that uniform quantization is not an ideal choice.
Other methods, such as SKVQ Duanmu et al. (2024) and QAQ Dong et al. (2024), also propose low-bit quantization for the KV cache, leveraging mixed-precision settings. SKVQ, for example, quantizes the key cache to 2 bits and the value cache to 1.5 bits, demonstrating minimal accuracy loss; however, it relies on fixed hyperparameter configurations that do not deeply analyze the key and value cache’ internal characteristics. QAQ similarly introduces quality-adaptive quantization to highlight the heightened sensitivity of key matrices to quantization errors, yet does not thoroughly investigate the underlying matrix properties, making generalization across diverse model architectures and tasks less certain.
By contrast, the proposed approach offers a more general framework of cache quantization that explicitly accounts for the distinct characteristics of key and value cache. This strategy reduces error propagation across transformer layers while achieving considerable memory savings. Additionally, it remains adaptable to further enhancements, such as calibration, grouping, and clustering, allowing the integration of complementary techniques to improve KV cache compression. This adaptability renders the framework orthogonal to existing methods, providing a more comprehensive solution to the challenges associated with KV cache quantization.
4 Experiments
A series of experiments is designed to investigate singular value patterns in KV cache and to demonstrate how these patterns can be leveraged to enhance practical quantization strategies. The experimental testbed consists of a diverse set of transformer models. Two distinct data setups are employed: one for analyzing the singular value distributions and another for evaluating the effectiveness of the proposed quantization approach.
4.1 Experiment Settings
Singular Value and Norm Analysis
This section examines structural differences between key and value cache by analyzing their singular values and norms. The process begins by feeding random batches (10 samples per batch) from the C4, MMLU, and GSM8K datasets into the model for each query, where every sample is padded to match the longest sequence to facilitate observations across diverse inputs. For each newly generated token (up to 1,000 tokens), the KV cache from each attention mechanism is collected, yielding matrices $\mathbf{K},\mathbf{V}\in\mathbb{R}^{d_{\text{head}}\times\text{seq\_len}}$ for each head and layer. Singular value decomposition is then performed offline in float32 precision using the PyTorch torch.linalg.svd operator, and the resulting singular values, along with the spectral norms ($\|\cdot\|_{2}$) and Frobenius norms ($\|\cdot\|_{F}$), are computed and averaged across all heads and layers to capture global patterns. This approach reveals how $\|\mathbf{K}\|_{2}$ and $\|\mathbf{V}\|_{2}$ evolve in deeper layers, sheds light on the total “energy” in these matrices, and clarifies the distribution of larger and smaller singular values, offering insights into their susceptibility to quantization errors.
Quantization Error
Hardware typically supports 2-bit, 4-bit, and 8-bit quantization, but in theory, any bit-width can balance compression and accuracy. We measure quantization error across bit-widths from 2 to 8 by quantizing and dequantizing key-value matrices, finding that MSE (and thus the Frobenius norm) behaves similarly across all bit-widths. As shown in Figure 5, the key matrix consistently exhibits higher error.
Bit-Width Configurations.
The effect of different bit-widths used to represent each entry in $\mathbf{K}$ and $\mathbf{V}$ is quantified. Uniform settings include bit-widths ranging from 2 to 8 bits for both keys and values (e.g., $(2,2)$, $(4,4)$, or $(8,8)$). Mixed-precision settings compare $(2,4)$ and $(4,2)$, assigning 2 bits to either values or keys and 4 bits to the other.
Figure 5: Mean squared error (MSE) of KV cache quantization for the Llama 3.3 70B model on the C4 dataset, using quantization bit-widths ranging from 2 to 8. The x-axis represents the quantization bit-width, while the y-axis shows the MSE on a logarithmic scale. A logarithmic scale is used to highlight differences at higher bit-widths, where MSE values decrease significantly and approach zero, particularly at 8-bit. Solid lines indicate the mean MSE across layers, while the shaded regions represent the min-max range of errors.
Quantization Procedure.
Hugging Face’s Optimum library is used for low-level CUDA kernels. No additional outlier or chunk-based grouping techniques are applied, so as to isolation of the effect of $\mathbf{K}$-vs-$\mathbf{V}$ bit allocation.
Error Metric: MSE.
To measure the quantization distortion, we record
$\text{MSE}(\mathbf{K},\widetilde{\mathbf{K}})\;=\;\frac{1}{d_{\text{head}}\,\times\,\text{seq\_len}}\,\|\mathbf{K}-\widetilde{\mathbf{K}}\|_{F}^{2},$
and likewise for $\mathbf{V}$. In addition to per layer and per head comparison, we average MSE across all heads, layers, and tokens in a batch. This approach yields a global view of how severely each bit-width compresses $\mathbf{K}$ vs. $\mathbf{V}$.
Experimental Protocol.
For each model and dataset subset (C4, MMLU, GSM8K):
  1. 1.
    Run a forward pass to produce unquantized KV caches in bfloat16.
  2. 2.
    Quantize to a chosen bit-width, then immediately dequantize to bfloat16.
  3. 3.
    Compare the dequantized version $\widetilde{\mathbf{K}}/\widetilde{\mathbf{V}}$ to the original $\mathbf{K}/\mathbf{V}$ in MSE terms.
  4. 4.
    Repeat for all bit-width settings of interest.
Table 1: Quantization error (mean ± std) for key and value cache at 2-bit, 3-bit, and 4-bit quantization ($K_{i}$ and $V_{i}$) across different datasets.
Dataset Model 2-bit 3-bit 4-bit
$K_{2}$ $V_{2}$ $K_{3}$ $V_{3}$ $K_{4}$ $V_{4}$
MMLU Llama3.2-1B 4.851 ± 1.037 0.127 ± 0.101 1.037 ± 0.265 0.021 ± 0.015 0.227 ± 0.059 0.005 ± 0.003
Llama3.1-8B-it 6.003 ± 1.782 0.187 ± 0.127 1.082 ± 0.244 0.028 ± 0.019 0.235 ± 0.055 0.006 ± 0.004
Llama3.3-70B-it 4.883 ± 1.106 0.112 ± 0.093 0.942 ± 0.198 0.016 ± 0.012 0.206 ± 0.043 0.003 ± 0.003
Phi4 5.929 ± 1.545 0.657 ± 0.472 1.306 ± 0.231 0.103 ± 0.070 0.286 ± 0.050 0.022 ± 0.015
Mistral0.3-7B 4.718 ± 1.340 0.398 ± 0.405 0.941 ± 0.240 0.059 ± 0.059 0.206 ± 0.053 0.013 ± 0.013
Qwen2.5-14B 5.184 ± 2.241 1.270 ± 1.547 1.005 ± 0.288 0.182 ± 0.221 0.223 ± 0.067 0.040 ± 0.052
DeepSeekR1Q-14B 5.126 ± 2.375 1.406 ± 1.609 0.900 ± 0.269 0.198 ± 0.226 0.199 ± 0.062 0.044 ± 0.052
C4 Llama3.2-1B 4.885 ± 1.056 0.207 ± 0.166 1.074 ± 0.289 0.030 ± 0.024 0.233 ± 0.062 0.006 ± 0.005
Llama3.1-8B-it 6.262 ± 1.789 0.254 ± 0.185 1.128 ± 0.249 0.036 ± 0.026 0.247 ± 0.056 0.008 ± 0.005
Llama3.3-70B-it 4.391 ± 1.027 0.121 ± 0.097 0.847 ± 0.175 0.017 ± 0.013 0.186 ± 0.038 0.004 ± 0.003
Phi4 5.715 ± 1.442 0.850 ± 0.684 1.316 ± 0.245 0.124 ± 0.093 0.291 ± 0.056 0.027 ± 0.020
Mistral0.3-7B 5.027 ± 1.332 0.543 ± 0.493 1.014 ± 0.269 0.079 ± 0.068 0.223 ± 0.060 0.017 ± 0.015
Qwen2.5-14B 4.382 ± 2.170 1.544 ± 1.872 0.846 ± 0.250 0.220 ± 0.265 0.187 ± 0.060 0.048 ± 0.060
DeepSeekR1Q-14B 4.832 ± 2.354 1.651 ± 1.914 0.927 ± 0.283 0.232 ± 0.267 0.201 ± 0.061 0.051 ± 0.060
GSM8K Llama3.2-1B 5.703 ± 1.557 0.179 ± 0.136 1.213 ± 0.352 0.026 ± 0.020 0.266 ± 0.078 0.005 ± 0.004
Llama3.1-8B-it 6.445 ± 1.837 0.213 ± 0.161 1.184 ± 0.268 0.030 ± 0.022 0.257 ± 0.060 0.007 ± 0.005
Llama3.3-70B-it 4.967 ± 1.127 0.113 ± 0.091 0.978 ± 0.203 0.016 ± 0.012 0.214 ± 0.044 0.004 ± 0.003
Phi4 6.610 ± 1.624 0.785 ± 0.598 1.498 ± 0.293 0.116 ± 0.082 0.330 ± 0.064 0.025 ± 0.017
Mistral0.3-7B 5.308 ± 1.367 0.461 ± 0.434 1.065 ± 0.288 0.067 ± 0.061 0.232 ± 0.061 0.015 ± 0.013
Qwen2.5-14B 4.829 ± 2.179 1.736 ± 2.659 0.979 ± 0.264 0.241 ± 0.372 0.214 ± 0.061 0.051 ± 0.077
DeepSeekR1Q-14B 4.477 ± 2.176 1.424 ± 1.752 0.830 ± 0.256 0.200 ± 0.242 0.181 ± 0.058 0.044 ± 0.056
Downstream Evaluation.
While MSE is a reasonable proxy for overall distortion, the ultimate question is whether the quantized KV cache preserves model quality on realistic tasks. We therefore evaluate the final model performance on GSM8K in two prompting modes: 1-shot and 8-shot prompting using 4 quantization settings:
  • $\mathbf{K}_{\mathbf{(2bit)}}\mathbf{V}_{\mathbf{(2bit)}}$: Aggressive uniform quantization.
  • $\mathbf{K}_{\mathbf{(4bit)}}\mathbf{V}_{\mathbf{(4bit)}}$: More conservative uniform quantization.
  • $\mathbf{K}_{\mathbf{(2bit)}}\mathbf{V}_{\mathbf{(4bit)}}$ vs. $\mathbf{K}_{\mathbf{(4bit)}}\mathbf{V}_{\mathbf{(2bit)}}$: Mixed-precision assignments, giving 2 bits to either $\mathbf{K}$ or $\mathbf{V}$ and 4 bits to the other.
4.2 Experiment Results
Table 2: GSM8K 1-shot and 8-shot performance of various models under different key-value quantization settings. $\mathbf{K}_{i}\mathbf{V}_{j}$ denotes $i$-bit quantization for the K cache and $j$-bit for the V cache. The $\mathbf{K}_{\mathbf{(4bit)}}\mathbf{V}_{\mathbf{(2bit)}}$ column, highlighted below, represents our main finding.
1-shot 8-shot
Model $\mathbf{K}_{\mathbf{2}}\mathbf{V}_{\mathbf{2}}$ $\mathbf{K}_{\mathbf{2}}\mathbf{V}_{\mathbf{4}}$ $\mathbf{K}_{\mathbf{4}}\mathbf{V}_{\mathbf{2}}$ $\mathbf{K}_{\mathbf{4}}\mathbf{V}_{\mathbf{4}}$ $\mathbf{K}_{\mathbf{2}}\mathbf{V}_{\mathbf{2}}$ $\mathbf{K}_{\mathbf{2}}\mathbf{V}_{\mathbf{4}}$ $\mathbf{K}_{\mathbf{4}}\mathbf{V}_{\mathbf{2}}$ $\mathbf{K}_{\mathbf{4}}\mathbf{V}_{\mathbf{4}}$
Llama3.2-1B-it 0.033 0.035 0.338 0.357 0.031 0.031 0.289 0.369
Llama3.1-8B-it 0.511 0.547 0.752 0.754 0.408 0.441 0.770 0.782
Phi4-14B 0.759 0.783 0.913 0.923 0.771 0.815 0.927 0.931
DeepSeekR1Q-14B 0.772 0.775 0.865 0.867 0.763 0.792 0.876 0.875
Table 1 shows that across datasets and models, quantization error for the key matrices is consistently and substantially higher than that for the value matrices. For instance, in the MMLU dataset, Llama3.2-1B exhibits an error of $4.85\pm 1.04$ for $\mathbf{K}_{2}$ versus only $0.13\pm 0.10$ for $\mathbf{V}_{2}$. This trend persists for both the 3-bit and 4-bit settings, where increasing the bit-width reduces the error for both matrices, yet the keys still incur a notably higher error. Similar patterns are evident in the C4 and GSM8K datasets, highlighting that key matrices are inherently more sensitive to quantization noise. Furthermore, Figure 6 in the Appendix B illustrates the quantization error per attention head in each layer for the Llama3.1-8B model on C4, MMLU, and GSM8K, reinforcing our observation (see Appendix D for extended results). Table 2 shows accuracy on GSM8K under 1-shot and 8-shot prompting. Uniform K${}_{(2\text{bit})}$${}_{(2\text{bit})}$ settings degrade accuracy noticeably, especially in the longer 8-shot context. By contrast, K${}_{(4\text{bit})}$${}_{(2\text{bit})}$ nearly matches the unquantized baseline, confirming that providing higher precision to keys preserves model quality while still compressing values aggressively. The reversed assignment K${}_{(2\text{bit})}$${}_{(4\text{bit})}$ consistently underperforms, underscoring that allocating fewer bits to key matrices adversely affects accuracy.
5 Related Work
Model Quantization.
for large models such as LLMs Wan et al. (2024), which is why post-training quantization (PTQ) techniques are generally preferred for quantizing these models.
Quantization reduces the computational cost of neural network inference by lowering model bit-precision Han et al. (2016); Nagel et al. (2021). It can be divided into Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT). PTQ requires no re-training or labeled data Jacob et al. (2017); Nagel et al. (2019) and is preferred for large models like LLMs, whereas QAT demands fine-tuning with labeled data Nagel et al. (2021); Gholami et al. (2022) but scales poorly for such models Wan et al. (2024).
LLMs Quantization.
Researchers consider three post-training quantization settings for LLMs: weight-only, weight-activation, and KV cache quantization. Weight-only quantization focuses solely on weights Frantar et al. (2023a, b); Lin et al. (2024), while weight-activation quantization targets both weights and activations Dettmers et al. (2022); Xiao et al. (2023); Shao et al. (2024). Since longer inputs inflate the KV cache’s memory overhead Pope et al. (2023), KV cache quantization offers memory savings similar to weight-activation methods while maintaining performance near that of weight-only quantization Yue et al. (2024).
KV Cache Quantization.
Recent KV cache quantization methods reduce LLM memory and accelerate inference, falling into three categories: outlier redistribution, fixed-precision, and mixed-precision Li et al. (2025). Outliers in KV cache are addressed by redistributing them or applying smoothing transformations Xiao et al. (2023); Lin et al. (2024); Shao et al. (2024). Fixed-precision methods use a uniform bit-width Yao et al. (2022); Sheng et al. (2023); Liu et al. (2024) but can overlook token importance and outliers. In contrast, mixed-precision allocates higher precision to critical tokens and lower precision elsewhere Hooper et al. (2024); Yue et al. (2024); Li et al. (2024); Dong et al. (2024); Duanmu et al. (2024). Existing methods like SKVQ and QAQ rely on specific engineering solutions or predefined parameters, whereas KV-AdaQuant uses a principled, analysis-driven framework that leverages the intrinsic properties of key and value caches for a more robust and adaptable quantization approach.
6 Conclusion
In this work, we systematically analyze KV cache matrices from the perspectives of linear algebra and information theory to uncover distinct patterns in the information carried by these matrices. Building on the singular value distribution of the KV cache, we propose an information-aware quantization framework for KV cache compression. We evaluate the effectiveness of our approach on two widely used datasets, GSM8K, demonstrating its practical impact. While our current study focuses on the per-token quantization methods, the proposed framework is general and can be applied to any key-value quantization technique.
Limitations
Although our information-aware quantization framework is effective, there are aspects that warrant further consideration. First, our experiments utilize a maximum context length of 2K tokens. While this setting aligns with many practical inference scenarios, it may not fully capture the demands of tasks requiring larger context windows (e.g., 4K or 8K tokens). Extending our approach to support longer contexts could reveal additional challenges, such as increased quantization sensitivity and greater memory management overhead.
Ethical Considerations
Our research aims to reduce the memory footprint and computational overhead of LLM in inference time through KV cache quantization. Such advancements can potentially lower energy consumption and expand access to resource-constrained communities, contributing to more environmentally sustainable and inclusive deployment of language modeling based productions. However, the use of proposal KV cache compression method must be carefully monitored to ensure that accuracy trade-offs do not disproportionately affect critical applications such as healthcare, legal, and financial contexts.
Appendix A Quantization and Dequantization Error Bounds
In this appendix we establishes upper bounds on the quantization error incurred when a matrix is represented using a finite bit-width in two’s-complement format. Two theorems are presented: one that characterizes the error in terms of the spectral norm and another in terms of the Frobenius norm.
The first theorem demonstrates that the spectral norm of the quantization error is bounded by
$\|A-\widehat{A}\|_{2}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{2}.$
This result implies that, for a given bit-width $b$, matrices with larger spectral norms incur proportionally larger quantization errors. Consequently, a matrix that exhibits a larger spectral norm is more susceptible to quantization errors. To control the error propagation in such matrices, a higher bit-width is necessary.
The second theorem provides an analogous bound for the Frobenius norm,
$\|A-\widehat{A}\|_{F}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{F}.$
Similar to the spectral norm result, this bound indicates that the quantization error, measured in the Frobenius norm, is directly proportional to the norm of the original matrix. Hence, matrices with larger Frobenius norms are also more vulnerable to quantization errors and would benefit from a higher precision during quantization.
Preliminaries and Notation
Let $A\in\mathbb{R}^{m\times n}$ denote a real matrix whose entries are to be stored using a fixed number of bits in two’s-complement representation. The following notation is adopted:
  • Bit Depth. Given $b$ bits in two’s-complement format, each representable integer $q$ lies in the interval
    $q\in\{-2^{b-1},-2^{b-1}+1,\dots,2^{b-1}-1\}.$
  • Maximum Entry Magnitude. Define
[…]
  • Scale Factor. Set
    $\alpha\;=\;\frac{M}{2^{b-1}-1}.$
    This choice ensures that the scaled entries $A_{ij}/\alpha$ lie within the representable range.
Quantization. Define the integer matrix $Q\in\mathbb{Z}^{m\times n}$ by
$Q_{ij}$ $=\mathrm{round}\!\Bigl{(}\frac{A_{ij}}{\alpha}\Bigr{)},$ (3)
$\text{with}\quad Q_{ij}$$\in\{-2^{b-1},\dots,2^{b-1}-1\}.$
Dequantization (Reconstruction). The reconstructed matrix $\widehat{A}$ is given by
$\widehat{A}_{ij}\;=\;\alpha\,Q_{ij}.$
Objective. The aim is to bound the errors
$\|A-\widehat{A}\|_{2}\quad\text{and}\quad\|A-\widehat{A}\|_{F},$
in terms of $b$, $m$, $n$, and the norms of $A$.
Spectral Norm Bound
Theorem 1 (Spectral Norm Quantization Error).
Let $A\in\mathbb{R}^{m\times n}$ and $b\in\mathbb{N}$ be given. Consider two’s-complement quantization with scale factor
$\alpha\;=\;\frac{M}{2^{b-1}-1},\quad M\;=\;\max_{i,j}|A_{ij}|.$
Define
$Q_{ij}\;=\;\mathrm{round}\!\Bigl{(}\frac{A_{ij}}{\alpha}\Bigr{)}\quad\text{and}\quad\widehat{A}_{ij}\;=\;\alpha\,Q_{ij}.$
Then, the following bound holds:
$M$ $\leq\|A-\widehat{A}\|_{2}\leq\sqrt{m\,n}\,\frac{M}{2(2^{b-1}-1)}$ (4)
$\leq\sqrt{m\,n}\,\frac{\|A\|_{2}}{2(2^{b-1}-1)}.$
In approximate form for large $b$,
$\|A-\widehat{A}\|_{2}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{2}.$
Proof.
Entrywise Bound. By construction,
$\left|\frac{A_{ij}}{\alpha}-Q_{ij}\right|\;\leq\;\frac{1}{2}.$
Multiplying by $\alpha$ gives
$|A_{ij}-\widehat{A}_{ij}|\;\leq\;\frac{\alpha}{2}\;=\;\frac{M}{2\,(2^{b-1}-1)}.$
Thus,
$\max_{i,j}\,|A_{ij}-\widehat{A}_{ij}|\;\leq\;\frac{M}{2\,(2^{b-1}-1)}.$
Conversion to the Spectral Norm. Using the inequality
$\|B\|_{2}\;\leq\;\sqrt{m\,n}\,\max_{i,j}|B_{ij}|,$
with $B=A-\widehat{A}$, it follows that
$\|A-\widehat{A}\|_{2}\;\leq\;\sqrt{m\,n}\,\frac{M}{2\,(2^{b-1}-1)}.$
Relating $M$ to $\|A\|_{2}$. Since
$M\;=\;\max_{i,j}|A_{ij}|\;\leq\;\|A\|_{2},$
the bound can be written as
$\|A-\widehat{A}\|_{2}\;\leq\;\sqrt{m\,n}\,\frac{\|A\|_{2}}{2\,(2^{b-1}-1)}.$
For large $b$, where $2^{b-1}-1\approx 2^{b-1}$, the bound becomes
$\|A-\widehat{A}\|_{2}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{2}.$
Frobenius Norm Bound
Theorem 2 (Frobenius Norm Quantization Error).
Under the same setup, the Frobenius norm of the quantization error satisfies
$\begin{split}\|A-\widehat{A}\|_{F}&\;\leq\;\sqrt{m\,n}\,{\textstyle\frac{M}{2(2^{b-1}-1)}}\\ &\;\leq\;\sqrt{m\,n}\,{\textstyle\frac{\|A\|_{F}}{2(2^{b-1}-1)}}.\end{split}$ (5)
In approximate form,
$\|A-\widehat{A}\|_{F}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{F}.$
Proof.
Entrywise Bound. As established,
$|A_{ij}-\widehat{A}_{ij}|\;\leq\;\frac{M}{2\,(2^{b-1}-1)}\quad\text{for all }i,j.$
Conversion to the Frobenius Norm. By definition,
$\|A-\widehat{A}\|_{F}^{2}\;=\;\sum_{i=1}^{m}\sum_{j=1}^{n}(A_{ij}-\widehat{A}_{ij})^{2},$
which yields
$\|A-\widehat{A}\|_{F}^{2}\;\leq\;m\,n\left(\frac{M}{2\,(2^{b-1}-1)}\right)^{2}.$
Taking square roots leads to
$\|A-\widehat{A}\|_{F}\;\leq\;\sqrt{m\,n}\,\frac{M}{2\,(2^{b-1}-1)}.$
Relating $M$ to $\|A\|_{F}$. Since
$M^{2}\;\leq\;\sum_{i=1}^{m}\sum_{j=1}^{n}A_{ij}^{2}\;=\;\|A\|_{F}^{2},$
it follows that $M\leq\|A\|_{F}$ and hence
$\|A-\widehat{A}\|_{F}\;\leq\;\sqrt{m\,n}\,\frac{\|A\|_{F}}{2\,(2^{b-1}-1)}.$
For large $b$, this simplifies to
$\|A-\widehat{A}\|_{F}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{F}.$
Remark. The results indicate that both the spectral norm and Frobenius norm errors satisfy similar approximate bounds:
$\|A-\widehat{A}\|_{2}$ $\lesssim\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{2},$ (6)
$\|A-\widehat{A}\|_{F}$ $\lesssim\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{F}.$
Quantizing KV Cache
Consider the key and value cache
$\mathbf{V}\in\mathbb{R}^{L\times d_{\text{head}}},\quad\mathbf{K}\in\mathbb{R}^{L\times d_{\text{head}}},$
with quantization bit-widths denoted by $b_{V}$ and $b_{K}$, respectively. Let $\widehat{\mathbf{V}}$ and $\widehat{\mathbf{K}}$ denote the dequantized matrices, and define the quantization errors as
$\mathbf{E}_{V}\;=\;\mathbf{V}-\widehat{\mathbf{V}},\quad\mathbf{E}_{K}\;=\;\mathbf{K}-\widehat{\mathbf{K}}.$
An empirical observation is that
$\|\mathbf{V}\|_{\ast}\;<\;\|\mathbf{K}\|_{\ast},$
where $\ast$ denotes either the spectral norm ($\|\cdot\|_{2}$) or the Frobenius norm ($\|\cdot\|_{F}$). In practice, $\mathbf{K}$ typically exhibits a larger norm than $\mathbf{V}$.
Spectral-Norm Perspective
Standard quantization error bounds yield
$\|\mathbf{E}_{V}\|_{2}$ $\lesssim\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{V}}}\;\|\mathbf{V}\|_{2},$ (7)
$\|\mathbf{E}_{K}\|_{2}$ $\lesssim\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{K}}}\;\|\mathbf{K}\|_{2}.$
To achieve comparable spectral-norm errors (i.e., $\|\mathbf{E}_{V}\|_{2}\approx\|\mathbf{E}_{K}\|_{2}$), it is necessary that
$\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{V}}}\;\|\mathbf{V}\|_{2}\;\approx\;\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{K}}}\;\|\mathbf{K}\|_{2}.$
Cancelling the common factor $\sqrt{L\,d_{\text{head}}}$ yields
[…]
or equivalently,
$2^{\,b_{K}-b_{V}}\;\approx\;\frac{\|\mathbf{V}\|_{2}}{\|\mathbf{K}\|_{2}}.$
Since $\|\mathbf{V}\|_{2}<\|\mathbf{K}\|_{2}$, it follows that $b_{K}>b_{V}$.
Frobenius-Norm (MSE) Perspective
Similarly, the Frobenius norm error bounds are given by
$\|\mathbf{E}_{V}\|_{F}$ $\;\lesssim\;\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{V}}}\;\|\mathbf{V}\|_{F},$ (8)
$\|\mathbf{E}_{K}\|_{F}$ $\;\lesssim\;\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{K}}}\;\|\mathbf{K}\|_{F}.$
To achieve comparable Frobenius errors,
$\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{V}}}\;\|\mathbf{V}\|_{F}\;\approx\;\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{K}}}\;\|\mathbf{K}\|_{F},$
which implies
[…]
and therefore,
$2^{\,b_{K}-b_{V}}\;\approx\;\frac{\|\mathbf{V}\|_{F}}{\|\mathbf{K}\|_{F}}.$
Again, because $\|\mathbf{V}\|_{F}<\|\mathbf{K}\|_{F}$, it follows that $b_{K}>b_{V}$.
Appendix B Quantization Error Across Layers and Heads
Figure 6 displays the MSE of the quantized key and value caches in the Llama 3.1 8B model, evaluated layer by layer across 32 layers and differentiated by attention head. Results are presented for three datasets: C4, MMLU, and GSM8k. Each subfigure illustrates the 2-bit and 4-bit quantization configurations for the key and value caches. The error patterns vary across layers, suggesting that certain layers or heads may be more sensitive to aggressive quantization; however, the overall pattern indicates that assigning 4-bit to the key cache is comparable to assigning 2-bit to the value cache.
(a) C4 dataset
(b) MMLU dataset
(c) GSM8k dataset
Figure 6: Quantization error (MSE) of K and V caches in the Llama 3.1 8B model across 32 layers for the (a) C4, (b) MMLU, and (c) GSM8k datasets. The top row in each plot shows 2-bit quantization (k2, v2), while the bottom row shows 4-bit quantization (k4, v4). Each point represents a different attention head in the corresponding layer. The x-axis indicates the layer index, and the y-axis shows the quantization error in MSE.
Appendix C Extended Visualization of Singular Value Distributions
Figure 7 provides the full-range view of the singular value distribution for the key and value cache at various layers of the Llama 3.3 70B model on the C4 dataset. The horizontal axis enumerates the singular values in descending order, beginning with the largest (i.e., the spectral norm), and the vertical axis measures the corresponding magnitudes. The shaded region captures the minimum-to-maximum range of singular values across the attention heads within each layer, while the solid curves depict the mean singular value at each rank.
This figure underscores two main observations. First, key matrices systematically exhibit higher leading singular values than value matrices. In particular, the first singular value (the spectral norm) for key matrices is consistently larger, confirming that the key matrices possess a higher spectral norm. Second, this behavior (significant gap) remains evident across the entire spectrum of singular values, implying that the sum of the squares of these values (i.e., the Frobenius norm) is also larger for key matrices. Consequently, the stronger representational capacity of key matrices, as reflected by both their spectral and Frobenius norms, renders them more sensitive to quantization errors.
Appendix D Extended Quantization Error Results
Tables 3, 4, and 5 present extended quantization error results across a wider range of models and settings for the MMLU, C4, and GSM8K datasets, respectively. These tables offer a detailed breakdown of the error metrics under various bit-width configurations, complementing the analysis presented in the main text and in Table 1.
Figure 7: Complete singular value distribution of key and value cache for the Llama 3.3 70B model on the C4 dataset. The x-axis denotes the singular value indices, ordered from the largest (spectral norm) to the smallest, while the y-axis represents the corresponding magnitudes. The shaded region illustrates the range between the minimum and maximum singular values across attention heads within each layer, and the solid lines indicate the mean singular value magnitude at each index. This full-spectrum view highlights that key matrices consistently maintain significantly higher singular values throughout the entire distribution, further reinforcing their dominant representational capacity compared to value matrices.
Table 3: Quantization error (mean ± std) for the key and value caches ($K_{i}$ and $V_{i}$) at 2-bit, 3-bit, and 4-bit quantization, evaluated on the MMLU dataset.
Model K2 V2 K3 V3 K4 V4
Llama3.2-1B 4.851 ± 1.037 0.127 ± 0.101 1.037 ± 0.265 0.021 ± 0.015 0.227 ± 0.059 0.005 ± 0.003
Llama3.2-1B-it 4.373 ± 1.034 0.124 ± 0.090 0.879 ± 0.218 0.019 ± 0.013 0.192 ± 0.047 0.004 ± 0.003
Llama3.2-3B 3.943 ± 0.924 0.193 ± 0.096 0.849 ± 0.150 0.030 ± 0.015 0.183 ± 0.031 0.007 ± 0.003
Llama3.2-3B-it 4.487 ± 1.180 0.202 ± 0.100 0.894 ± 0.180 0.030 ± 0.015 0.193 ± 0.037 0.007 ± 0.003
Llama2-7B 3.190 ± 0.783 0.259 ± 0.184 0.769 ± 0.194 0.042 ± 0.030 0.168 ± 0.042 0.009 ± 0.006
Llama3.1-8B-it 6.003 ± 1.782 0.187 ± 0.127 1.082 ± 0.244 0.028 ± 0.019 0.235 ± 0.055 0.006 ± 0.004
Llama3.3-70B-it 4.883 ± 1.106 0.112 ± 0.093 0.942 ± 0.198 0.016 ± 0.012 0.206 ± 0.043 0.003 ± 0.003
Memotron3.1-it 5.125 ± 1.284 0.114 ± 0.094 0.985 ± 0.207 0.016 ± 0.012 0.216 ± 0.046 0.003 ± 0.003
Phi3-Medium-128K-it 5.063 ± 1.914 0.584 ± 0.559 1.000 ± 0.319 0.087 ± 0.083 0.217 ± 0.068 0.019 ± 0.018
Phi4 5.929 ± 1.545 0.657 ± 0.472 1.306 ± 0.231 0.103 ± 0.070 0.286 ± 0.050 0.022 ± 0.015
Mistral0.3-7B 4.718 ± 1.340 0.398 ± 0.405 0.941 ± 0.240 0.059 ± 0.059 0.206 ± 0.053 0.013 ± 0.013
Qwen2.5-14B 5.184 ± 2.241 1.270 ± 1.547 1.005 ± 0.288 0.182 ± 0.221 0.223 ± 0.067 0.040 ± 0.052
DeepSeekR1L-8B 5.502 ± 1.549 0.189 ± 0.118 0.955 ± 0.204 0.028 ± 0.017 0.209 ± 0.046 0.006 ± 0.004
DeepSeekR1Q-14B 5.126 ± 2.375 1.406 ± 1.609 0.900 ± 0.269 0.198 ± 0.226 0.199 ± 0.062 0.044 ± 0.052
Table 4: Quantization error (mean ± std) for the key and value caches ($K_{i}$ and $V_{i}$) at 2-bit, 3-bit, and 4-bit quantization, evaluated on the C4 dataset.
Model K2 V2 K3 V3 K4 V4
Llama3.2-1B 4.885 ± 1.056 0.207 ± 0.166 1.074 ± 0.289 0.030 ± 0.024 0.233 ± 0.062 0.006 ± 0.005
Llama3.2-1B-it 4.524 ± 1.108 0.193 ± 0.137 0.925 ± 0.235 0.028 ± 0.020 0.201 ± 0.050 0.006 ± 0.005
Llama3.2-3B 3.885 ± 0.777 0.282 ± 0.150 0.909 ± 0.168 0.042 ± 0.023 0.194 ± 0.032 0.009 ± 0.005
Llama3.2-3B-it 4.135 ± 1.088 0.274 ± 0.137 0.912 ± 0.176 0.039 ± 0.020 0.195 ± 0.036 0.009 ± 0.004
Llama2-7B 6.337 ± 1.710 0.456 ± 0.247 1.054 ± 0.263 0.071 ± 0.038 0.213 ± 0.052 0.015 ± 0.008
Llama3.1-8B-it 6.262 ± 1.789 0.254 ± 0.185 1.128 ± 0.249 0.036 ± 0.026 0.247 ± 0.056 0.008 ± 0.005
Llama3.3-70B-it 4.391 ± 1.027 0.121 ± 0.097 0.847 ± 0.175 0.017 ± 0.013 0.186 ± 0.038 0.004 ± 0.003
Memotron3.1-it 5.367 ± 1.332 0.127 ± 0.105 1.049 ± 0.222 0.018 ± 0.013 0.231 ± 0.049 0.004 ± 0.003
Phi3-Medium-128K-it 4.831 ± 1.759 0.788 ± 0.726 1.022 ± 0.306 0.109 ± 0.097 0.220 ± 0.064 0.023 ± 0.021
Phi4 5.715 ± 1.442 0.850 ± 0.684 1.316 ± 0.245 0.124 ± 0.093 0.291 ± 0.056 0.027 ± 0.020
Mistral0.3-7B 5.027 ± 1.332 0.543 ± 0.493 1.014 ± 0.269 0.079 ± 0.068 0.223 ± 0.060 0.017 ± 0.015
Qwen2.5-14B 4.382 ± 2.170 1.544 ± 1.872 0.846 ± 0.250 0.220 ± 0.265 0.187 ± 0.060 0.048 ± 0.060
DeepSeekR1L-8B 4.575 ± 1.122 0.204 ± 0.134 0.817 ± 0.141 0.030 ± 0.019 0.179 ± 0.033 0.006 ± 0.004
DeepSeekR1Q-14B 4.832 ± 2.354 1.651 ± 1.914 0.927 ± 0.283 0.232 ± 0.267 0.201 ± 0.061 0.051 ± 0.060
Table 5: Quantization error (mean ± std) for the key and value caches ($K_{i}$ and $V_{i}$) at 2-bit, 3-bit, and 4-bit quantization, evaluated on the GSM8K dataset.
Model K2 V2 K3 V3 K4 V4
Llama3.2-1B 5.703 ± 1.557 0.179 ± 0.136 1.213 ± 0.352 0.026 ± 0.020 0.266 ± 0.078 0.005 ± 0.004
Llama3.2-1B-it 5.002 ± 1.383 0.171 ± 0.130 1.024 ± 0.287 0.025 ± 0.020 0.223 ± 0.061 0.006 ± 0.004
Llama3.2-3B 4.840 ± 1.396 0.261 ± 0.136 1.045 ± 0.211 0.038 ± 0.021 0.226 ± 0.044 0.008 ± 0.005
Llama3.2-3B-it 3.604 ± 0.850 0.226 ± 0.129 0.790 ± 0.135 0.034 ± 0.019 0.171 ± 0.028 0.007 ± 0.004
Llama2-7B 5.081 ± 1.396 0.405 ± 0.231 0.969 ± 0.238 0.065 ± 0.037 0.205 ± 0.050 0.014 ± 0.008
Llama3.1-8B-it 6.445 ± 1.837 0.213 ± 0.161 1.184 ± 0.268 0.030 ± 0.022 0.257 ± 0.060 0.007 ± 0.005
Llama3.3-70B-it 4.967 ± 1.127 0.113 ± 0.091 0.978 ± 0.203 0.016 ± 0.012 0.214 ± 0.044 0.004 ± 0.003
Memotron3.1-it 4.752 ± 1.124 0.113 ± 0.089 0.940 ± 0.194 0.016 ± 0.012 0.206 ± 0.042 0.004 ± 0.003
Phi3-Medium-128K-it 4.940 ± 1.834 0.605 ± 0.579 1.042 ± 0.320 0.088 ± 0.082 0.227 ± 0.069 0.019 ± 0.018
Phi4 6.610 ± 1.624 0.785 ± 0.598 1.498 ± 0.293 0.116 ± 0.082 0.330 ± 0.064 0.025 ± 0.017
Mistral0.3-7B 5.308 ± 1.367 0.461 ± 0.434 1.065 ± 0.288 0.067 ± 0.061 0.232 ± 0.061 0.015 ± 0.013
Qwen2.5-14B 4.829 ± 2.179 1.736 ± 2.659 0.979 ± 0.264 0.241 ± 0.372 0.214 ± 0.061 0.051 ± 0.077
DeepSeekR1L-8B 5.547 ± 1.517 0.193 ± 0.129 1.000 ± 0.212 0.028 ± 0.018 0.218 ± 0.049 0.006 ± 0.004
DeepSeekR1Q-14B 4.477 ± 2.176 1.424 ± 1.752 0.830 ± 0.256 0.200 ± 0.242 0.181 ± 0.058 0.044 ± 0.056

### v2

@@
-More for Keys, Less for Values: Adaptive KV Cache Quantization
+Quantize What Counts: Bit Allocation Insights Informed by Spectral Gaps in Keys and Values
 Abstract
-This paper introduces an information-aware quantization framework that adaptively compresses the key-value (KV) cache in large language models (LLMs).
-Although prior work has underscored the distinct roles of key and value cache during inference, our systematic analysis—examining singular value distributions, spectral norms, and Frobenius norms—reveals, for the first time, that key matrices consistently exhibit higher norm values and are more sensitive to quantization than value matrices.
-Furthermore, our theoretical analysis shows that matrices with higher spectral norms amplify quantization errors more significantly.
-Motivated by these insights, we propose a mixed-precision quantization strategy, KV-AdaQuant, which allocates more bit-width for keys and fewer for values since key matrices have higher norm values.
-With the same total KV bit budget, this approach effectively mitigates error propagation across transformer layers while achieving significant memory savings.
-Our extensive experiments on multiple LLMs (1B - 70B) demonstrate that our mixed-precision quantization scheme maintains high model accuracy even under aggressive compression.
-For instance, using 4-bit for Key and 2-bit for Value achieves an accuracy of 75.2%, whereas reversing the assignment—2-bit for Key and 4-bit for Value—yields only 54.7% accuracy.
-The code is available at https://tinyurl.com/kv-adaquant.
-More for Keys, Less for Values: Adaptive KV Cache Quantization
-Mohsen Hariri1, Lam Nguyen1, Sixu Chen1, Shaochen Zhong2, Qifan Wang3, Xia Hu2, Xiaotian Han1, Vipin Chaudhary1 1Case Western Reserve University; 2Rice University; 3Meta mohsen.hariri@case.edu
+Large Language Models (LLMs) have introduced significant advancements to the capabilities of Natural Language Processing (NLP) in recent years.
+However, as these models continue to scale in size, memory constraints pose substantial challenge.
+Key and Value cache (KV cache) quantization has been well-documented as a promising solution to this limitation.
+In this work, we provide two novel theorems aimed at enhancing KV quantization methods.
+Our first theorem, termed Key-Value Norm Disparity, states that the key weight matrices by nature carry richer information compared to the value weight matrices, as evidenced by higher spectral and Frobenius norms across most of the layers.
+Our second theorem, Key-Driven Quantization, posits that prioritizing the quantization precision of keys over values induces significant improvements to the overall quantization performance.
+In particular, assigning greater precision to the keys compared to the values achieves a higher degree of precision reduction with minimal impact on model accuracy.
+We validate these theorems through theory and extensive experiments on several state-of-the-art LLM architectures and benchmarks.
+These findings offer valuable guidelines for improving KV cache quantization strategies, facilitating more efficient memory utilization without compromising model performance across diverse NLP tasks.
+Source code is available at https://github.com/mohsenhariri/spectral-kv.
 1 Introduction
+LLMs have introduced significant advancements to the capabilities of NLP via their ability to express and capture intricate ideas [1, 2], and a key reason for these advancements is model scaling.
+GPT has seen a tremendous increase across its four versions: GPT-1 at 117M [3], GPT-2 at 1.5B [4], GPT-3 at 175B [5], and GPT-4 at 1.8T parameters [6].
+In the open-source space, Llama has grown to 2T [7], Mistral Large is now 123B [8], and DeepSeek V3 is at 671B [9].
+However, with this growth in model size comes a prohibitive memory bottleneck in the KV cache [10].
+An increase in parameter count requires an increase in context-window length to accommodate chain-of-thought reasoning, which in turn expands the KV cache central to the attention unit.
+That expansion can impose a restrictive memory bottleneck on the LLM.
+Already, supported context windows in SOTA systems have grown significantly [11, 12, 13, 14], underscoring the need for a solution.
+One powerful remedy is quantization, wherein high-precision numbers are reduced to low-precision formats [15, 16].
+In LLMs, this can be applied to the KV-cache tensors (e.g., from BF16 to INT4) yielding substantial memory savings with minimal performance loss.
 Size K2V4 K4V2
@@
 14B 0.76 0.87
-Figure 1: Key cache needs more bits.
-(Left): Spectral norms of the key cache (blue) and value cache (orange) across layers in Llama3.3-70B show that key caches consistently exhibit higher norms.
-(Right): GSM8k accuracy for two schemes: $\text{K}_{2}\text{V}_{4}$, representing 2-bit allocation for the K cache and 4-bit allocation for the V cache and $\text{K}_{4}\text{V}_{2}$, representing 4-bit allocation for the K cache and 2-bit allocation for the V cache, demonstrates that allocating more bits to the key cache maintains strong performance, confirming the efficacy of norm-aware, mixed-precision quantization.
-Large language models (LLMs) have advanced NLP by enabling parallel processing and capturing complex dependencies Vaswani et al.
-(2017); Sutskever et al.
-(2014), which have scaled from GPT-1’s 117M Radford et al.
-(2018), GPT-2’s 1.5B Radford et al.
-(2019) to GPT-3’s 175B Brown et al.
-(2020) and GPT-4’s 1.8T parameters OpenAI et al.
-(2024).
-Open-source models like Llama have grown to 405B Touvron et al.
-(2023), with Mistral Large (123B) Mistral-AI (2023) and DeepSeek V3 (671B) DeepSeek-AI (2023) also emerging.
-In parallel with increasing parameter counts, longer context windows are becoming essential for chain-of-thought reasoning, necessitating an expansion of the KV cache and contributing to memory bottlenecks Wei et al.
-(2022).
-New systems support context windows of up to 1M tokens DeepMind (2025), 200k tokens OpenAI (2024, 2025), and 130k tokens DeepSeek-AI et al.
-(2025).
-NVIDIA’s Blackwell architecture introduces low-precision formats like FP8, FP4, and INT4 NVIDIA (2025), emphasizing the need for precise quantization methods.
-The KV cache is a critical component in these models and its effective operation is paramount and significant.
-To accommodate larger context windows, the KV cache can become a significant memory bottleneck.
-One effective solution is to quantize the key and value caches from high-precision formats (e.g., FP32 or BF16) to lower-precision ones (e.g., INT4 or INT2).
-However, we first discovered that the Key cache generally exhibits higher norms than the Value cache, making it more sensitive to uniform quantization.
-Since matrices with larger norms are more prone to quantization errors, bit-width allocation should be tailored to each matrix’s norm distribution.
-To confirm this, we conducted experiments with Llama3.3-70B (see Figures 1 (Left) and 2), as the Key cache (blue) consistently shows elevated norm values compared to the Value cache (orange).
-Furthermore, Figure 1 (Right) demonstrates that the K(4bit)V(2bit) scheme significantly outperforms the K(2bit)V(4bit) configuration on GSM8K dataset.
-These results underscore the need for a norm-aware, mixed-precision strategy that allocates higher precision to Keys and lower precision to Values.
-Guided by these inspiring observations, we propose a mixed-precision quantization strategy, KV-AdaQuant, which allocates different bit-width to more bits for keys and fewer bits for values.
-since key matrices consistently have higher norm values.
-With the same total KV bit budget, this approach effectively mitigates error propagation across transformer layers while achieving significant memory savings.
-KV-AdaQuant uses singular values, spectral norms, and Frobenius norms to assign more bits to Keys and fewer to Values, balancing memory savings with preserving critical information.
-This work makes the following contributions: Guided by these insights, we propose a mixed-precision quantization strategy, KV-AdaQuant, which assigns a higher bit-width to key caches and a lower bit-width to value caches—reflecting the consistently higher norm values observed in key matrices.
-KV-AdaQuant leverages metrics such as singular values, spectral norms, and Frobenius norms to identify the critical information in KV cache.
-By maintaining the same overall KV bit budget, our approach effectively mitigates error propagation across transformer layers while offering significant memory savings.
-Our contributions include:
-  • We systematically investigate the KV cache in transformers, uncovering structural patterns in how information is stored and retrieved.
-  • We introduce KV-AdaQuant, an information-aware quantization method that leverages singular value distributions to allocate mixed precision—assigning higher bit-widths to keys and lower bit-widths to values.
-  • We demonstrate the efficacy of our approach on the GSM8K benchmark, achieving substantial memory savings without compromising performance.
-2 Key and Value are Different
-This section investigates the differences between key and value caches to enhance quantization strategies.
-2.1 Singular Value Distribution of the KV Cache
-An extensive collection of models, ranging in size up to 70B parameters (Llama 3.3 70B Grattafiori et al.
-(2024)), is examined to explore the singular values of the KV cache.
-Random batches are sampled from the C4 Dodge et al.
-(2021), GSM8K Cobbe et al.
-(2021), and MMLU Hendrycks et al.
-(2021) datasets, and KV caches are extracted across all attention heads and layers.
-Figure 3 illustrates the distribution of singular values for the KV cache, starting from the 5th largest singular value onward, highlighting how keys and values diverge.
-Additionally, Figure 7 in Appendix C presents the full-range distribution, further revealing the discrepancy in larger singular values.
-Beyond individual singular values, the spectral norm ($\|M\|_{2}$) and the Frobenius norm ($\|M\|_{F}$) are computed, providing empirical evidence of the substantial gap in key and value matrix singular values and illustrating how this characteristic of the KV cache can inform the theoretical analysis of quantization error propagation in multi-layer attention networks.
-This discrepancy underscores the importance of careful precision allocation to mitigate performance degradation while saving more memory in the inference time.
-2.2 Rank of KV Cache
-The rank of a matrix corresponds to the number of nonzero singular values and determines the effective dimension of its column or row space.
-In multi-head self-attention, the key ($K$) and value ($V$) matrices possess one fixed dimension, $d_{\mathrm{head}}$, and one variable dimension corresponding to the sequence length, $\mathrm{seq\_len}$.
-Consequently, $K$ and $V$ can be analyzed in two distinct regimes: (a) when $\mathrm{seq\_len}<d_{\mathrm{head}}$, and (b) when $\mathrm{seq\_len}\geq d_{\mathrm{head}}$.
-In the first regime, the matrices are relatively short and thus offer limited opportunities for compression.
-Accordingly, many existing KV-cache compression methods either avoid compressing in this stage or employ specialized strategies that differ from those used later.
-When $\mathrm{seq\_len}\geq d_{\mathrm{head}}$, the $K$ and $V$ matrices assume rectangular shapes with dimensions $\mathrm{seq\_len}\times d_{\mathrm{head}}$.
-Although the exact rank is data-dependent (since $K=XW_{K}$, where $X$ is drawn from a high-dimensional embedding space), empirical observations suggest that both key and value cache typically approach a rank close to $d_{\mathrm{head}}$.
-Therefore, in practical settings it is reasonable to assume that there are approximately $d_{\mathrm{head}}$ nonzero singular values for both $K$ and $V$.
-In the remainder of this section, we focus on this regime to analyze how spectral and Frobenius norm considerations guide our quantization strategy.
-2.3 Stability of the Dequantized KV Cache: Error Propagation
-The spectral norm of a matrix $M$, denoted $\|M\|_{2}$, quantifies the maximum stretching factor that $M$ can impart on any vector.
-In a multi-layer transformer (e.g., Llama 3.3 Grattafiori et al.
-(2024) with 80 layers or Phi-4 Abdin et al.
-(2024) with 16 layers), the sequential application of transformations implies that quantization errors introduced at one layer can propagate and amplify in subsequent layers.
-To illustrate this effect, consider a simplified layer without normalization or additional scaling Elhage et al. (2021), where the hidden state at layer $(l+1)$ is given by
-$h_{l+1}=h_{l}+W_{l}h_{l},$ (1)
-Let $W_{l}$ denote the weight matrix at layer $l$.
-Define $\widetilde{W}_{l}$ as the result of quantizing the KV cache, and $W_{l}$ as the unquantized version.
-Then, the local error incurred at layer $l$ can be bounded as
-$\begin{split}\|h_{l+1}-\widetilde{h}_{l+1}\|_{2}&=\|(W_{l}-\widetilde{W}_{l})h_{l}\|_{2}\\ &\leq\|W_{l}-\widetilde{W}_{l}\|_{2}\|h_{l}\|_{2}.\end{split}$ (2)
-Propagating such errors through multiple layers yields a cumulative effect that is proportional to the product of the spectral norms at each layer.
-In other words, even small perturbations introduced by quantization can be significantly amplified if the weight matrices exhibit large singular values.
-Empirical observations indicate that key matrices consistently have higher spectral norms than value matrices.
-This suggests that errors in key matrices are more prone to propagate and magnify throughout the transformer stack.
-As a result, it is advisable to employ less aggressive quantization (e.g., 4-bit rather than 2-bit) for key matrices, while value matrices—characterized by lower spectral norms—can tolerate more aggressive quantization without incurring substantial distortion.
-2.4 Total Energy of KV Cache (MSE Perspective)
-Figure 2: Frobenius norm of key and value cache across layers for the Llama 3.3 70B model on the C4 dataset.
-The x-axis represents the layer index.
-The shaded region shows the min-max range across attention heads within each layer, and the solid line indicates the mean norm per layer.
-Key matrices exhibit significantly higher norms compared to value matrices, highlighting their stronger representational capacity.
-The Frobenius norm of a matrix, denoted by $\|M\|_{F}$, is defined as the square root of the sum of the squared entries:
-$\|M\|_{F}=\sqrt{\sum_{i,j}M_{i,j}^{2}}$
-Equivalently, it can be viewed as the square root of the sum of the eigenvalues of $M^{\top}M$, meaning it captures the total “energy” of $M$.
-In the context of mean-squared error (MSE), the Frobenius norm provides a direct indicator of how much overall distortion might occur when quantizing the matrix.
-Quantization Error and the Frobenius Norm.
-When a matrix $M$ is quantized to a lower-precision representation $\widetilde{M}$, the quantization error in MSE terms is often measured by the squared Frobenius norm:
-$\|M-\widetilde{M}\|_{F}^{2}=\sum_{i,j}(M_{i,j}-\widetilde{M}_{i,j})^{2}.$
-A common assumption is that this error scales with both the range of the entries in $M$ and the granularity determined by the bit precision. Formally, one may write
-$\mathbb{E}\bigl{[}\|M-\widetilde{M}\|_{F}^{2}\bigr{]}\;\propto\;\|M\|_{F}^{2}\times 2^{-2b},$
-where $b$ denotes the number of bits used per entry.
-[^1] The proportionality depends on the quantizer design and the distribution of entries in $M$.
-For more details see Appendix A.
-For matrices with the same dimensions (e.g., key and value cache), the MSE and squared Frobenius norm are directly proportional by a factor of $1/mn$.
-Consequently, when analyzing quantization error for matrices of identical size:
-  • Minimizing $\|M-\widetilde{M}\|_{F}^{2}$ is equivalent to minimizing $\text{MSE}(M,\widetilde{M})$, as the proportionality constant $1/mn$ does not affect optimization or relative comparisons.
-  • The scaling relationship $\mathbb{E}\bigl{[}\|M-\widetilde{M}\|_{F}^{2}\bigr{]}\propto\|M\|_{F}^{2}\times 2^{-2b}$ holds equivalently for MSE, since both metrics share the same dependence on $b$ and $\|M\|_{F}^{2}$.
-Thus, the Frobenius norm of the quantization and mean square of the quantization error can be used interchangeably for key and value cache, as they encode the same error structure up to a constant scaling factor.
-Figure 3: Singular value distribution of key and value cache for the Llama 3.3 70B model on the C4 dataset.
+Figure 1: Keys carry greater importance.
+(Left): Keys exhibit higher norms than values across all Llama 3.3-70B layers.
+(Right): GSM8K accuracy is higher under $\mathrm{K}_{2}\mathrm{V}_{4}$ (2-bit keys, 4-bit values) than $\mathrm{K}_{4}\mathrm{V}_{2}$.
+This process is known as KV cache quantization.
+Extensive work has been done in designing novel KV quantization frameworks, yet none so far have provided deep investigation into mixed-precision schemes, in particular those which focus on the bit allocations split between the Key and Value caches.
+Instead, previous works have treated the bit allocation as a hyperparameter, seeing it as an object for optimization via grid search rather than taking a deeper dive into the properties of Keys and Values.
+We provide further motivation for the study in the form of a full description of the current literature, which is detailed in Section 2.
+Therefore, there is a present need for a deep exploration into the inherent differences between the Key and the Value caches, as well as a change to the way that quantization is performed.
+This study aims to fill this need by uncovering fundamental theoretical ideas to improve upon KV cache quantization methods.
+A preview of these results is illustrated in Figure 1 (left), showing that the spectral norms of the Keys consistently exceed those of the values.
+Additionally, Figure 1 (right) demonstrates that accuracy across four different models is notably higher for the K4V2 quantization configuration compared to the K2V4 configuration.
+Our main contributions are summarized as follows:
+  • We prove that the expected values of the spectral norm and Frobenius norm are consistently higher for Key weight matrices compared to Value weight matrices (Key-Value Norm Disparity theorem), implying that Key matrices inherently carry more information.
+We provide detailed empirical validation through observational studies on multiple LLMs.
+  • Building on the Key-Value Norm Disparity, we introduce the Key-Driven Quantization theorem, positing that assigning higher bit precision to Keys over Values significantly enhances quantization efficiency.
+Specifically, this approach allows for greater precision reduction while maintaining minimal accuracy loss.
+  • We substantiate the Key-Driven Quantization theorem through two primary forms of evidence: (1) a theoretical analysis explicitly linking it to the previously established Key-Value Norm Disparity, and (2) comprehensive empirical evaluations involving multiple SOTA LLMs, three datasets, and two quantization frameworks.
+Section 2 addresses gaps in the existing literature on KV cache quantization, followed by Section 3.1, which introduces the Key-Value Norm Disparity theorem and its supporting evidence.
+Section 3.2 presents the Key-Driven Quantization theorem along with its corresponding proof, and Section 4 reports the empirical evaluation of both theorems.
+Finally, Section 5 summarizes the main contributions and discusses their limitations.
+2 Related Work
+KV Quantization Background.
+Quantization frameworks can be categorized in several ways.
+One way is the time at which the quantization is performed, which is grouped into training-time and post-training (PTQ) quantization techniques [17].
+Training-time quantization frameworks can struggle with the usual costs of neural network training, including the dependence on labelled data and longer training times, but these frameworks can achieve more accurate performance due to targeting the weights or activations during the training process.
+In contrast, post-training quantization techniques are the reverse, struggling with sometimes less accurate performance due to dealing with the values after the training has been completed but achieving lower computation overhead and no longer relying on labelled data which may not be available [16].
+There are several settings for PTQ: weight-only, weight-activation, and KV quantization.
+Weight-only quantization focuses on the weights and achieves strong performance but can struggle with memory overhead [18, 19, 20].
+Weight-activation quantization focuses on both the weights and the activations and avoids memory overhead but can struggle with achieving high accuracy [21, 22, 23].
+In contrast, KV quantization offers the memory savings seen in weight-activation methods while also maintaining the performance of weight-only quantization [24].
+This is due to the KV cache memory overhead inflation caused by longer sequences [25].
+KV quantization frameworks can be further categorized into three groups based on the exact quantization scheme being used for the keys and values: outlier redistribution, fixed-precision, and mixed-precision [26].
+In outlier redistribution, the outliers in the cache are moved elsewhere or smoothed out.
+Several outlier redistribution frameworks have been built to varying levels of success [22, 20, 23].
+In fixed-precision, the same number of bits is used to quantize certain parts of the cache while in mixed-precision, different numbers of bits are used.
+Fixed-precision frameworks have also been built to varying levels of success, though they all suffer from a similar problem of neglecting the inherent properties within the KV cache as well as the varying importance of tokens [27, 28, 29].
+More focus has been devoted to the third type of KV quantization framework, mixed-precision, but as yet, these works have not looked much into asymmetric allocation into keys and values [30, 24, 31, 32, 33, 34, 35].
+KVQuant, for example, focuses on defining an outlier for each vector, such as channel or token, to achieve what the authors term dense-and-sparse quantization.
+Only KVTuner, QAQ, and SKVQ discuss the asymmetric allocation in some detail, which is described below.
+KV Quantization Gaps.
+As described above, extensive work has been done in designing new KV quantization frameworks, yet none so far have conducted deeper investigation into mixed-precision schemes.
+In particular, no studies have focused on making a bit allocation split between the key matrix and the value matrix, nor have they provided a deep theoretical explanation for the inherent differences between the two matrices.
+Only three studies to date have touched upon this area, with none providing explanations for potential differences between the key and value matrix importance.
+The closest study has been a KV quantization framework called KVTuner, which employs layerwise quantization based on the sensitivity of the layers to the KV cache quantization mode and precision pairs [35].
+The study makes a brief mention of the difference in attention score errors and perplexity score errors, observing that precision pairs with more bits allocated for keys than for values achieves similar accuracy than for even-split pairs, such as K8V4 versus K8V8.
+However, the study makes no further analysis into this observation, providing no reason for why this difference is observed.
+In addition, the quantization approach does not use any underlying insight on this difference between key and values, instead opting for a straightforward grid search over a limited amount of precision pairs such as K8V4, K4V8, K4V2, and so on.
+Therefore, this allocation split of bits is not the core focus of the framework, with the focus instead being on deploying a layerwise quantization scheme.
+Two other studies also touch upon the key-value difference, but like KVTuner, do not provide much depth into the difference.
+One of these studies is a KV quantization framework called SKVQ, which deploys a sliding window filter along the input tokens to remove tokens that fall outside the window frame and that meet certain criteria [33].
+The quantization strategy achieves 2 bit key quantization and 1.5 bit value quantization, but the authors do not explore the possible reasons behind why this difference in bit allocation yielded such powerful results, instead arriving at this bit allocation via hyperparameter search.
+The other study is a KV quantization framework called QAQ, which uses the attention output to perform the key quantization and the self-attention output to perform the value quantization [32].
+Unlike SKVQ and all other KV quantization strategies, QAQ is the first to target the keys and values in different approaches.
+However, the difference in the approach lies in the type of data being used to perform the quantization, not in the exact number of bits being allocated to one versus the other.
+Finally, both SKVQ and QAQ are post-training quantization approaches, making them susceptible to overhead issues during the attention computation.
+Figure 2: Frobenius norm for Llama Family.
+The x-axis represents the layer index in the model while the y-axis represents the Frobenius norm magnitude.
+The norms are higher for the Key weights than for the Value weights across layers.
+3 Dynamics of the Norms of Key and Value
+3.1 Key-Value Norm Disparity
+Before formalizing our finding, a brief examination is presented on how the magnitudes of the key and value projection matrices influence KV cache quantization.
+In the self-attention mechanism, the key weights $W^{k}$ scale the token representations that enter the cache, whereas the value weights $W^{v}$ scale the representations that are read out once the attention scores are formed.
+Because quantization noise is approximately proportional to the dynamic range of the signal, the relative sizes of $\|W^{k}\|$ and $\|W^{v}\|$ determine which part of the cache is more sensitive to a given bit budget.
+Computing norms across four Llama-3 models (Fig.
+2) show that the Frobenius norm of $W^{k}$ consistently exceeds that of $W^{v}$ in almost every layer; the same ordering holds for the spectral norm.
+Intuitively, this systematic gap motivates the formal statement captured in Theorem Key-Value Norm Disparity below.
+Theorem 1 (Key-Value Norm Disparity).
+The expectation of the Frobenius norm for the key weight matrices are greater than that for the value weight matrices.
+$\mathbb{E}\!\bigl{[}\|W^{K}\|_{F}\bigr{]}\;>\;\mathbb{E}\!\bigl{[}\|W^{V}\|_{F}\bigr{]}.$ (1)
+The proof of Theorem 1 is deferred to Appendix B, but its core argument can be sketched as follows: Starting from a standard Xavier initialization[36], the analysis tracks how stochastic-gradient descent updates the Frobenius norms of $W^{k}$ and $W^{v}$.
+Under isotropic-input assumptions, the cross-term $\langle W^{m},\,\partial\mathcal{L}/\partial W^{m}\rangle_{F}$ becomes negligible, so the expected norm growth is governed by the squared gradient magnitudes.
+A careful decomposition of those gradients shows that the Key weight update contains an additional factor of the query weight matrix $Q$, whose own norm increases during training.
+Consequently, the expected squared gradient of $W^{k}$ dominates that of $W^{v}$, causing $\mathbb{E}\bigl{[}\|W^{k}\|_{F}\bigr{]}$ (and, by monotonicity, the spectral norm) to outpace its value counterpart.
+This asymptotic divergence establishes the persistent norm gap asserted Key-Value Norm Disparity.
+Although Fig.
+2 illustrates this disparity for the Llama family, supplementary plots for Mistral models in Appendix D confirm that the same pattern holds across diverse architectures [8].
+The next subsection investigates how this norm disparity influences quantization error, laying the groundwork for data-driven bit-allocation strategies.
+3.2 Key-Driven Quantization
+Theorem Key-Value Norm Disparity establishes a persistent norm gap between the Key and Value weight matrices.
+Since quantization noise scales with the dynamic range of the signal, weight matrices of larger norm inevitably introduce larger absolute errors under identical bit precision.
+Consequently, assigning equal quantization budgets to the Key and Value caches is sub-optimal.
+Consider an additive-residual Transformer block (Layer Norm omitted for clarity [37]),
+$h_{l+1}=h_{l}+W_{l}h_{l},$
+and let $\Delta_{l}$ denote the quantization perturbation applied to $W_{l}$. A single block incurs
+$\|h_{l+1}-\tilde{h}_{l+1}\|_{2}=\|\Delta_{l}h_{l}\|_{2}\leq\|\Delta_{l}\|_{2}\,\|h_{l}\|_{2}.$
+After $L$ layers the worst-case deviation satisfies
+$\|h_{L}-\tilde{h}_{L}\|_{2}\;\leq\;\Bigl{(}\prod_{i=1}^{L}\|W_{i}\|_{2}\Bigr{)}\;\max_{1\leq i\leq L}\|\Delta_{i}\|_{2}.$
+From  Key-Value Norm Disparity:
+$\mathbb{E}\!\bigl{[}\|W^{K}\|_{F}^{2}\bigr{]}>\mathbb{E}\!\bigl{[}\|W^{V}\|_{F}^{2}\bigr{]}.$ (i)
+Hence any distortion injected on the key path is amplified more strongly throughout the layers.
+Let $X\in\mathbb{R}^{\mathrm{seq}\times d_{\mathrm{model}}}$ be the hidden-state matrix at a given layer. The key and value caches are produced by the same input but different projections:
+$K=X\,W^{K},\qquad V=X\,W^{V}.$
+Because $X$ is common to both paths, multiplying (i) on the left by $X$ and using the sub-multiplicativity of the Frobenius norm, $\|AB\|_{F}\leq\|A\|_{F}\,\|B\|_{2}$, yields
+$\mathbb{E}\!\bigl{[}\|K\|_{F}^{2}\bigr{]}>\mathbb{E}\!\bigl{[}\|V\|_{F}^{2}\bigr{]},$ (ii)
+i.e. the norm gap persists in the caches themselves.
+Quantization error vs. Frobenius norm.
+For a matrix $M\in\mathbb{R}^{m\times n}$ the squared Frobenius norm $\|M\|_{F}^{2}$ equals the total energy $\sum_{i,j}M_{ij}^{2}$, or equivalently the sum of eigenvalues of $M^{\top}M$ [^1] Because $\operatorname{tr}(AB)=\operatorname{tr}(BA)$, one has $\operatorname{tr}\!\bigl{(}M^{\top}M\bigr{)}=\operatorname{tr}\!\bigl{(}MM^{\top}\bigr{)}$.
+The non-zero eigenvalues of $M^{\top}M$ and $MM^{\top}$ are identical and equal to the squared singular values $\sigma_{i}^{2}$ of $M$; hence $\|M\|_{F}^{2}=\sum_{i}\sigma_{i}^{2}=\operatorname{tr}\!\bigl{(}M^{\top}M\bigr{)}$..
+When $M$ is quantized to $\tilde{M}$ with $b$ bits, uniform-scalar quantizers yield:
+[…]
+where [^2] Big-$\Theta$ notation: $f(n)=\Theta(g(n))$ iff $\exists\,c_{1},c_{2}>0,\;n_{0}$ such that $c_{1}g(n)\leq f(n)\leq c_{2}g(n)$ for all $n\geq n_{0}$.
+the hidden constants depend only on the quantizer variant.
+Since Key and Value caches share the same shape, minimizing mean square error reduces to minimizing $\|\,\cdot\,\|_{F}^{2}$.
+When $\|K\|_{F}\gg\|V\|_{F}$ and equal bit-widths are used,
+$\mathbb{E}\!\bigl{[}\|K-\tilde{K}\|_{F}^{2}\bigr{]}\;\gg\;\mathbb{E}\!\bigl{[}\|V-\tilde{V}\|_{F}^{2}\bigr{]},$
+making the key cache the dominant source of degradation, formally:
+Theorem 2 (Key-Driven Quantization).
+Let $(b_{K},b_{V})$ denote the bit allocations assigned to the key and value caches, respectively, under a symmetric uniform-scalar quantizer.
+For any pair $(b_{K},b_{V})$ and $b_{K}>b_{V}$, the expected inference accuracy is strictly higher than for the swapped allocation $(b_{V},b_{K})$, provided that $\mathbb{E}\!\bigl{[}\|K\|_{F}^{2}\bigr{]}>\mathbb{E}\!\bigl{[}\|V\|_{F}^{2}\bigr{]}$.
+Figure 3: Singular value distribution of key and value activations for Llama 3.3-70B on C4.
 The x-axis represents the singular value indices, ordered from the 5th largest to the smallest, while the y-axis denotes the magnitude of the singular values.
@@
 This zoomed-in view reveals that, beyond the largest singular value (spectral norm), key matrices exhibit consistently higher singular values even in the lower ranges, underscoring their greater representational significance compared to value matrices.
-Consequently, key matrices should be quantized with less aggressive compression.
-See Figure 7 in Appendix C for the complete range.
-Larger Frobenius Norms Require Higher Precision.
-As we discussed in  2.1 empirical measurements across multiple transformer models reveal that key matrices have significantly larger Frobenius norms than value matrices.
-This disparity suggests that if key matrices are quantized too aggressively (e.g., with fewer bits), the MSE incurred could be disproportionately high.
-Formally, for two matrices $K$ (key) and $V$ (value), if
-$\|K\|_{F}\;\gg\;\|V\|_{F},$
-then allocating the same limited bit precision to both $K$ and $V$ would cause
-$\|K-\widetilde{K}\|_{F}^{2}\;\gg\;\|V-\widetilde{V}\|_{F}^{2},$
-since the error term for $K$ is amplified by its larger energy.
-Furthermore, errors introduced in earlier layers propagate through subsequent layers in a multi-layer transformer architecture, compounding the negative impact of coarse quantization.
-3 Method
-3.1 Adaptive KV Cache Quantization (KV-AdaQuant)
-Building on the theoretical findings that matrices with higher spectral norms propagate quantization error more strongly and that larger Frobenius norms incur greater quantization error, along with empirical evidence indicating that key matrices exhibit higher norms than value matrices, a straightforward strategy emerges: quantize keys with higher precision and values with fewer bits.
-In particular, allocating 4-bit precision to keys and 2-bit precision to values effectively controls the overall mean squared error (Figure 4), reducing memory usage while mitigating error propagation through successive transformer layers.
-(a) Mean4squared quantization error across layers for 2-bit quantization.
-(b) Mean squared quantization error across layers for 4-bit quantization.
-Figure 4: Quantization error (mean squared error) of key and value cache for the Llama 3.3 70B model on the C4 dataset across all transformer layers.
-The x-axis represents the layer index, and the y-axis indicates the mean squared quantization error.
-Key matrices exhibit consistently higher quantization error compared to value matrices, with larger discrepancies observed in lower-bit quantization.
-3.2 Differences between KV-AdaQuant and Other Methods
-The Hugging Face Transformers library implements a KV cache quantization technique that applies uniform precision to KV cache.
-Although straightforward, this approach can be suboptimal due to the distinct sensitivities of keys and values to quantization errors.
-Consequently, assigning the same bit-width to both key and value often introduces notable performance degradation.
-Furthermore, increasing the bit-width for value matrices offers minimal or no additional benefits, indicating that uniform quantization is not an ideal choice.
-Other methods, such as SKVQ Duanmu et al.
-(2024) and QAQ Dong et al.
-(2024), also propose low-bit quantization for the KV cache, leveraging mixed-precision settings.
-SKVQ, for example, quantizes the key cache to 2 bits and the value cache to 1.5 bits, demonstrating minimal accuracy loss; however, it relies on fixed hyperparameter configurations that do not deeply analyze the key and value cache’ internal characteristics.
-QAQ similarly introduces quality-adaptive quantization to highlight the heightened sensitivity of key matrices to quantization errors, yet does not thoroughly investigate the underlying matrix properties, making generalization across diverse model architectures and tasks less certain.
-By contrast, the proposed approach offers a more general framework of cache quantization that explicitly accounts for the distinct characteristics of key and value cache.
-This strategy reduces error propagation across transformer layers while achieving considerable memory savings.
-Additionally, it remains adaptable to further enhancements, such as calibration, grouping, and clustering, allowing the integration of complementary techniques to improve KV cache compression.
-This adaptability renders the framework orthogonal to existing methods, providing a more comprehensive solution to the challenges associated with KV cache quantization.
-4 Experiments
-A series of experiments is designed to investigate singular value patterns in KV cache and to demonstrate how these patterns can be leveraged to enhance practical quantization strategies.
-The experimental testbed consists of a diverse set of transformer models.
-Two distinct data setups are employed: one for analyzing the singular value distributions and another for evaluating the effectiveness of the proposed quantization approach.
-4.1 Experiment Settings
-Singular Value and Norm Analysis
-This section examines structural differences between key and value cache by analyzing their singular values and norms.
-The process begins by feeding random batches (10 samples per batch) from the C4, MMLU, and GSM8K datasets into the model for each query, where every sample is padded to match the longest sequence to facilitate observations across diverse inputs.
-For each newly generated token (up to 1,000 tokens), the KV cache from each attention mechanism is collected, yielding matrices $\mathbf{K},\mathbf{V}\in\mathbb{R}^{d_{\text{head}}\times\text{seq\_len}}$ for each head and layer.
-Singular value decomposition is then performed offline in float32 precision using the PyTorch torch.linalg.svd operator, and the resulting singular values, along with the spectral norms ($\|\cdot\|_{2}$) and Frobenius norms ($\|\cdot\|_{F}$), are computed and averaged across all heads and layers to capture global patterns.
-This approach reveals how $\|\mathbf{K}\|_{2}$ and $\|\mathbf{V}\|_{2}$ evolve in deeper layers, sheds light on the total “energy” in these matrices, and clarifies the distribution of larger and smaller singular values, offering insights into their susceptibility to quantization errors.
-Quantization Error
-Hardware typically supports 2-bit, 4-bit, and 8-bit quantization, but in theory, any bit-width can balance compression and accuracy.
-We measure quantization error across bit-widths from 2 to 8 by quantizing and dequantizing key-value matrices, finding that MSE (and thus the Frobenius norm) behaves similarly across all bit-widths.
-As shown in Figure 5, the key matrix consistently exhibits higher error.
-Bit-Width Configurations.
-The effect of different bit-widths used to represent each entry in $\mathbf{K}$ and $\mathbf{V}$ is quantified.
-Uniform settings include bit-widths ranging from 2 to 8 bits for both keys and values (e.g., $(2,2)$, $(4,4)$, or $(8,8)$).
-Mixed-precision settings compare $(2,4)$ and $(4,2)$, assigning 2 bits to either values or keys and 4 bits to the other.
-Figure 5: Mean squared error (MSE) of KV cache quantization for the Llama 3.3 70B model on the C4 dataset, using quantization bit-widths ranging from 2 to 8.
-The x-axis represents the quantization bit-width, while the y-axis shows the MSE on a logarithmic scale.
-A logarithmic scale is used to highlight differences at higher bit-widths, where MSE values decrease significantly and approach zero, particularly at 8-bit.
-Solid lines indicate the mean MSE across layers, while the shaded regions represent the min-max range of errors.
-Quantization Procedure.
-Hugging Face’s Optimum library is used for low-level CUDA kernels.
-No additional outlier or chunk-based grouping techniques are applied, so as to isolation of the effect of $\mathbf{K}$-vs-$\mathbf{V}$ bit allocation.
-Error Metric: MSE.
-To measure the quantization distortion, we record
-$\text{MSE}(\mathbf{K},\widetilde{\mathbf{K}})\;=\;\frac{1}{d_{\text{head}}\,\times\,\text{seq\_len}}\,\|\mathbf{K}-\widetilde{\mathbf{K}}\|_{F}^{2},$
-and likewise for $\mathbf{V}$.
-In addition to per layer and per head comparison, we average MSE across all heads, layers, and tokens in a batch.
-This approach yields a global view of how severely each bit-width compresses $\mathbf{K}$ vs.
-$\mathbf{V}$.
-Experimental Protocol.
-For each model and dataset subset (C4, MMLU, GSM8K):
-  1. 1.
-    Run a forward pass to produce unquantized KV caches in bfloat16.
-  2. 2.
-    Quantize to a chosen bit-width, then immediately dequantize to bfloat16.
-  3. 3.
-    Compare the dequantized version $\widetilde{\mathbf{K}}/\widetilde{\mathbf{V}}$ to the original $\mathbf{K}/\mathbf{V}$ in MSE terms.
-  4. 4.
-    Repeat for all bit-width settings of interest.
-Table 1: Quantization error (mean ± std) for key and value cache at 2-bit, 3-bit, and 4-bit quantization ($K_{i}$ and $V_{i}$) across different datasets.
+For the complete range, see Figure 5 in Appendix E.
+Appendix 3 and  4 establishes a norm-dependent upper bound on the quantization error of an arbitrary matrix $M$, and Appendix C.6 applies this bound to show that key caches incur proportionally larger errors than value caches.
+To empirically validate the theoretical predictions, we examined the singular value spectra of the K anc V caches across a broad set of models, up to 70B, by extracting every attention head in every layer (Section 4.1).
+Figure 3 displays the singular values from the fifth largest onward, revealing a systematic separation between keys and values; the complete spectra in Appendix E (Figure 5) show the same trend across all modes.
+Aggregate norms reinforce this observation: Figures 7 and 7 plot the layer-wise Frobenius and spectral norms of the $K\!V$ cache in Llama-3.3-70B, consistently indicating larger magnitudes for keys than for values.
+4 Experimental Results
+4.1 Experiment Setup
+We conduct two experiments: (i) a quantization-error analysis and (ii) a downstream evaluation.
+For the quantization-error analysis, we use PyTorch’s weight-packing quantization to measure reconstruction error.
+For the downstream evaluation, we employ hardware-aligned quantization via Hugging Face’s Optimum Quanto library as back-end (token-wise, 2/4-bit quantization) on Llama-3.2-1B, Llama-3.1-8B [38], Phi-4-14B [39] and DeepSeek-R1-Qwen-14B, and the Half-Quadratic Quantization (HQQ) [40] back-end (channel-wise, 1–8-bit quantization) on Llama-3.1-8B, Llama-3.2-1B, Llama-3.2-3B, and Qwen3-0.6B/1.7B/4B/8B [41].
+The quantization-error sweep additionally includes Phi-3-Medium-128K [42], Llama-3.1-Nemotron-70B [43], Mistral-0.3-7B [8], Qwen-2.5-14B and Llama-3.3-70B.
+See Appendix G for full details on compute resources, software tools, and reproducibility.
+Table 1: Quantization errors for KV cache under various configurations.
+The quantization error is expressed as mean ± std.
+We use 2-bit, 3-bit, and 4-bit quantization for both $K_{i}$ and $V_{i}$ across three datasets.
 Dataset Model 2-bit 3-bit 4-bit
@@
 DeepSeekR1Q-14B 4.477 ± 2.176 1.424 ± 1.752 0.830 ± 0.256 0.200 ± 0.242 0.181 ± 0.058 0.044 ± 0.056
-Downstream Evaluation.
-While MSE is a reasonable proxy for overall distortion, the ultimate question is whether the quantized KV cache preserves model quality on realistic tasks.
-We therefore evaluate the final model performance on GSM8K in two prompting modes: 1-shot and 8-shot prompting using 4 quantization settings:
-  • $\mathbf{K}_{\mathbf{(2bit)}}\mathbf{V}_{\mathbf{(2bit)}}$: Aggressive uniform quantization.
-  • $\mathbf{K}_{\mathbf{(4bit)}}\mathbf{V}_{\mathbf{(4bit)}}$: More conservative uniform quantization.
-  • $\mathbf{K}_{\mathbf{(2bit)}}\mathbf{V}_{\mathbf{(4bit)}}$ vs.
-$\mathbf{K}_{\mathbf{(4bit)}}\mathbf{V}_{\mathbf{(2bit)}}$: Mixed-precision assignments, giving 2 bits to either $\mathbf{K}$ or $\mathbf{V}$ and 4 bits to the other.
-4.2 Experiment Results
-Table 2: GSM8K 1-shot and 8-shot performance of various models under different key-value quantization settings.
-$\mathbf{K}_{i}\mathbf{V}_{j}$ denotes $i$-bit quantization for the K cache and $j$-bit for the V cache.
-The $\mathbf{K}_{\mathbf{(4bit)}}\mathbf{V}_{\mathbf{(2bit)}}$ column, highlighted below, represents our main finding.
+4.2 Quantization Error
+In several studies (Table 3 of KIVI[29], Table 8 of GEAR[44], and Table 4 of RotateKV[45]), quantized KV cache representations have been shown to outperform their BF16 full-precision counterparts, an outcome that may seem counterintuitive.
+Taking BF16 as the gold standard, we attribute this phenomenon to the implicit regularization effect introduced by quantization: by discretizing activation values, quantization injects a small, structured noise into the model’s intermediate representations.
+Much like dropout or weight decay, this noise can smooth overly sharp decision boundaries and mitigate over-fitting, thereby improving generalization on held-out tasks.
+To isolate and measure this effect, our first experiment focuses on characterizing the quantization error in the KV cache, without grouping tokens, or using residual tokens in each quantization turn.
+In addition to per layer and per head comparison, we average MSE across all heads, layers, and tokens in a batch.
+This approach yields a global view of how severely each bit-width compresses $\mathbf{K}$ vs.
+$\mathbf{V}$.
+The evaluation uses ten random sequences from C4 [46], MMLU [47], and GSM8K [48], padded to the longest length and generated autoregressively for up to 1 000 tokens.
+Table 1 presents 2-, 3-, and 4-bit errors for seven representative models (Llama3, Phi4 [39], Mistral, Qwen, DeepSeek-Distilled, and others); the complete results appear in Appendix F.
+Figure 8 plots the full error curve for Llama-3.3-70B on C4, while Figure 10 shows the per-layer breakdown for Llama-3.1-8B.
+4.3 Evaluation
+Experiments in Section 4.2 established that K cache incur larger MSE than V cache at every bit-width, suggesting that $\mathbf{K}$ contains proportionally richer information.
+However, ultimately model quality is measured by downstream metrics such as accuracy rather than reconstruction error.
+Our second experiments therefore measures how the observed error asymmetry translates into task performance under realistic, deployment-oriented constraints.
+KV cache quantization methods rarely de-quantize the entire cache after every generated token because the associated memory traffic defeats the purpose of quantization.
+Instead, they maintain a residual buffer that holds the most recent tokens in full-precision and flushes it en bloc once full.
+We replicate this strategy by (i) reserving a fixed-size residual cache of 64 tokens and (ii) grouping activations into 128-element blocks—settings adopted by popular open-source implementations such as KIVI, Flash-Decoding [49] and Marlin [50].
+Two hardware-aligned quantization kernels were selected to span the design space:
+  • Quanto: token-wise (per-row) quantization applied to the key and value caches, supporting $\{2,4\}$-bit precision.
+  • HQQ: channel-wise (per-column) quantization applied to the key and value caches, supporting $\{1,2,4,6,8\}$-bit precision.
+For downstream evaluation, we employ three generative datasets that require the model to produce extended outputs rather than select among fixed choices: the GSM8K, the CoQA [51] conversational question answering dataset, and the EQ-Bench [52] long-form structured QA suite.
+Results.
+Tables 2 and 3 corroborate the norm-based hypothesis by showing that downstream accuracy is far more sensitive to the precision assigned to the key cache than to the value cache.
+  • Quanto backend.
+Across four models and two prompt lengths, allocating $4$ bits to keys and $2$ bits to values ($\mathbf{K}_{4}\mathbf{V}_{2}$) recovers on average $94\%$ of the full-precision score, whereas the inverted scheme ($\mathbf{K}_{2}\mathbf{V}_{4}$) loses up to $30$ percentage points (pp).
+The margin widens with model capacity: in 1-shot GSM8K, $\mathbf{K}_{4}\mathbf{V}_{2}$ outperforms $\mathbf{K}_{2}\mathbf{V}_{4}$ by $+0.30$ pp on Llama-3.2-1B and by $+0.16$ pp on Phi-4-14B.
+Notably, $\mathbf{K}_{4}\mathbf{V}_{2}$ attains almost the same accuracy as the symmetric $4$-bit baseline.
+  • HQQ backend.
+The broader sweep over $\{1,2,4,6,8\}$ bits and three datasets paints the same picture.
+On GSM8K, granting even a single extra bit to keys yields double-digit gains; e.g., Qwen3-8B improves from $0.08$ to $0.54$ accuracy ($+46$ pp) when the extra bit is moved from values to keys.
+Gaps persist on EQ-Bench and CoQA, where key-centric quantization delivers up to $+62$ pp (Qwen3-0.6B, EQ-Parseable, $1$-bit) and never underperforms the converse setting.
+Advantages taper but do not vanish at higher precisions: at $6$ bits the key-first policy still adds $7\!-\!12$ pp for 8–14B models.
+Table 2: KV cache quantization on GSM8K 1-shot and 8-shot performance employing Optimum Quanto backend.
+$\mathbf{K}_{i}\mathbf{V}_{j}$ denotes $i$-bit quantization for the K cache and $j$-bit for the V cache using Quanto backend.
+The $\mathbf{K}_{\mathbf{(4bit)}}\mathbf{V}_{\mathbf{(2bit)}}$ column, highlighted below, shows the optimal bit allocation setting.
 1-shot 8-shot
@@
 DeepSeekR1Q-14B 0.772 0.775 0.865 0.867 0.763 0.792 0.876 0.875
-Table 1 shows that across datasets and models, quantization error for the key matrices is consistently and substantially higher than that for the value matrices.
-For instance, in the MMLU dataset, Llama3.2-1B exhibits an error of $4.85\pm 1.04$ for $\mathbf{K}_{2}$ versus only $0.13\pm 0.10$ for $\mathbf{V}_{2}$.
-This trend persists for both the 3-bit and 4-bit settings, where increasing the bit-width reduces the error for both matrices, yet the keys still incur a notably higher error.
-Similar patterns are evident in the C4 and GSM8K datasets, highlighting that key matrices are inherently more sensitive to quantization noise.
-Furthermore, Figure 6 in the Appendix B illustrates the quantization error per attention head in each layer for the Llama3.1-8B model on C4, MMLU, and GSM8K, reinforcing our observation (see Appendix D for extended results).
-Table 2 shows accuracy on GSM8K under 1-shot and 8-shot prompting.
-Uniform K${}_{(2\text{bit})}$${}_{(2\text{bit})}$ settings degrade accuracy noticeably, especially in the longer 8-shot context.
-By contrast, K${}_{(4\text{bit})}$${}_{(2\text{bit})}$ nearly matches the unquantized baseline, confirming that providing higher precision to keys preserves model quality while still compressing values aggressively.
-The reversed assignment K${}_{(2\text{bit})}$${}_{(4\text{bit})}$ consistently underperforms, underscoring that allocating fewer bits to key matrices adversely affects accuracy.
-5 Related Work
-Model Quantization.
-for large models such as LLMs Wan et al. (2024), which is why post-training quantization (PTQ) techniques are generally preferred for quantizing these models.
-Quantization reduces the computational cost of neural network inference by lowering model bit-precision Han et al.
-(2016); Nagel et al.
-(2021).
-It can be divided into Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT).
-PTQ requires no re-training or labeled data Jacob et al.
-(2017); Nagel et al.
-(2019) and is preferred for large models like LLMs, whereas QAT demands fine-tuning with labeled data Nagel et al.
-(2021); Gholami et al.
-(2022) but scales poorly for such models Wan et al.
-(2024).
-LLMs Quantization.
-Researchers consider three post-training quantization settings for LLMs: weight-only, weight-activation, and KV cache quantization.
-Weight-only quantization focuses solely on weights Frantar et al.
-(2023a, b); Lin et al.
-(2024), while weight-activation quantization targets both weights and activations Dettmers et al.
-(2022); Xiao et al.
-(2023); Shao et al.
-(2024).
-Since longer inputs inflate the KV cache’s memory overhead Pope et al.
-(2023), KV cache quantization offers memory savings similar to weight-activation methods while maintaining performance near that of weight-only quantization Yue et al.
-(2024).
-KV Cache Quantization.
-Recent KV cache quantization methods reduce LLM memory and accelerate inference, falling into three categories: outlier redistribution, fixed-precision, and mixed-precision Li et al.
-(2025).
-Outliers in KV cache are addressed by redistributing them or applying smoothing transformations Xiao et al.
-(2023); Lin et al.
-(2024); Shao et al.
-(2024).
-Fixed-precision methods use a uniform bit-width Yao et al.
-(2022); Sheng et al.
-(2023); Liu et al.
-(2024) but can overlook token importance and outliers.
-In contrast, mixed-precision allocates higher precision to critical tokens and lower precision elsewhere Hooper et al.
-(2024); Yue et al.
-(2024); Li et al.
-(2024); Dong et al.
-(2024); Duanmu et al.
-(2024).
-Existing methods like SKVQ and QAQ rely on specific engineering solutions or predefined parameters, whereas KV-AdaQuant uses a principled, analysis-driven framework that leverages the intrinsic properties of key and value caches for a more robust and adaptable quantization approach.
+22footnotetext: Although Chain-of-Thought prompting is generally discouraged for reasoning models and can even impair their performance, we deliberately apply 1-shot and 5-shot CoT prompts here to examine the effect of longer input prompts on KV cache quantization.
+Table 3: KV cache quantization comparison across GSM8K, EQ-Parseable, and CoQA datasets.
+The results are obtained on the HQQ quantization backend.
+Each cell shows accuracies for $\mathbf{K}_{i}\mathbf{V}_{x}$ and $\mathbf{K}_{x}\mathbf{V}_{i}$, where $x$ denotes the mean accuracy across quantization bit-widths $\{1,2,4,6,8\}$, with the percentage-point gap annotated above the comparison sign.
+See Appendix H for all settings.
+GSM8K $\mathbf{K}_{1}\mathbf{V}_{x}$ vs.
+$\mathbf{K}_{x}\mathbf{V}_{1}$ $\mathbf{K}_{2}\mathbf{V}_{x}$ vs.
+$\mathbf{K}_{x}\mathbf{V}_{2}$ $\mathbf{K}_{4}\mathbf{V}_{x}$ vs.
+$\mathbf{K}_{x}\mathbf{V}_{4}$ $\mathbf{K}_{6}\mathbf{V}_{x}$ vs.
+$\mathbf{K}_{x}\mathbf{V}_{6}$ $\mathbf{K}_{8}\mathbf{V}_{x}$ vs.
+$\mathbf{K}_{x}\mathbf{V}_{8}$
+Qwen3-0.6B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+3\%}}{{<}}\mathbf{0.03}$ $0.00\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{<}}\mathbf{0.16}$ $0.04\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{<}}\mathbf{0.17}$ $\mathbf{0.34}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.18$ $\mathbf{0.34}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.18$
+Llama-3.2-1B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+5\%}}{{<}}\mathbf{0.05}$ $0.01\stackrel{{\scriptstyle\scriptscriptstyle+18\%}}{{<}}\mathbf{0.19}$ $\mathbf{0.27}\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{>}}0.20$ $\mathbf{0.28}\stackrel{{\scriptstyle\scriptscriptstyle+8\%}}{{>}}0.20$ $\mathbf{0.28}\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{>}}0.21$
+Llama-3.2-3B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+19\%}}{{<}}\mathbf{0.19}$ $0.27\stackrel{{\scriptstyle\scriptscriptstyle+17\%}}{{<}}\mathbf{0.44}$ $\mathbf{0.57}\stackrel{{\scriptstyle\scriptscriptstyle+11\%}}{{>}}0.46$ $\mathbf{0.58}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.46$ $\mathbf{0.59}\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{>}}0.46$
+Qwen3-4B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+41\%}}{{<}}\mathbf{0.41}$ $0.01\stackrel{{\scriptstyle\scriptscriptstyle+49\%}}{{<}}\mathbf{0.50}$ $\mathbf{0.80}\stackrel{{\scriptstyle\scriptscriptstyle+29\%}}{{>}}0.51$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.51$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.51$
+Llama-3.1-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{<}}\mathbf{0.31}$ $0.35\stackrel{{\scriptstyle\scriptscriptstyle+17\%}}{{<}}\mathbf{0.52}$ $\mathbf{0.70}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.54$ $\mathbf{0.70}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.54$ $\mathbf{0.70}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.54$
+Qwen3-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+48\%}}{{<}}\mathbf{0.48}$ $0.08\stackrel{{\scriptstyle\scriptscriptstyle+46\%}}{{<}}\mathbf{0.54}$ $\mathbf{0.86}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.55$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+32\%}}{{>}}0.55$ $\mathbf{0.86}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.55$
+Qwen3-14B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+50\%}}{{<}}\mathbf{0.50}$ $0.26\stackrel{{\scriptstyle\scriptscriptstyle+32\%}}{{<}}\mathbf{0.58}$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}0.60$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}0.60$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}0.60$
+Qwen3-32B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+42\%}}{{<}}\mathbf{0.42}$ $0.25\stackrel{{\scriptstyle\scriptscriptstyle+23\%}}{{<}}\mathbf{0.48}$ $\mathbf{0.73}\stackrel{{\scriptstyle\scriptscriptstyle+23\%}}{{>}}0.50$ $\mathbf{0.71}\stackrel{{\scriptstyle\scriptscriptstyle+21\%}}{{>}}0.50$ $\mathbf{0.72}\stackrel{{\scriptstyle\scriptscriptstyle+21\%}}{{>}}0.51$
+EQ-Parseable
+Qwen3-0.6B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+10\%}}{{<}}\mathbf{10.06}$ $0.00\stackrel{{\scriptstyle\scriptscriptstyle+53\%}}{{<}}\mathbf{52.51}$ $51.23\stackrel{{\scriptstyle\scriptscriptstyle+2\%}}{{<}}\mathbf{52.87}$ $\mathbf{84.21}\stackrel{{\scriptstyle\scriptscriptstyle+32\%}}{{>}}52.63$ $\mathbf{84.91}\stackrel{{\scriptstyle\scriptscriptstyle+33\%}}{{>}}52.28$
+Llama-3.2-1B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+34\%}}{{<}}\mathbf{33.68}$ $25.61\stackrel{{\scriptstyle\scriptscriptstyle+37\%}}{{<}}\mathbf{61.52}$ $\mathbf{87.25}\stackrel{{\scriptstyle\scriptscriptstyle+22\%}}{{>}}65.38$ $\mathbf{90.06}\stackrel{{\scriptstyle\scriptscriptstyle+24\%}}{{>}}66.32$ $\mathbf{90.06}\stackrel{{\scriptstyle\scriptscriptstyle+25\%}}{{>}}66.08$
+Llama-3.2-3B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+36\%}}{{<}}\mathbf{35.44}$ $68.19\stackrel{{\scriptstyle\scriptscriptstyle+5\%}}{{<}}\mathbf{73.10}$ $\mathbf{90.06}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}76.49$ $\mathbf{89.12}\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{>}}76.02$ $\mathbf{89.94}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}76.26$
+Qwen3-4B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+54\%}}{{<}}\mathbf{47.60}$ $29.01\stackrel{{\scriptstyle\scriptscriptstyle+33\%}}{{<}}\mathbf{58.01}$ $\mathbf{84.09}\stackrel{{\scriptstyle\scriptscriptstyle+28\%}}{{>}}59.88$ $\mathbf{85.96}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}59.06$ $\mathbf{85.38}\stackrel{{\scriptstyle\scriptscriptstyle+29\%}}{{>}}59.88$
+Llama-3.1-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+61\%}}{{<}}\mathbf{60.58}$ $\mathbf{77.89}\stackrel{{\scriptstyle\scriptscriptstyle+2\%}}{{>}}75.79$ $\mathbf{95.67}\stackrel{{\scriptstyle\scriptscriptstyle+19\%}}{{>}}77.19$ $\mathbf{96.96}\stackrel{{\scriptstyle\scriptscriptstyle+20\%}}{{>}}77.31$ $\mathbf{97.43}\stackrel{{\scriptstyle\scriptscriptstyle+21\%}}{{>}}77.08$
+Qwen3-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+62\%}}{{<}}\mathbf{58.48}$ $54.39\stackrel{{\scriptstyle\scriptscriptstyle+17\%}}{{<}}\mathbf{70.41}$ $\mathbf{95.32}\stackrel{{\scriptstyle\scriptscriptstyle+25\%}}{{>}}71.35$ $\mathbf{96.26}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}70.76$ $\mathbf{96.14}\stackrel{{\scriptstyle\scriptscriptstyle+26\%}}{{>}}71.11$
+Qwen3-32B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+60\%}}{{<}}\mathbf{47.37}$ $56.14\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{<}}\mathbf{61.75}$ $\mathbf{80.12}\stackrel{{\scriptstyle\scriptscriptstyle+22\%}}{{>}}62.46$ $\mathbf{80.47}\stackrel{{\scriptstyle\scriptscriptstyle+23\%}}{{>}}62.34$ $\mathbf{79.77}\stackrel{{\scriptstyle\scriptscriptstyle+22\%}}{{>}}62.57$
+CoQA
+Qwen3-0.6B $0.20\stackrel{{\scriptstyle\scriptscriptstyle+33\%}}{{<}}\mathbf{0.44}$ $0.32\stackrel{{\scriptstyle\scriptscriptstyle+29\%}}{{<}}\mathbf{0.52}$ $\mathbf{0.65}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.53$ $\mathbf{0.69}\stackrel{{\scriptstyle\scriptscriptstyle+23\%}}{{>}}0.53$ $\mathbf{0.69}\stackrel{{\scriptstyle\scriptscriptstyle+23\%}}{{>}}0.53$
+Llama-3.2-1B $0.22\stackrel{{\scriptstyle\scriptscriptstyle+29\%}}{{<}}\mathbf{0.42}$ $0.48\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{<}}\mathbf{0.57}$ $\mathbf{0.67}\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{>}}0.58$ $\mathbf{0.68}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.58$ $\mathbf{0.68}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.58$
+Llama-3.2-3B $0.21\stackrel{{\scriptstyle\scriptscriptstyle+50\%}}{{<}}\mathbf{0.60}$ $\mathbf{0.73}\stackrel{{\scriptstyle\scriptscriptstyle+6\%}}{{>}}0.68$ $\mathbf{0.79}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.68$ $\mathbf{0.79}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.68$ $\mathbf{0.79}\stackrel{{\scriptstyle\scriptscriptstyle+15\%}}{{>}}0.67$
+Qwen3-4B $0.34\stackrel{{\scriptstyle\scriptscriptstyle+34\%}}{{<}}\mathbf{0.62}$ $0.68\stackrel{{\scriptstyle\scriptscriptstyle+3\%}}{{<}}\mathbf{0.70}$ $\mathbf{0.81}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.71$ $\mathbf{0.81}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.71$ $\mathbf{0.81}\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{>}}0.70$
+Llama-3.1-8B $0.21\stackrel{{\scriptstyle\scriptscriptstyle+50\%}}{{<}}\mathbf{0.60}$ $\mathbf{0.72}\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{>}}0.66$ $\mathbf{0.78}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.67$ $\mathbf{0.78}\stackrel{{\scriptstyle\scriptscriptstyle+15\%}}{{>}}0.67$ $\mathbf{0.78}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.67$
+Qwen3-8B $0.38\stackrel{{\scriptstyle\scriptscriptstyle+37\%}}{{<}}\mathbf{0.68}$ $\mathbf{0.73}\stackrel{{\scriptstyle\scriptscriptstyle+1\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$
+Qwen3-14B $0.32\stackrel{{\scriptstyle\scriptscriptstyle+45\%}}{{<}}\mathbf{0.69}$ $\mathbf{0.79}\stackrel{{\scriptstyle\scriptscriptstyle+8\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$
+Qwen3-32B $0.32\stackrel{{\scriptstyle\scriptscriptstyle+44\%}}{{<}}\mathbf{0.68}$ $\mathbf{0.79}\stackrel{{\scriptstyle\scriptscriptstyle+9\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.73$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.73$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.73$
 5 Conclusion
-In this work, we systematically analyze KV cache matrices from the perspectives of linear algebra and information theory to uncover distinct patterns in the information carried by these matrices.
-Building on the singular value distribution of the KV cache, we propose an information-aware quantization framework for KV cache compression.
-We evaluate the effectiveness of our approach on two widely used datasets, GSM8K, demonstrating its practical impact.
-While our current study focuses on the per-token quantization methods, the proposed framework is general and can be applied to any key-value quantization technique.
-Limitations
-Although our information-aware quantization framework is effective, there are aspects that warrant further consideration.
-First, our experiments utilize a maximum context length of 2K tokens.
-While this setting aligns with many practical inference scenarios, it may not fully capture the demands of tasks requiring larger context windows (e.g., 4K or 8K tokens).
-Extending our approach to support longer contexts could reveal additional challenges, such as increased quantization sensitivity and greater memory management overhead.
-Ethical Considerations
-Our research aims to reduce the memory footprint and computational overhead of LLM in inference time through KV cache quantization.
-Such advancements can potentially lower energy consumption and expand access to resource-constrained communities, contributing to more environmentally sustainable and inclusive deployment of language modeling based productions.
-However, the use of proposal KV cache compression method must be carefully monitored to ensure that accuracy trade-offs do not disproportionately affect critical applications such as healthcare, legal, and financial contexts.
-Appendix A Quantization and Dequantization Error Bounds
+Large scale language models increasingly spend the majority of their inference footprint on the KV cache, making cache compression a decisive factor in practical deployment.
+This work supplied a principled answer to the bit-allocation problem by unifying theory and experiment.
+  • Key–Value Norm Disparity. We proved that Key weights exhibit larger Frobenius norms than Value weights throughout training, implying a higher information density in the key path.
+  • Key-Driven Quantization.
+Leveraging this norm gap, a second theorem shows that assigning the finer bit-width to keys and the coarser bit-width to values strictly minimises quantization error, thereby preserving downstream accuracy.
+  • Empirical Validation.
+Experiments across nine model families, three generative benchmarks, and two hardware-aligned back-ends confirm the theory: a $4\!\times\!2$ (keys / values) split consistently recovers $\geq 94\%$ of full-precision accuracy while halving KV-cache memory.
+These findings elevate bit allocation from a hyper-parameter tuned by grid search to a design principle derived from model geometry, offering immediate guidance for practitioners and a foundation for hardware–algorithm co-design.
+We used a context length of 2K tokens, a common setting, but will use longer contexts in future.
+Appendix
+Appendix A Dynamics of the Norms of Key and Value Weight Matrices
+A.1 Preliminary: Dimensions
+$\begin{array}[]{@{}ll@{}}c&\text{sequence length (context size)},\\[4.0pt] d_{m}&\text{embedding dimension},\\[4.0pt] d_{v}=d_{k}=d_{m}/n_{h}=d_{m}&\text{for a single head ($n_{h}=1$)},\\[4.0pt] n_{h}&\text{number of attention heads}.\end{array}$
+A.2 Symbols
+$\begin{array}[]{@{}ll@{}}X&\text{input matrix }(c\times d_{m}),\\ X^{O}&\text{output matrix }(c\times d_{m}),\\ W^{V}&\text{value weights }(d_{m}\times d_{m}),\\ W^{K}&\text{key weights }(d_{m}\times d_{m}),\\ W^{Q}&\text{query weights }(d_{m}\times d_{m}),\\ K=XW^{K}&(c\times d_{m}),\\ Q=XW^{Q}&(c\times d_{m}),\\ V=XW^{V}&(c\times d_{m}),\\ S=QK^{\top}/\sqrt{d_{m}}&(c\times c),\\ A=\operatorname{softmax}(S)&(c\times c),\\ \mathcal{L}&\text{loss function.}\end{array}$
+Appendix B Norm Dynamics
+In this appendix, we compare the long–time behavior of the Frobenius norms of the key ($W^{K}$) and value ($W^{V}$) weight matrices of a single–head self–attention layer trained with stochastic gradient descent (SGD).
+We show that, under standard isotropic assumptions, $\|W^{K}\|_{F}$ grows faster than $\|W^{V}\|_{F}$.
+B.1 Update rule
+At update step $i$ the weights follow
+$W^{m}_{i+1}\;=\;W^{m}_{i}-\eta\,\frac{\partial\mathcal{L}}{\partial W^{m}_{i}},\qquad m\in\{K,V\},$ (2)
+with learning rate $\eta$. Squaring the Frobenius norm of both sides gives
+$\|W^{m}_{i+1}\|_{F}^{2}$ $=\|W^{m}_{i}\|_{F}^{2}+\eta^{2}\Bigl{\|}\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr{\|}_{F}^{2}-2\eta\,\Bigl{\langle}W^{m}_{i},\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr{\rangle}_{F},$ (3)
+where
+$\|A\|_{F}$ $\equiv\sqrt{\operatorname{tr}(A^{\top}A)},$ (4)
+$\langle A,B\rangle_{F}$ $\equiv\operatorname{tr}(A^{\top}B).$ (5)
+B.2 Expectation over mini-batches
+Taking an expectation over mini-batches yields
+$\Delta\mathbb{E}\bigl{[}\|W^{m}_{i}\|_{F}^{2}\bigr{]}$ $\equiv\mathbb{E}\bigl{[}\|W^{m}_{i+1}\|_{F}^{2}\bigr{]}-\mathbb{E}\bigl{[}\|W^{m}_{i}\|_{F}^{2}\bigr{]}$ (6)
+[…]
+In high–dimensional weight space the gradient is almost orthogonal to the current weights, making the second term negligible. Hence,
+$\Delta\mathbb{E}\bigl{[}\|W^{m}_{i}\|_{F}^{2}\bigr{]}\;\approx\;\eta^{2}\,\mathbb{E}\Bigl{[}\Bigl{\|}\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr{\|}_{F}^{2}\Bigr{]}.$ (8)
+B.3 Initialization
+Throughout we assume Xavier initialization: each entry of $W^{K}$ and $W^{V}$ is drawn i.i.d. from a zero–mean distribution whose variance preserves the input scale [36].
+B.4 Gradients
+The attention outputs are
+$X^{O}\;=\;A\,V\,W^{O},$ (9)
+where $A=\operatorname{softmax}\!\bigl{(}QK^{\top}/\sqrt{d_{m}}\bigr{)}$ and $V=XW^{V}$. Differentials give
+$d\mathcal{L}(X^{O})$ $=\operatorname{tr}\!\bigl{[}(\partial\mathcal{L}/\partial X^{O})^{\!\top}\,AX\,dW^{V}W^{O}\bigr{]}=\operatorname{tr}\!\bigl{[}(AX)^{\top}\,\frac{\partial\mathcal{L}}{\partial X^{O}}\,(W^{O})^{\top}\,dW^{V}\bigr{]},$ (10)
+so that
+$\frac{\partial\mathcal{L}}{\partial W^{V}}=(AX)^{\top}\,\frac{\partial\mathcal{L}}{\partial X^{O}}\,(W^{O})^{\top}.$ (11)
+Similarly, writing $S=QK^{\top}/\sqrt{d_{m}}$,
+$d\mathcal{L}(S)$ $=\operatorname{tr}\!\bigl{[}(\partial\mathcal{L}/\partial S)^{\!\top}\,dS\bigr{]}=\frac{1}{\sqrt{d_{m}}}\,\operatorname{tr}\!\bigl{[}X^{\top}\,(\partial\mathcal{L}/\partial S)^{\!\top}\,Q\,dW^{K}\bigr{]},$ (12)
+yielding
+$\frac{\partial\mathcal{L}}{\partial W^{K}}=\frac{1}{\sqrt{d_{m}}}\,X^{\top}\,(\partial\mathcal{L}/\partial S)^{\!\top}\,Q.$ (13)
+B.5 Expected gradient norms
+Squaring the Frobenius norm of (11) and taking an expectation under the isotropic-input assumption,
+$\mathbb{E}\!\bigl{[}\|\partial\mathcal{L}/\partial W^{V}\|^{2}_{F}\bigr{]}$ $=c\,\sigma_{x}^{2}\,\sigma_{o}^{2},$ (14)
+where $\mathbb{E}[X^{\top}X]=c\sigma_{x}^{2}I$ and $\mathbb{E}\Bigl{[}\Bigl{(}A^{\top}\tfrac{\partial\mathcal{L}}{\partial X^{O}}(W^{O})^{\top}\Bigr{)}\Bigl{(}A^{\top}\tfrac{\partial\mathcal{L}}{\partial X^{O}}(W^{O})^{\top}\Bigr{)}^{\top}\Bigr{]}=\sigma_{o}^{2}I$.
+For the key weights,
+$\mathbb{E}\!\bigl{[}\|\partial\mathcal{L}/\partial W^{K}\|^{2}_{F}\bigr{]}=\frac{\sigma_{x}^{2}\,\sigma_{s}^{2}}{d_{m}}\,\|Q\|_{F}^{2},$ (15)
+with $\sigma_{s}^{2}$ the entry-wise variance of $\partial\mathcal{L}/\partial S$.
+B.6 Asymptotic comparison
+Substituting these expectations into (8) shows that
+$\mathbb{E}\!\bigl{[}\|W^{K}\|^{2}_{F}\bigr{]}\text{grows as}\;\eta^{2}\,\frac{\sigma_{x}^{2}\sigma_{s}^{2}}{d_{m}}\,\|Q\|_{F}^{2},\quad\text{whereas}\quad\mathbb{E}\!\bigl{[}\|W^{V}\|^{2}_{F}\bigr{]}\;\text{grows as}\;\eta^{2}\,c\sigma_{x}^{2}\sigma_{o}^{2}.$ (16)
+Because $\|Q\|_{F}^{2}$ itself increases during training, $\|W^{K}\|_{F}$ eventually dominates:
+$\mathbb{E}\!\bigl{[}\|W^{K}\|_{F}\bigr{]}\;>\;\mathbb{E}\!\bigl{[}\|W^{V}\|_{F}\bigr{]}.$ (17)
+As shown in Equation 17, the expectation of the Frobenius norm of $W^{k}$ is greater than that of $W^{v}$.
+[^3] While this analysis assumes a single-head, decoder-only transformer for clarity, the extension to multi-head attention is direct.
+The same conclusion also holds for group-query and multi-query attention.
+Appendix C Quantization and Dequantization Error Bounds
 In this appendix we establishes upper bounds on the quantization error incurred when a matrix is represented using a finite bit-width in two’s-complement format.
@@
 Hence, matrices with larger Frobenius norms are also more vulnerable to quantization errors and would benefit from a higher precision during quantization.
-Preliminaries and Notation
+C.1 Preliminaries and Notation
 Let $A\in\mathbb{R}^{m\times n}$ denote a real matrix whose entries are to be stored using a fixed number of bits in two’s-complement representation. The following notation is adopted:
@@
 in terms of $b$, $m$, $n$, and the norms of $A$.
-Spectral Norm Bound
+C.2 Spectral Norm Bound
 Theorem 3 (Spectral Norm Quantization Error).
@@
 $\|A-\widehat{A}\|_{2}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{2}.$
-Frobenius Norm Bound
+C.3 Frobenius Norm Bound
 Theorem 4 (Frobenius Norm Quantization Error).
@@
 $\|A-\widehat{A}\|_{F}$ $\lesssim\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{F}.$
-Quantizing KV Cache
+C.4 Quantizing KV Cache
 Consider the key and value cache
@@
 where $\ast$ denotes either the spectral norm ($\|\cdot\|_{2}$) or the Frobenius norm ($\|\cdot\|_{F}$). In practice, $\mathbf{K}$ typically exhibits a larger norm than $\mathbf{V}$.
-Spectral-Norm Perspective
+C.5 Spectral-Norm Perspective
 Standard quantization error bounds yield
@@
 Since $\|\mathbf{V}\|_{2}<\|\mathbf{K}\|_{2}$, it follows that $b_{K}>b_{V}$.
-Frobenius-Norm (MSE) Perspective
+C.6 Frobenius-Norm (MSE) Perspective
 Similarly, the Frobenius norm error bounds are given by
@@
 Again, because $\|\mathbf{V}\|_{F}<\|\mathbf{K}\|_{F}$, it follows that $b_{K}>b_{V}$.
-Appendix B Quantization Error Across Layers and Heads
-Figure 6 displays the MSE of the quantized key and value caches in the Llama 3.1 8B model, evaluated layer by layer across 32 layers and differentiated by attention head.
-Results are presented for three datasets: C4, MMLU, and GSM8k.
-Each subfigure illustrates the 2-bit and 4-bit quantization configurations for the key and value caches.
-The error patterns vary across layers, suggesting that certain layers or heads may be more sensitive to aggressive quantization; however, the overall pattern indicates that assigning 4-bit to the key cache is comparable to assigning 2-bit to the value cache.
-(a) C4 dataset
-(b) MMLU dataset
-(c) GSM8k dataset
-Figure 6: Quantization error (MSE) of K and V caches in the Llama 3.1 8B model across 32 layers for the (a) C4, (b) MMLU, and (c) GSM8k datasets.
-The top row in each plot shows 2-bit quantization (k2, v2), while the bottom row shows 4-bit quantization (k4, v4).
-Each point represents a different attention head in the corresponding layer.
-The x-axis indicates the layer index, and the y-axis shows the quantization error in MSE.
-Appendix C Extended Visualization of Singular Value Distributions
+Appendix D Supplementary Key-Value Weight Norm Analysis
+Here we present additional evidence supporting the Key-Value Norm Disparity theorem established in Section 3.1.
+While Fig.
+2 demonstrates this observation for the Llama model family, the same pattern consistently emerges across Mistral family.
+Figure 4: Frobenius norm plot for Mistral Family.
+The x-axis represents the layer index in the model while the y-axis represents the Frobenius norm magnitude.
+The spectral norms are higher for the key weights than for the value weights across layers.
+Appendix E Extended Visualization of Singular Value Distributions
 Figure 5 provides the full-range view of the singular value distribution for the key and value cache at various layers of the Llama 3.3 70B model on the C4 dataset.
@@
 The shaded region captures the minimum-to-maximum range of singular values across the attention heads within each layer, while the solid curves depict the mean singular value at each rank.
-This figure underscores two main observations.
-First, key matrices systematically exhibit higher leading singular values than value matrices.
-In particular, the first singular value (the spectral norm) for key matrices is consistently larger, confirming that the key matrices possess a higher spectral norm.
-Second, this behavior (significant gap) remains evident across the entire spectrum of singular values, implying that the sum of the squares of these values (i.e., the Frobenius norm) is also larger for key matrices.
-Consequently, the stronger representational capacity of key matrices, as reflected by both their spectral and Frobenius norms, renders them more sensitive to quantization errors.
-Appendix D Extended Quantization Error Results
-Tables 3, 4, and 5 present extended quantization error results across a wider range of models and settings for the MMLU, C4, and GSM8K datasets, respectively.
-These tables offer a detailed breakdown of the error metrics under various bit-width configurations, complementing the analysis presented in the main text and in Table 1.
 Figure 5: Complete singular value distribution of key and value cache for the Llama 3.3 70B model on the C4 dataset.
@@
 This full-spectrum view highlights that key matrices consistently maintain significantly higher singular values throughout the entire distribution, further reinforcing their dominant representational capacity compared to value matrices.
+Figure 6: Frobenius norm of key and value cache across layers for the Llama 3.3 70B model on the C4 dataset.
+The x-axis represents the layer index.
+The shaded region shows the min-max range across attention heads within each layer, and the solid line indicates the mean norm per layer.
+Key matrices exhibit significantly higher norms compared to value matrices, highlighting their stronger representational capacity.
+Figure 7: Spectral norm of key and value cache across layers for the Llama 3.3 70B model on the C4 dataset.
+The x-axis represents the layer index.
+The shaded region shows the min-max range across attention heads within each layer, and the solid line indicates the mean norm per layer.
+Key matrices exhibit significantly higher norms compared to value matrices, highlighting their stronger representational capacity.
+Appendix F Extended Quantization Error Results
 Table 4: Quantization error (mean ± std) for the key and value caches ($K_{i}$ and $V_{i}$) at 2-bit, 3-bit, and 4-bit quantization, evaluated on the MMLU dataset.
@@
 Llama3.3-70B-it 4.883 ± 1.106 0.112 ± 0.093 0.942 ± 0.198 0.016 ± 0.012 0.206 ± 0.043 0.003 ± 0.003
-Memotron3.1-it 5.125 ± 1.284 0.114 ± 0.094 0.985 ± 0.207 0.016 ± 0.012 0.216 ± 0.046 0.003 ± 0.003
+Nemotron3.1-it 5.125 ± 1.284 0.114 ± 0.094 0.985 ± 0.207 0.016 ± 0.012 0.216 ± 0.046 0.003 ± 0.003
 Phi3-Medium-128K-it 5.063 ± 1.914 0.584 ± 0.559 1.000 ± 0.319 0.087 ± 0.083 0.217 ± 0.068 0.019 ± 0.018
@@
 Llama3.3-70B-it 4.391 ± 1.027 0.121 ± 0.097 0.847 ± 0.175 0.017 ± 0.013 0.186 ± 0.038 0.004 ± 0.003
-Memotron3.1-it 5.367 ± 1.332 0.127 ± 0.105 1.049 ± 0.222 0.018 ± 0.013 0.231 ± 0.049 0.004 ± 0.003
+Nemotron3.1-it 5.367 ± 1.332 0.127 ± 0.105 1.049 ± 0.222 0.018 ± 0.013 0.231 ± 0.049 0.004 ± 0.003
 Phi3-Medium-128K-it 4.831 ± 1.759 0.788 ± 0.726 1.022 ± 0.306 0.109 ± 0.097 0.220 ± 0.064 0.023 ± 0.021
@@
 Llama3.3-70B-it 4.967 ± 1.127 0.113 ± 0.091 0.978 ± 0.203 0.016 ± 0.012 0.214 ± 0.044 0.004 ± 0.003
-Memotron3.1-it 4.752 ± 1.124 0.113 ± 0.089 0.940 ± 0.194 0.016 ± 0.012 0.206 ± 0.042 0.004 ± 0.003
+Nemotron3.1-it 4.752 ± 1.124 0.113 ± 0.089 0.940 ± 0.194 0.016 ± 0.012 0.206 ± 0.042 0.004 ± 0.003
 Phi3-Medium-128K-it 4.940 ± 1.834 0.605 ± 0.579 1.042 ± 0.320 0.088 ± 0.082 0.227 ± 0.069 0.019 ± 0.018
@@
 DeepSeekR1Q-14B 4.477 ± 2.176 1.424 ± 1.752 0.830 ± 0.256 0.200 ± 0.242 0.181 ± 0.058 0.044 ± 0.056
+Figure 8: Mean squared error (MSE) of KV quantization for Llama 3.3-70B on C4.
+We use quantization bit-widths ranging from 2 to 8.
+The x-axis represents the quantization bit-width, while the y-axis shows the MSE on a logarithmic scale.
+A logarithmic scale is used to highlight differences at higher bit-widths, where MSE values decrease significantly and approach zero, particularly at 8-bit.
+Solid lines indicate the mean MSE across layers, while the shaded regions represent the min-max range of errors.
+(a) C4 dataset
+(b) MMLU dataset
+(c) GSM8k dataset
+Figure 9: Quantization error (MSE) of K and V caches in the Llama 3.1 8B model across 32 layers for the (a) C4, (b) MMLU, and (c) GSM8k datasets.
+The top row in each plot shows 2-bit quantization (k2, v2), while the bottom row shows 4-bit quantization (k4, v4).
+Each point represents a different attention head in the corresponding layer.
+The x-axis indicates the layer index, and the y-axis shows the quantization error in MSE.
+(a) C4 dataset
+(b) MMLU dataset
+(c) GSM8k dataset
+Figure 10: Quantization error (MSE) of K and V caches in the Llama 3.1 8B model across 32 layers for the (a) C4, (b) MMLU, and (c) GSM8k datasets.
+The top row in each plot shows 2-bit quantization (k2, v2), while the bottom row shows 4-bit quantization (k4, v4).
+Each point represents a different attention head in the corresponding layer.
+The x-axis indicates the layer index, and the y-axis shows the quantization error in MSE.
+Appendix G Reproducibility and Resources
+Inference was performed using the Hugging Face Transformers [53] and Accelerate [54] libraries, together with FlashAttention [55, 56], to optimize token-generation throughput.
+We integrated both Quanto and HQQ into the Language Model Evaluation Harness [57] to enable systematic and reproducible evaluation of model performance under varying quantization schemes.
+All runs were executed on GPU clusters equipped with NVIDIA A100 (80 GB) using the Quanto backend and NVIDIA H200 (141 GB) using the HQQ backend.
+Appendix H Extended Evaluation Results
+Extended evaluation results for KV cache quantization under the HQQ backend are presented in Tables 910 (GSM8K flexible and strict accuracies), Tables 8 (EQ-Parseable parseable-pass rates and exact-match scores), and Table 7 (CoQA F1 scores).
+Table 7: Full CoQA Accuracies F1.
+Variant Qwen3-0.6B Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct Qwen3-4B Llama-3.1-8B-Instruct Qwen3-8B Qwen3-14B Qwen3-32B
+k1v1 0.129657 0.147672 0.127875 0.222218 0.127204 0.358652 0.272695 0.231077
+k1v2 0.226803 0.236351 0.222816 0.360911 0.206763 0.384316 0.329965 0.327598
+k1v4 0.226441 0.236751 0.229168 0.375805 0.241965 0.385504 0.323921 0.34914
+k1v6 0.214566 0.242716 0.229515 0.378638 0.220753 0.378799 0.34787 0.351566
+k1v8 0.219723 0.240167 0.221966 0.364095 0.241671 0.382309 0.341797 0.3545
+k2v1 0.252746 0.189839 0.576534 0.463689 0.565141 0.598736 0.73744 0.732328
+k2v2 0.334262 0.525527 0.760021 0.72138 0.757424 0.766938 0.801972 0.808949
+k2v4 0.333106 0.558613 0.765639 0.732337 0.762769 0.760775 0.799323 0.808594
+k2v6 0.339966 0.564876 0.764105 0.730226 0.763797 0.765849 0.797442 0.80742
+k2v8 0.331698 0.567622 0.763972 0.733084 0.765828 0.769881 0.79908 0.803599
+k4v1 0.496626 0.575004 0.768665 0.803298 0.773274 0.815164 0.823299 0.817609
+k4v2 0.666081 0.693266 0.797161 0.805846 0.785901 0.822081 0.819957 0.819761
+k4v4 0.691803 0.701052 0.797872 0.806737 0.788327 0.818848 0.816583 0.824813
+k4v6 0.688022 0.701648 0.797325 0.809305 0.784273 0.817312 0.818533 0.823784
+k4v8 0.686547 0.700206 0.79702 0.807328 0.782431 0.817872 0.818202 0.822701
+k6v1 0.649633 0.597133 0.770615 0.801156 0.765218 0.812864 0.817002 0.816527
+k6v2 0.692575 0.699732 0.798369 0.80343 0.784922 0.820546 0.821854 0.825656
+k6v4 0.701547 0.705158 0.79613 0.806657 0.786 0.820736 0.822193 0.824622
+k6v6 0.700741 0.705549 0.795475 0.807525 0.790379 0.81943 0.821026 0.825455
+k6v8 0.693932 0.702742 0.79642 0.806614 0.789155 0.81943 0.820462 0.825005
+k8v1 0.650554 0.598797 0.768704 0.802301 0.767131 0.814654 0.818909 0.818461
+k8v2 0.691099 0.696408 0.798332 0.802632 0.787855 0.822552 0.822365 0.824697
+k8v4 0.70116 0.703779 0.794958 0.807328 0.788884 0.820734 0.822508 0.825976
+k8v6 0.698706 0.70461 0.795362 0.808917 0.786832 0.818963 0.820002 0.825695
+k8v8 0.696944 0.70377 0.794663 0.808457 0.785888 0.820653 0.819874 0.826971
+bf16 0.707878 0.70117 0.794959 0.807871 0.785389 0.820364 0.81985 0.826088
+Table 8: Full EQ-Bench Parseable.
+Variant Qwen3-0.6B Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct Qwen3-4B Llama-3.1-8B-Instruct Qwen3-8B Qwen3-32B
+k1v1 0 0 0 0 0 0 0
+k1v2 0 0 0 0 0 0 0
+k1v4 0 0 0 0 0 0 0
+k1v6 0 0 0 0 0 0 0
+k1v8 0 0 0 0 0 0 0
+k2v1 0 0 8.187 0 35.673 0 2.339
+k2v2 0 16.374 80.702 27.485 85.965 65.497 67.836
+k2v4 0 34.503 84.795 39.766 88.889 70.175 70.175
+k2v6 0 39.181 83.041 36.842 90.059 67.836 69.591
+k2v8 0 38.012 84.211 40.936 88.889 68.421 70.760
+k4v1 0 47.953 58.480 77.778 83.626 98.830 77.778
+k4v2 63.743 97.076 95.322 87.719 97.661 95.322 80.702
+k4v4 65.497 97.076 98.830 84.795 99.415 95.322 80.702
+k4v6 64.328 97.076 98.830 85.380 98.830 93.567 80.702
+k4v8 62.573 97.076 98.830 84.795 98.830 93.567 80.702
+k6v1 23.392 60.234 53.801 81.287 90.643 96.491 79.532
+k6v2 99.415 97.076 94.737 87.719 97.661 94.737 80.702
+k6v4 99.415 97.661 99.415 87.135 98.830 95.906 80.702
+k6v6 99.415 97.661 98.830 86.550 98.830 96.491 80.702
+k6v8 99.415 97.661 98.830 87.135 98.830 97.661 80.702
+k8v1 26.901 60.234 56.725 78.947 92.983 97.076 77.193
+k8v2 99.415 97.076 94.737 87.135 97.661 96.491 79.532
+k8v4 99.415 97.661 99.415 87.719 98.830 95.322 80.702
+k8v6 99.415 97.661 99.415 86.550 98.830 95.906 80.702
+k8v8 99.415 97.661 99.415 86.550 98.830 95.906 80.702
+bf16 100 97.661 99.415 87.719 98.830 94.737 79.532
+Table 9: Full GSM8K Accuracies (Flexible).
+Variant Qwen3-0.6B Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct Qwen3-4B Llama-3.1-8B-Instruct Qwen3-8B Qwen3-14B Qwen3-32B
+k1v1 0.0106141 0.0121304 0.0159212 0.00985595 0.0159212 0.0106141 0.0144049 0.0204701
+k1v2 0.00833965 0.0144049 0.015163 0.00985595 0.0197119 0.0128886 0.0174375 0.0181956
+k1v4 0.0090978 0.0197119 0.025019 0.0144049 0.0204701 0.0075815 0.00985595 0.0106141
+k1v6 0.00985595 0.0106141 0.0272934 0.0174375 0.0197119 0.0075815 0.0045489 0.0159212
+k1v8 0.00833965 0.015163 0.0280516 0.0090978 0.015163 0.0136467 0.0113723 0.0106141
+k2v1 0.0030326 0.0181956 0.015163 0.00833965 0.0500379 0.0128886 0.0204701 0.0144049
+k2v2 0.0030326 0.0288097 0.278241 0.0265353 0.38514 0.128886 0.327521 0.38514
+k2v4 0.0030326 0.0197119 0.351024 0.0432146 0.45489 0.170584 0.412434 0.43442
+k2v6 0.00379075 0.0257771 0.3442 0.040182 0.445792 0.169067 0.396513 0.419257
+k2v8 0.00379075 0.0280516 0.354814 0.0348749 0.460197 0.166793 0.394238 0.448825
+k4v1 0.0227445 0.0727824 0.281274 0.576952 0.517058 0.77862 0.70887 0.721001
+k4v2 0.040182 0.294162 0.639121 0.808188 0.756634 0.87718 0.804397 0.648976
+k4v4 0.040182 0.332828 0.652767 0.830174 0.777862 0.87718 0.823351 0.629265
+k4v6 0.055345 0.33207 0.66793 0.831691 0.762699 0.88097 0.828658 0.619409
+k4v8 0.0416983 0.341168 0.658074 0.826384 0.75815 0.877938 0.835481 0.631539
+k6v1 0.0932525 0.0962851 0.307809 0.632297 0.521607 0.795299 0.742987 0.701289
+k6v2 0.367703 0.3116 0.644428 0.830174 0.761941 0.880212 0.824867 0.59818
+k6v4 0.404852 0.333586 0.653525 0.836998 0.770281 0.877938 0.827142 0.610311
+k6v6 0.420015 0.336619 0.662623 0.843821 0.771797 0.884003 0.8279 0.595906
+k6v8 0.41395 0.341926 0.664898 0.840788 0.771039 0.886277 0.826384 0.608036
+k8v1 0.0879454 0.0955269 0.304018 0.636846 0.529947 0.797574 0.742987 0.708112
+k8v2 0.389689 0.313116 0.64746 0.825625 0.74981 0.882487 0.822593 0.605762
+k8v4 0.416224 0.340409 0.660349 0.832449 0.770281 0.874905 0.831691 0.620925
+k8v6 0.419257 0.341926 0.667172 0.843063 0.767248 0.880212 0.839272 0.608036
+k8v8 0.413192 0.333586 0.666414 0.843063 0.758908 0.880212 0.839272 0.603487
+bf16 0.412434 0.328279 0.663381 0.845337 0.759666 0.87718 0.835481 0.603487
+Table 10: Full GSM8K Accuracies (Strict).
+Variant Qwen3-0.6B Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct Qwen3-4B Llama-3.1-8B-Instruct Qwen3-8B Qwen3-14B Qwen3-32B
+k1v1 0 0 0 0 0 0 0 0
+k1v2 0 0 0 0 0 0 0 0
+k1v4 0 0 0 0 0 0 0.00075815 0
+k1v6 0 0 0.0015163 0 0 0.00075815 0 0
+k1v8 0 0 0 0 0 0 0.00075815 0
+k2v1 0 0.00075815 0.00530705 0 0.0280516 0 0.00075815 0.0015163
+k2v2 0 0.0113723 0.275967 0.00379075 0.379833 0.0667172 0.260045 0.234268
+k2v4 0 0.00985595 0.350265 0.0090978 0.443518 0.103867 0.343442 0.334344
+k2v6 0 0.0121304 0.3442 0.00985595 0.438969 0.115239 0.348749 0.329795
+k2v8 0 0.0136467 0.353298 0.00530705 0.448825 0.109174 0.35254 0.347233
+k4v1 0.0060652 0.0614102 0.292646 0.64746 0.500379 0.785444 0.824867 0.701289
+k4v2 0.0424564 0.297195 0.627748 0.824867 0.750569 0.875663 0.882487 0.726308
+k4v4 0.0439727 0.331312 0.642911 0.843821 0.758908 0.875663 0.877938 0.733889
+k4v6 0.055345 0.331312 0.656558 0.839272 0.747536 0.880212 0.887036 0.733131
+k4v8 0.0485216 0.341926 0.648976 0.834723 0.745262 0.87718 0.887036 0.736922
+k6v1 0.0833965 0.0826384 0.326005 0.698256 0.505686 0.802881 0.830174 0.702047
+k6v2 0.369219 0.307809 0.639121 0.839272 0.74981 0.879454 0.884003 0.70887
+k6v4 0.404852 0.333586 0.643669 0.844579 0.755876 0.876422 0.884761 0.718726
+k6v6 0.421531 0.335102 0.6558 0.852919 0.751327 0.881729 0.884003 0.71721
+k6v8 0.41395 0.341168 0.657316 0.84837 0.752843 0.884761 0.884003 0.721001
+k8v1 0.0849128 0.0841547 0.32373 0.704321 0.512509 0.805155 0.829416 0.695224
+k8v2 0.388173 0.313874 0.639879 0.829416 0.739196 0.881729 0.886277 0.71721
+k8v4 0.416983 0.340409 0.652009 0.845337 0.755876 0.874905 0.88097 0.721001
+k8v6 0.419257 0.339651 0.661107 0.84837 0.749052 0.878696 0.886277 0.722517
+k8v8 0.415466 0.33207 0.659591 0.850644 0.743745 0.876422 0.886277 0.723275
+bf16 0.413192 0.327521 0.657316 0.852161 0.741471 0.874147 0.890826 0.717968

### v3

@@
-Quantize What Counts: Bit Allocation Insights Informed by Spectral Gaps in Keys and Values
+Quantize What Counts: More for Keys, Less for Values
 Abstract
-Large Language Models (LLMs) have introduced significant advancements to the capabilities of Natural Language Processing (NLP) in recent years.
-However, as these models continue to scale in size, memory constraints pose substantial challenge.
-Key and Value cache (KV cache) quantization has been well-documented as a promising solution to this limitation.
-In this work, we provide two novel theorems aimed at enhancing KV quantization methods.
-Our first theorem, termed Key-Value Norm Disparity, states that the key weight matrices by nature carry richer information compared to the value weight matrices, as evidenced by higher spectral and Frobenius norms across most of the layers.
-Our second theorem, Key-Driven Quantization, posits that prioritizing the quantization precision of keys over values induces significant improvements to the overall quantization performance.
-In particular, assigning greater precision to the keys compared to the values achieves a higher degree of precision reduction with minimal impact on model accuracy.
-We validate these theorems through theory and extensive experiments on several state-of-the-art LLM architectures and benchmarks.
-These findings offer valuable guidelines for improving KV cache quantization strategies, facilitating more efficient memory utilization without compromising model performance across diverse NLP tasks.
+Large Language Models (LLMs) suffer inference-time memory bottlenecks dominated by the attention Key-Value (KV) cache, which scales with model size and context length.
+While KV-cache quantization alleviates this cost, bit allocation between keys and values is often tuned heuristically, lacking theoretical grounding and generalizability.
+This paper proposes two theorems that anchor mixed-precision KV quantization in the intrinsic geometry of Transformer models.
+First, key projections systematically have larger spectral and Frobenius norms than value matrices, implying higher information density along the key path.
+Second, for any given memory budget, prioritizing precision for keys over values strictly reduces quantization error and better preserves accuracy.
+Empirical evaluations across various prominent LLMs and benchmarks show that key-favored allocations (e.g., 4-bit keys, 2-bit values) retain up to 98.3% accuracy compared to uniform allocations (e.g., 4-bit for both), while conserving memory.
+These results transform bit allocation from ad hoc tuning into a theoretically grounded, geometry-driven design principle for efficient LLM inference.
 Source code is available at https://github.com/mohsenhariri/spectral-kv.
+Quantize What Counts: More for Keys, Less for Values
+Mohsen Hariri1, Alan Luo1, Weicong Chen1, Shaochen Zhong2, Tianyi Zhang2, Qifan Wang3, Xia Hu2, Xiaotian Han1, Vipin Chaudhary1 1Case Western Reserve University; 2Rice University; 3Meta mohsen.hariri@case.edu
 1 Introduction
-LLMs have introduced significant advancements to the capabilities of NLP via their ability to express and capture intricate ideas [1, 2], and a key reason for these advancements is model scaling.
-GPT has seen a tremendous increase across its four versions: GPT-1 at 117M [3], GPT-2 at 1.5B [4], GPT-3 at 175B [5], and GPT-4 at 1.8T parameters [6].
-In the open-source space, Llama has grown to 2T [7], Mistral Large is now 123B [8], and DeepSeek V3 is at 671B [9].
-However, with this growth in model size comes a prohibitive memory bottleneck in the KV cache [10].
-An increase in parameter count requires an increase in context-window length to accommodate chain-of-thought reasoning, which in turn expands the KV cache central to the attention unit.
-That expansion can impose a restrictive memory bottleneck on the LLM.
-Already, supported context windows in SOTA systems have grown significantly [11, 12, 13, 14], underscoring the need for a solution.
-One powerful remedy is quantization, wherein high-precision numbers are reduced to low-precision formats [15, 16].
-In LLMs, this can be applied to the KV-cache tensors (e.g., from BF16 to INT4) yielding substantial memory savings with minimal performance loss.
+Large Language Models (LLMs) have rapidly scaled in recent years, driving major advances in generative capabilities and reasoning performance Vaswani et al.
+(2017); Sutskever et al.
+(2014).
+Model size has increased by several orders of magnitude: GPT has grown from 117M parameters in GPT-1 Radford et al.
+(2018) to 1.5B in GPT-2 Radford et al.
+(2019), 175B in GPT-3 Brown et al.
+(2020), and 1.8T in GPT-4 OpenAI et al.
+(2024).
+Open-source models have followed a similar trajectory, with Llama reaching 2T parameters Meta AI (2025b), Mistral Large scaling to 123B Mistral AI (2024), and DeepSeek V3 to 671B DeepSeek-AI et al.
+(2025b).
 Size K2V4 K4V2
@@
 14B 0.78 0.91
-14B 0.76 0.87
-Figure 1: Keys carry greater importance.
-(Left): Keys exhibit higher norms than values across all Llama 3.3-70B layers.
-(Right): GSM8K accuracy is higher under $\mathrm{K}_{2}\mathrm{V}_{4}$ (2-bit keys, 4-bit values) than $\mathrm{K}_{4}\mathrm{V}_{2}$.
-This process is known as KV cache quantization.
-Extensive work has been done in designing novel KV quantization frameworks, yet none so far have provided deep investigation into mixed-precision schemes, in particular those which focus on the bit allocations split between the Key and Value caches.
-Instead, previous works have treated the bit allocation as a hyperparameter, seeing it as an object for optimization via grid search rather than taking a deeper dive into the properties of Keys and Values.
-We provide further motivation for the study in the form of a full description of the current literature, which is detailed in Section 2.
-Therefore, there is a present need for a deep exploration into the inherent differences between the Key and the Value caches, as well as a change to the way that quantization is performed.
-This study aims to fill this need by uncovering fundamental theoretical ideas to improve upon KV cache quantization methods.
-A preview of these results is illustrated in Figure 1 (left), showing that the spectral norms of the Keys consistently exceed those of the values.
-Additionally, Figure 1 (right) demonstrates that accuracy across four different models is notably higher for the K4V2 quantization configuration compared to the K2V4 configuration.
-Our main contributions are summarized as follows:
-  • We prove that the expected values of the spectral norm and Frobenius norm are consistently higher for Key weight matrices compared to Value weight matrices (Key-Value Norm Disparity theorem), implying that Key matrices inherently carry more information.
-We provide detailed empirical validation through observational studies on multiple LLMs.
-  • Building on the Key-Value Norm Disparity, we introduce the Key-Driven Quantization theorem, positing that assigning higher bit precision to Keys over Values significantly enhances quantization efficiency.
-Specifically, this approach allows for greater precision reduction while maintaining minimal accuracy loss.
-  • We substantiate the Key-Driven Quantization theorem through two primary forms of evidence: (1) a theoretical analysis explicitly linking it to the previously established Key-Value Norm Disparity, and (2) comprehensive empirical evaluations involving multiple SOTA LLMs, three datasets, and two quantization frameworks.
-Section 2 addresses gaps in the existing literature on KV cache quantization, followed by Section 3.1, which introduces the Key-Value Norm Disparity theorem and its supporting evidence.
-Section 3.2 presents the Key-Driven Quantization theorem along with its corresponding proof, and Section 4 reports the empirical evaluation of both theorems.
-Finally, Section 5 summarizes the main contributions and discusses their limitations.
-2 Related Work
-KV Quantization Background.
-Quantization frameworks can be categorized in several ways.
-One way is the time at which the quantization is performed, which is grouped into training-time and post-training (PTQ) quantization techniques [17].
-Training-time quantization frameworks can struggle with the usual costs of neural network training, including the dependence on labelled data and longer training times, but these frameworks can achieve more accurate performance due to targeting the weights or activations during the training process.
-In contrast, post-training quantization techniques are the reverse, struggling with sometimes less accurate performance due to dealing with the values after the training has been completed but achieving lower computation overhead and no longer relying on labelled data which may not be available [16].
-There are several settings for PTQ: weight-only, weight-activation, and KV quantization.
-Weight-only quantization focuses on the weights and achieves strong performance but can struggle with memory overhead [18, 19, 20].
-Weight-activation quantization focuses on both the weights and the activations and avoids memory overhead but can struggle with achieving high accuracy [21, 22, 23].
-In contrast, KV quantization offers the memory savings seen in weight-activation methods while also maintaining the performance of weight-only quantization [24].
-This is due to the KV cache memory overhead inflation caused by longer sequences [25].
-KV quantization frameworks can be further categorized into three groups based on the exact quantization scheme being used for the keys and values: outlier redistribution, fixed-precision, and mixed-precision [26].
-In outlier redistribution, the outliers in the cache are moved elsewhere or smoothed out.
-Several outlier redistribution frameworks have been built to varying levels of success [22, 20, 23].
-In fixed-precision, the same number of bits is used to quantize certain parts of the cache while in mixed-precision, different numbers of bits are used.
-Fixed-precision frameworks have also been built to varying levels of success, though they all suffer from a similar problem of neglecting the inherent properties within the KV cache as well as the varying importance of tokens [27, 28, 29].
-More focus has been devoted to the third type of KV quantization framework, mixed-precision, but as yet, these works have not looked much into asymmetric allocation into keys and values [30, 24, 31, 32, 33, 34, 35].
-KVQuant, for example, focuses on defining an outlier for each vector, such as channel or token, to achieve what the authors term dense-and-sparse quantization.
-Only KVTuner, QAQ, and SKVQ discuss the asymmetric allocation in some detail, which is described below.
-KV Quantization Gaps.
-As described above, extensive work has been done in designing new KV quantization frameworks, yet none so far have conducted deeper investigation into mixed-precision schemes.
-In particular, no studies have focused on making a bit allocation split between the key matrix and the value matrix, nor have they provided a deep theoretical explanation for the inherent differences between the two matrices.
-Only three studies to date have touched upon this area, with none providing explanations for potential differences between the key and value matrix importance.
-The closest study has been a KV quantization framework called KVTuner, which employs layerwise quantization based on the sensitivity of the layers to the KV cache quantization mode and precision pairs [35].
-The study makes a brief mention of the difference in attention score errors and perplexity score errors, observing that precision pairs with more bits allocated for keys than for values achieves similar accuracy than for even-split pairs, such as K8V4 versus K8V8.
-However, the study makes no further analysis into this observation, providing no reason for why this difference is observed.
-In addition, the quantization approach does not use any underlying insight on this difference between key and values, instead opting for a straightforward grid search over a limited amount of precision pairs such as K8V4, K4V8, K4V2, and so on.
-Therefore, this allocation split of bits is not the core focus of the framework, with the focus instead being on deploying a layerwise quantization scheme.
-Two other studies also touch upon the key-value difference, but like KVTuner, do not provide much depth into the difference.
-One of these studies is a KV quantization framework called SKVQ, which deploys a sliding window filter along the input tokens to remove tokens that fall outside the window frame and that meet certain criteria [33].
-The quantization strategy achieves 2 bit key quantization and 1.5 bit value quantization, but the authors do not explore the possible reasons behind why this difference in bit allocation yielded such powerful results, instead arriving at this bit allocation via hyperparameter search.
-The other study is a KV quantization framework called QAQ, which uses the attention output to perform the key quantization and the self-attention output to perform the value quantization [32].
-Unlike SKVQ and all other KV quantization strategies, QAQ is the first to target the keys and values in different approaches.
-However, the difference in the approach lies in the type of data being used to perform the quantization, not in the exact number of bits being allocated to one versus the other.
-Finally, both SKVQ and QAQ are post-training quantization approaches, making them susceptible to overhead issues during the attention computation.
-Figure 2: Frobenius norm for Llama Family.
-The x-axis represents the layer index in the model while the y-axis represents the Frobenius norm magnitude.
-The norms are higher for the Key weights than for the Value weights across layers.
-3 Dynamics of the Norms of Key and Value
-3.1 Key-Value Norm Disparity
-Before formalizing our finding, a brief examination is presented on how the magnitudes of the key and value projection matrices influence KV cache quantization.
-In the self-attention mechanism, the key weights $W^{k}$ scale the token representations that enter the cache, whereas the value weights $W^{v}$ scale the representations that are read out once the attention scores are formed.
-Because quantization noise is approximately proportional to the dynamic range of the signal, the relative sizes of $\|W^{k}\|$ and $\|W^{v}\|$ determine which part of the cache is more sensitive to a given bit budget.
-Computing norms across four Llama-3 models (Fig.
-2) show that the Frobenius norm of $W^{k}$ consistently exceeds that of $W^{v}$ in almost every layer; the same ordering holds for the spectral norm.
-Intuitively, this systematic gap motivates the formal statement captured in Theorem Key-Value Norm Disparity below.
-Theorem 1 (Key-Value Norm Disparity).
-The expectation of the Frobenius norm for the key weight matrices are greater than that for the value weight matrices.
-$\mathbb{E}\!\bigl{[}\|W^{K}\|_{F}\bigr{]}\;>\;\mathbb{E}\!\bigl{[}\|W^{V}\|_{F}\bigr{]}.$ (1)
-The proof of Theorem 1 is deferred to Appendix B, but its core argument can be sketched as follows: Starting from a standard Xavier initialization[36], the analysis tracks how stochastic-gradient descent updates the Frobenius norms of $W^{k}$ and $W^{v}$.
-Under isotropic-input assumptions, the cross-term $\langle W^{m},\,\partial\mathcal{L}/\partial W^{m}\rangle_{F}$ becomes negligible, so the expected norm growth is governed by the squared gradient magnitudes.
-A careful decomposition of those gradients shows that the Key weight update contains an additional factor of the query weight matrix $Q$, whose own norm increases during training.
-Consequently, the expected squared gradient of $W^{k}$ dominates that of $W^{v}$, causing $\mathbb{E}\bigl{[}\|W^{k}\|_{F}\bigr{]}$ (and, by monotonicity, the spectral norm) to outpace its value counterpart.
-This asymptotic divergence establishes the persistent norm gap asserted Key-Value Norm Disparity.
-Although Fig.
-2 illustrates this disparity for the Llama family, supplementary plots for Mistral models in Appendix D confirm that the same pattern holds across diverse architectures [8].
-The next subsection investigates how this norm disparity influences quantization error, laying the groundwork for data-driven bit-allocation strategies.
-3.2 Key-Driven Quantization
-Theorem Key-Value Norm Disparity establishes a persistent norm gap between the Key and Value weight matrices.
-Since quantization noise scales with the dynamic range of the signal, weight matrices of larger norm inevitably introduce larger absolute errors under identical bit precision.
-Consequently, assigning equal quantization budgets to the Key and Value caches is sub-optimal.
-Consider an additive-residual Transformer block (Layer Norm omitted for clarity [37]),
-$h_{l+1}=h_{l}+W_{l}h_{l},$
-and let $\Delta_{l}$ denote the quantization perturbation applied to $W_{l}$. A single block incurs
+70B 0.76 0.87
+Figure 1: Key cache needs more bits.
+(Left): Spectral norms of the key cache (blue) and value cache (orange) across layers in Llama3.3-70B show that key caches consistently exhibit higher norms.
+(Right): GSM8k accuracy for two schemes: $\text{K}_{2}\text{V}_{4}$, representing 2-bit allocation for the K cache and 4-bit allocation for the V cache and $\text{K}_{4}\text{V}_{2}$, representing 4-bit allocation for the K cache and 2-bit allocation for the V cache, demonstrates that allocating more bits to the key cache maintains strong performance, confirming the efficacy of norm-aware, mixed-precision quantization.
+However, this rapid growth has introduced severe inference-time memory bottlenecks, primarily due to the Key-Value (KV) cache Wei et al.
+(2022).
+As parameter counts increase, context lengths must also grow to support more complex reasoning, which further expands the KV cache and strains GPU memory Sheng et al.
+(2023a).
+Modern systems already support extremely long contexts Google DeepMind (2025); OpenAI (2024, 2025); DeepSeek-AI et al.
+(2025a), reaching up to 10 million tokens Meta AI (2025a), making memory-efficient methods essential.
+KV cache quantization refers to reducing the precision of KV tensors (e.g., BF16 to INT4), which can provide substantial memory savings while maintaining controlled accuracy degradation, provided applied strategically Han et al.
+(2016); Nagel et al.
+(2021).
+Although many KV quantization methods have been proposed, most determine key–value bit splits through ad hoc hyperparameter tuning on cache statistics (e.g., inference-time activations
+) rather than grounding them in intrinsic model properties (e.g., model weights). This raises a fundamental question: How should bits be allocated in a principled and generalizable way?
+To answer, Figure 1 illustrates two key observations.
+First, key caches ($\mathbf{K}$) consistently have larger spectral norms than value caches ($\mathbf{V}$).
+Second, assigning more bits to keys (e.g., $\mathrm{K}_{4}\mathrm{V}_{2}$) improves accuracy across multiple models compared to key-underprovisioned allocations ($\mathrm{K}_{2}\mathrm{V}_{4}$).
+These motivate a deeper investigation into the distinct roles of key-value weights (i.e., $\mathbf{W}^{\mathbf{K}}$ and $\mathbf{W}^{\mathbf{V}}$) in attention mechanisms and their implications for quantization performance.
+Toward this, our contributions are:
+  • We propose the Key-Value Norm Disparity theorem, proving that the expected spectral and Frobenius norms of $\mathrm{W}^{\mathrm{K}}$ predominantly exceed those of $\mathrm{W}^{\mathrm{V}}$ across prominent LLM models (i.e., Llama3 and Mistral herds).
+We then derive the Key-Prioritized Quantization theorem, establishing the theoretical foundation on why assigning higher precision to $\mathrm{K}$ than $\mathrm{V}$ strictly reduces quantization error, enabling greater KV-cache compression while maintaining accuracy.
+  • We corroborate and operationalize these theorems across a diverse set of models (Llama-3.2-1B/3B/8B, Llama-3.3-70B, Phi-4-14B, Qwen3-0.6B/1.7B/4B/8B, DeepSeek-R1, Mistral-0.3-7B), datasets (C4, MMLU, GSM8K, EQ-Bench, CoQA, and LongBench [^1] We purposefully select generative tasks rather than commonsense reasoning tasks to better isolate quantization effects; see Section C.1 for details.), and two quantization backends (Optimum Quanto and HQQ).
+    Notably, $\mathrm{K}_{4}\mathrm{V}_{2}$ retains 98.3% (1-shot) and 94.1% (5-shot) accuracy of $\mathrm{K}_{4}\mathrm{V}_{4}$ accuracy while reducing KV-cache memory by 25%, demonstrating both the theoretical soundness and practical effectiveness of the proposed strategy.
+  • Owing to its efficient one-off tunability, we show that our geometry-driven mixed-precision strategy is orthogonal to existing inference-time KV quantization methods and can be seamlessly integrated to yield synergistic gains.
+In a case study with rotation-based outlier redistribution, combining a key-prioritized quantization ($\mathrm{K}_{4}\mathrm{V}_{2}$) with key-only rotation outperforms $\mathrm{K}_{4}\mathrm{V}_{4}$ by 4.4-18% in accuracy across tasks.
+In contrast, rotating value caches is unnecessary and sometimes detrimental
+2 Background and Related Work
+KV Quantization.
+Quantization methods for LLMs can be categorized by timing: training-time and post-training (PTQ) Gholami et al.
+(2021).
+Training-time quantization integrates quantization into model training, typically achieving higher accuracy by quantizing weights or activations during the optimization process.
+However, it requires labeled data and incurs significant training overhead.
+PTQ applies quantization after training, avoiding retraining costs and labeled data requirements Nagel et al.
+(2021), but sometimes yields lower accuracy.
+PTQ can target different model components: weights, activations, or the KV cache.
+Weight-only quantization Frantar et al.
+(2023a, b); Lin et al.
+(2024) achieves strong accuracy but does not reduce activation or KV memory.
+Weight-activation quantization Dettmers et al.
+(2022); Xiao et al.
+(2023); Shao et al.
+(2024) reduces overall memory but often sacrifices accuracy.
+KV quantization offers the best of both: it targets the rapidly growing KV cache Pope et al.
+(2023), providing activation-level memory savings while maintaining the accuracy of weight-only methods Yue et al.
+(2024).
+Existing KV Quantization Schemes.
+KV quantization methods can be categorized by how they treat keys and values Li et al. (2025a):
+  • Outlier redistribution.
+These methods smooth or relocate outliers (i.e., unusually large activation values that dominate quantization ranges) in KV tensors, e.g., SmoothQuant Xiao et al.
+(2023), AWQ Lin et al.
+(2024), and OmniQuant Shao et al.
+(2024).
+  • Fixed-precision. A single bit-width is used for both keys and values, ignoring their different roles and statistical properties Yao et al. (2022); Sheng et al. (2023b); Liu et al. (2024).
+  • Mixed-precision.
+Different bit-widths are assigned to different parts of the cache Hooper et al.
+(2024); Yue et al.
+(2024); Li et al.
+(2024); Dong et al.
+(2024); Duanmu et al.
+(2024); Zhang et al.
+(2024); Li et al.
+(2025b).
+While mixed-precision schemes are the most flexible, existing methods have not systematically explored asymmetric bit allocation between keys and values.
+KVQuant Hooper et al.
+(2024), for example, focuses on vector-wise outlier handling rather than analyzing keys and values separately.
+Only a few methods have attempted to address this issue, and even then, only superficial insights have been gained.
+Needs and Gaps.
+Despite rapid progress in KV quantization, several needs from the LLM research community remain unmet.
+First, there is a need for principled strategies to guide bit allocation.
+Current frameworks either treat the key-value split as a hyperparameter tuned through grid search Li et al.
+(2025b) or rely on heuristics derived from cache statistics (e.g., activation ranges or distributions collected at inference time) Zhang et al.
+(2023); Liu et al.
+(2025); Yang et al.
+(2024).
+These approaches are costly, model- or data-specific, and provide little theoretical insight into the inherent differences between keys and values.
+Second, there is a need to understand and exploit key-value asymmetry.
+Existing works such as KVTuner Li et al.
+(2025b), SKVQ Duanmu et al.
+(2024), and QAQ Dong et al.
+(2024) briefly observe that allocating more bits to keys can preserve accuracy, but none explain why or propose a generalizable strategy.
+KVTuner reports differences in attention vs.
+perplexity errors across bit pairs without analysis; SKVQ finds asymmetric allocations (e.g., 2-bit keys, 1.5-bit values) through hyperparameter search rather than model structure; and QAQ focuses on different data types for keys and values rather than bit-width asymmetry.
+Third, there is a need for lightweight, modular methods that integrate seamlessly with existing KV quantization frameworks.
+The studies mentioned above often involve complex, runtime-dependent procedures that are difficult to generalize or combine, limiting their drop-in applicability and composability with other techniques.
+Our Perspective.
+We address these needs by deriving bit allocation strategies directly from model weights, analyzing spectral and Frobenius norms of key and value matrices.
+Our approach is lightweight, incurring only a one-off analytical cost per model without any inference-time introspection; generalizable, since weight statistics are invariant across inputs and tasks; and principled, grounding allocation in linear algebraic properties rather than heuristic search.
+Moreover, our mix-precision quantization oracle is orthogonal to existing KV quantization techniques, paving a foundation that others can build upon.
+For example, pairing our mixed-precision allocation with rotation-based outlier redistribution techniques yields complementary improvements in both accuracy and memory savings.
+In this way, we elevate bit allocation from an ad hoc design choice to a geometry-informed building block for future KV quantization frameworks.
+3 Norm Dynamics of KV Weights
+We now establish a theoretical foundation for mixed-precision KV-cache quantization by analyzing the intrinsic geometry of the key and value projection weights.
+Our analysis proceeds in two steps.
+First, we prove that key weights systematically exhibit larger spectral and Frobenius norms than value weights (Key-Value Norm Disparity), a property that is preserved in the resulting key and value caches.
+Second, we demonstrate that this norm gap directly impacts quantization error, providing a formal guarantee that assigning higher bit precision to keys yields strictly lower distortion and higher inference accuracy than symmetric or inverted allocations (Key-Prioritized Quantization).
+3.1 Key-Value Norm Disparity: Key Weights Dominate in Norm
+The key weight matrix $\mathrm{W}^{K}$ maps hidden states into the key cache, whereas the value weight matrix $\mathrm{W}^{V}$ determines the representations retrieved during attention.
+Because quantization error scales approximately with the dynamic range of the signal, the relative magnitudes of $\|\mathrm{W}^{K}\|$ and $\|\mathrm{W}^{V}\|$, e.g., their Frobenius or spectral norms, indicate which cache is more sensitive to quantization.
+Figure 2: Frobenius norms of key and value weight matrices across the Llama 3 family.
+$\|W^{K}\|_{F}$ consistently exceeds $\|W^{V}\|_{F}$ across nearly all layers, with the exception occurring in early layers of the 70B variant.
+Empirical measurements across four Llama-3 herds (Figure 2) show that the Frobenius norm of $\mathrm{W}^{K}$ consistently exceeds that of $\mathrm{W}^{V}$ in nearly every layer; the same ordering holds for spectral norms.
+This persistent gap motivates the following theorem.
+Theorem 1 (Key-Value Norm Disparity) Let $\mathrm{W}^{K}$ and $\mathrm{W}^{V}$ denote the key and value projection matrices in a Transformer.
+Then $\mathbb{E}\bigl[\|\mathrm{W}^{K}\|_{F}\bigr]>\mathbb{E}\bigl[\|\mathrm{W}^{V}\|_{F}\bigr].$
+The detailed proof is provided in Appendix A.2.
+The core idea here is to examine how the Frobenius norms of the key and value weight matrices evolve during training.
+The analysis begins with Xavier initialization Glorot and Bengio (2010), where all projection matrices have identical expected norms, and tracks their evolution under stochastic gradient descent (SGD).
+Intuitively, $\mathrm{W}^{K}$ play a dual role: they shape the attention map and determine what representations are stored in the cache.
+Each input is multiplied by $\mathrm{W}^{K}$ to produce keys that interact multiplicatively with the queries derived from $\mathrm{W}^{Q}$ (i.e., the query projection).
+As training proceeds, $\mathrm{W}^{Q}$ typically grows to sharpen attention, amplifying the gradient signals backpropagated into $\mathrm{W}^{K}$.
+By contrast, $\mathrm{W}^{V}$ only influences post-attention representations, so its gradients lack this multiplicative amplification.
+This architectural asymmetry causes $\mathrm{W}^{K}$ to receive systematically larger updates, leading to persistently larger norms over time.
+This phenomenon is ubiquitous.
+Appendix C.2 offers analogous results for Mistral family Mistral AI (2024), demonstrating the generality of the pattern.
+We next show that this norm disparity has direct consequences for quantization.
+3.2 Key-Prioritized Quantization: Key-Favored Allocation Minimizes Quantization Error
+Theorem 3.1 implies that, on average, $\mathrm{W}^{K}$ and their resulting activations, $\mathrm{K}$, have larger magnitude than their value counterparts.
+Since quantization error under uniform scalar quantization scales with the signal energy, assigning equal bit precision to both is sub-optimal.
+Consider an additive-residual Transformer block (layer normalization omitted for clarity Elhage et al. (2021)):
+$h_{l+1}=h_{l}+W_{l}h_{l}.$
+Quantizing $W_{l}$ to $\tilde{W}_{l}=W_{l}+\Delta_{l}$ introduces an error bounded by
 $\|h_{l+1}-\tilde{h}_{l+1}\|_{2}=\|\Delta_{l}h_{l}\|_{2}\leq\|\Delta_{l}\|_{2}\|h_{l}\|_{2}.$
-After $L$ layers the worst-case deviation satisfies
+After $L$ layers, the worst-case deviation accumulates multiplicatively:
 $\|h_{L}-\tilde{h}_{L}\|_{2}\leq\Bigl(\prod_{i=1}^{L}\|W_{i}\|_{2}\Bigr)\max_{1\leq i\leq L}\|\Delta_{i}\|_{2}.$
-From  Key-Value Norm Disparity:
-$\mathbb{E}\!\bigl{[}\|W^{K}\|_{F}^{2}\bigr{]}>\mathbb{E}\!\bigl{[}\|W^{V}\|_{F}^{2}\bigr{]}.$ (i)
-Hence any distortion injected on the key path is amplified more strongly throughout the layers.
-Let $X\in\mathbb{R}^{\mathrm{seq}\times d_{\mathrm{model}}}$ be the hidden-state matrix at a given layer. The key and value caches are produced by the same input but different projections:
+Because the expected norms satisfy
+$\mathbb{E}\bigl[\|W^{K}\|_{F}^{2}\bigr]>\mathbb{E}\bigl[\|W^{V}\|_{F}^{2}\bigr],$
+any quantization noise injected along the key path is amplified more strongly through the network.
+Let $X\in\mathbb{R}^{\text{seq}\times d_{\text{model}}}$ be the hidden-state matrix at a given layer. The same input generates both caches via
 $K=XW^{K},\qquad V=XW^{V}.$
-Because $X$ is common to both paths, multiplying (i) on the left by $X$ and using the sub-multiplicativity of the Frobenius norm, $\|AB\|_{F}\leq\|A\|_{F}\,\|B\|_{2}$, yields
-$\mathbb{E}\!\bigl{[}\|K\|_{F}^{2}\bigr{]}>\mathbb{E}\!\bigl{[}\|V\|_{F}^{2}\bigr{]},$ (ii)
-i.e. the norm gap persists in the caches themselves.
-Quantization error vs. Frobenius norm.
-For a matrix $M\in\mathbb{R}^{m\times n}$ the squared Frobenius norm $\|M\|_{F}^{2}$ equals the total energy $\sum_{i,j}M_{ij}^{2}$, or equivalently the sum of eigenvalues of $M^{\top}M$ [^1] Because $\operatorname{tr}(AB)=\operatorname{tr}(BA)$, one has $\operatorname{tr}\!\bigl{(}M^{\top}M\bigr{)}=\operatorname{tr}\!\bigl{(}MM^{\top}\bigr{)}$.
-The non-zero eigenvalues of $M^{\top}M$ and $MM^{\top}$ are identical and equal to the squared singular values $\sigma_{i}^{2}$ of $M$; hence $\|M\|_{F}^{2}=\sum_{i}\sigma_{i}^{2}=\operatorname{tr}\!\bigl{(}M^{\top}M\bigr{)}$..
-When $M$ is quantized to $\tilde{M}$ with $b$ bits, uniform-scalar quantizers yield:
-[…]
-where [^2] Big-$\Theta$ notation: $f(n)=\Theta(g(n))$ iff $\exists\,c_{1},c_{2}>0,\;n_{0}$ such that $c_{1}g(n)\leq f(n)\leq c_{2}g(n)$ for all $n\geq n_{0}$.
-the hidden constants depend only on the quantizer variant.
-Since Key and Value caches share the same shape, minimizing mean square error reduces to minimizing $\|\,\cdot\,\|_{F}^{2}$.
-When $\|K\|_{F}\gg\|V\|_{F}$ and equal bit-widths are used,
-$\mathbb{E}\!\bigl{[}\|K-\tilde{K}\|_{F}^{2}\bigr{]}\;\gg\;\mathbb{E}\!\bigl{[}\|V-\tilde{V}\|_{F}^{2}\bigr{]},$
-making the key cache the dominant source of degradation, formally:
-Theorem 2 (Key-Driven Quantization).
-Let $(b_{K},b_{V})$ denote the bit allocations assigned to the key and value caches, respectively, under a symmetric uniform-scalar quantizer.
-For any pair $(b_{K},b_{V})$ and $b_{K}>b_{V}$, the expected inference accuracy is strictly higher than for the swapped allocation $(b_{V},b_{K})$, provided that $\mathbb{E}\!\bigl{[}\|K\|_{F}^{2}\bigr{]}>\mathbb{E}\!\bigl{[}\|V\|_{F}^{2}\bigr{]}$.
-Figure 3: Singular value distribution of key and value activations for Llama 3.3-70B on C4.
-The x-axis represents the singular value indices, ordered from the 5th largest to the smallest, while the y-axis denotes the magnitude of the singular values.
-The shaded region indicates the minimum-to-maximum range across attention heads within each layer, and the solid lines represent the mean singular value magnitude at each index.
-This zoomed-in view reveals that, beyond the largest singular value (spectral norm), key matrices exhibit consistently higher singular values even in the lower ranges, underscoring their greater representational significance compared to value matrices.
-For the complete range, see Figure 5 in Appendix E.
-Appendix 3 and  4 establishes a norm-dependent upper bound on the quantization error of an arbitrary matrix $M$, and Appendix C.6 applies this bound to show that key caches incur proportionally larger errors than value caches.
-To empirically validate the theoretical predictions, we examined the singular value spectra of the K anc V caches across a broad set of models, up to 70B, by extracting every attention head in every layer (Section 4.1).
-Figure 3 displays the singular values from the fifth largest onward, revealing a systematic separation between keys and values; the complete spectra in Appendix E (Figure 5) show the same trend across all modes.
-Aggregate norms reinforce this observation: Figures 7 and 7 plot the layer-wise Frobenius and spectral norms of the $K\!V$ cache in Llama-3.3-70B, consistently indicating larger magnitudes for keys than for values.
-4 Experimental Results
-4.1 Experiment Setup
-We conduct two experiments: (i) a quantization-error analysis and (ii) a downstream evaluation.
-For the quantization-error analysis, we use PyTorch’s weight-packing quantization to measure reconstruction error.
-For the downstream evaluation, we employ hardware-aligned quantization via Hugging Face’s Optimum Quanto library as back-end (token-wise, 2/4-bit quantization) on Llama-3.2-1B, Llama-3.1-8B [38], Phi-4-14B [39] and DeepSeek-R1-Qwen-14B, and the Half-Quadratic Quantization (HQQ) [40] back-end (channel-wise, 1–8-bit quantization) on Llama-3.1-8B, Llama-3.2-1B, Llama-3.2-3B, and Qwen3-0.6B/1.7B/4B/8B [41].
-The quantization-error sweep additionally includes Phi-3-Medium-128K [42], Llama-3.1-Nemotron-70B [43], Mistral-0.3-7B [8], Qwen-2.5-14B and Llama-3.3-70B.
-See Appendix G for full details on compute resources, software tools, and reproducibility.
-Table 1: Quantization errors for KV cache under various configurations.
-The quantization error is expressed as mean ± std.
-We use 2-bit, 3-bit, and 4-bit quantization for both $K_{i}$ and $V_{i}$ across three datasets.
+Multiplying the previous inequality by $X$ and applying sub-multiplicativity of the Frobenius norm yields
+$\mathbb{E}\bigl[\|K\|_{F}^{2}\bigr]>\mathbb{E}\bigl[\|V\|_{F}^{2}\bigr],$
+i.e., the norm gap persists in the caches.
+Quantization Error and Norm Magnitude.
+For a matrix $M\in\mathbb{R}^{m\times n}$, the squared Frobenius norm equals the total signal energy and the sum of squared singular values.
+When quantized with $b$-bit uniform scalar quantization, the expected mean-square error satisfies
+$\mathbb{E}\bigl[\|M-\tilde{M}\|_{F}^{2}\bigr]=\Theta\bigl(\|M\|_{F}^{2}\,2^{-2b}\bigr),$
+with constants depending only on the quantizer.
+Since key and value caches have identical shape, minimizing quantization error reduces to allocating bits in proportion to their energy.
+When $\|K\|_{F}\gg\|V\|_{F}$ and equal bit-widths are used, key-cache error dominates:
+$\mathbb{E}\bigl[\|K-\tilde{K}\|_{F}^{2}\bigr]\gg\mathbb{E}\bigl[\|V-\tilde{V}\|_{F}^{2}\bigr].$
+This asymmetry in quantization error directly motivates an asymmetric bit allocation strategy, formalized as the following theorem:
+Theorem 2 (Key-Prioritized Quantization) Let $(b_{K},b_{V})$ denote the bit allocations for key and value caches under a uniform scalar quantizer.
+For any pair with $b_{K}>b_{V}$, the expected inference accuracy is strictly higher than for the swapped allocation $(b_{V},b_{K})$, provided that $\mathbb{E}[\|K\|_{F}^{2}]>\mathbb{E}[\|V\|_{F}^{2}]$.
+Figure 3 provides empirical evidence for this effect.
+For Llama 3.3-70B on C4, the singular value spectra of the key caches consistently exceed those of the value caches beyond the top singular mode, indicating higher representational significance throughout the spectrum.
+Appendix C.3 extends this analysis to the full singular value range, where Figure 7 shows layer-wise Frobenius and spectral norms, all consistently revealing larger magnitudes for keys than for values.
+The theoretical underpinnings of this phenomenon are established in Theorems B.2 of Appendix B.2 and B.3 of Appendix B.3, which derive a norm-dependent upper bound on the quantization error of an arbitrary matrix M.
+Appendix B.4 then applies this bound to the key and value caches, showing that the larger norms of the key caches translate into proportionally higher quantization errors under equal bit allocations.
+Figure 3: Singular value spectra of key and value activations in Llama 3.3-70B on C4 benchmark dataset.
+The x-axis shows singular value indices, ordered from the 5th largest onward for cleaner illustration, and the y-axis shows their magnitudes.
+Shaded regions mark the minimum-maximum range across attention heads within each layer, while dashed lines indicate the mean at each index.
+Beyond the top singular value (i.e., the spectral norm), key activations consistently exhibit larger singular values than value activations across the spectrum, highlighting their greater representational capacity.
+Full spectra are provided in Figure 6 of Appendix C.3.
+4 Results
+4.1 Experimental Setup
+Three evaluations are conducted: (i) a quantization-error analysis, which measures reconstruction error across different bit-widths; (ii) a downstream task accuracy evaluation, which assesses how mixed-precision KV cache quantization affects model accuracy on practical benchmarks; and (iii) an integration case study with rotation-based outlier distribution methods, which investigates the effect of combining mixed precisions with rotation strategies.
+Full details on compute resources and software configurations are supplied in Appendix D.
+Table 1: Quantization error (lower is better) of $K$ and $V$ caches under matched bit-widths.
+Each cell reports the mean $\pm$ standard deviation of the reconstruction MSE between the dequantized cache and its BF16 reference, averaged over layers, heads, and tokens (10 sequences per dataset up to 1,000 tokens each).
+Both caches are quantized to the same precision, i.e., $\mathbf{K}_{2}\mathbf{V}_{2}$, $\mathbf{K}_{3}\mathbf{V}_{3}$, and $\mathbf{K}_{4}\mathbf{V}_{4}$, to isolate their intrinsic sensitivity at equal bit budgets.
+Across model families (Llama, Phi, Mistral, Qwen, DeepSeek) and datasets (C4, MMLU, GSM8K), $\mathbf{K}$ consistently exhibits larger reconstruction error than $\mathbf{V}$ at the same bit-width, and the error decreases monotonically with increasing precision, indicating that keys are the dominant source of quantization distortion.
 Dataset Model 2-bit 3-bit 4-bit
-$K_{2}$ $V_{2}$ $K_{3}$ $V_{3}$ $K_{4}$ $V_{4}$
+$\mathbf{K}_{\mathbf{2}}$ $\mathbf{V}_{\mathbf{2}}$ $\mathbf{K}_{\mathbf{3}}$ $\mathbf{V}_{\mathbf{3}}$ $\mathbf{K}_{\mathbf{4}}$ $\mathbf{V}_{\mathbf{4}}$
 MMLU Llama3.2-1B 4.851 ± 1.037 0.127 ± 0.101 1.037 ± 0.265 0.021 ± 0.015 0.227 ± 0.059 0.005 ± 0.003
@@
 Qwen2.5-14B 5.184 ± 2.241 1.270 ± 1.547 1.005 ± 0.288 0.182 ± 0.221 0.223 ± 0.067 0.040 ± 0.052
-DeepSeekR1Q-14B 5.126 ± 2.375 1.406 ± 1.609 0.900 ± 0.269 0.198 ± 0.226 0.199 ± 0.062 0.044 ± 0.052
+R1Q-14B 5.126 ± 2.375 1.406 ± 1.609 0.900 ± 0.269 0.198 ± 0.226 0.199 ± 0.062 0.044 ± 0.052
 C4 Llama3.2-1B 4.885 ± 1.056 0.207 ± 0.166 1.074 ± 0.289 0.030 ± 0.024 0.233 ± 0.062 0.006 ± 0.005
@@
 DeepSeekR1Q-14B 4.477 ± 2.176 1.424 ± 1.752 0.830 ± 0.256 0.200 ± 0.242 0.181 ± 0.058 0.044 ± 0.056
-4.2 Quantization Error
-In several studies (Table 3 of KIVI[29], Table 8 of GEAR[44], and Table 4 of RotateKV[45]), quantized KV cache representations have been shown to outperform their BF16 full-precision counterparts, an outcome that may seem counterintuitive.
-Taking BF16 as the gold standard, we attribute this phenomenon to the implicit regularization effect introduced by quantization: by discretizing activation values, quantization injects a small, structured noise into the model’s intermediate representations.
-Much like dropout or weight decay, this noise can smooth overly sharp decision boundaries and mitigate over-fitting, thereby improving generalization on held-out tasks.
-To isolate and measure this effect, our first experiment focuses on characterizing the quantization error in the KV cache, without grouping tokens, or using residual tokens in each quantization turn.
-In addition to per layer and per head comparison, we average MSE across all heads, layers, and tokens in a batch.
-This approach yields a global view of how severely each bit-width compresses $\mathbf{K}$ vs.
-$\mathbf{V}$.
-The evaluation uses ten random sequences from C4 [46], MMLU [47], and GSM8K [48], padded to the longest length and generated autoregressively for up to 1 000 tokens.
-Table 1 presents 2-, 3-, and 4-bit errors for seven representative models (Llama3, Phi4 [39], Mistral, Qwen, DeepSeek-Distilled, and others); the complete results appear in Appendix F.
-Figure 8 plots the full error curve for Llama-3.3-70B on C4, while Figure 10 shows the per-layer breakdown for Llama-3.1-8B.
-4.3 Evaluation
-Experiments in Section 4.2 established that K cache incur larger MSE than V cache at every bit-width, suggesting that $\mathbf{K}$ contains proportionally richer information.
-However, ultimately model quality is measured by downstream metrics such as accuracy rather than reconstruction error.
-Our second experiments therefore measures how the observed error asymmetry translates into task performance under realistic, deployment-oriented constraints.
-KV cache quantization methods rarely de-quantize the entire cache after every generated token because the associated memory traffic defeats the purpose of quantization.
-Instead, they maintain a residual buffer that holds the most recent tokens in full-precision and flushes it en bloc once full.
-We replicate this strategy by (i) reserving a fixed-size residual cache of 64 tokens and (ii) grouping activations into 128-element blocks—settings adopted by popular open-source implementations such as KIVI, Flash-Decoding [49] and Marlin [50].
-Two hardware-aligned quantization kernels were selected to span the design space:
-  • Quanto: token-wise (per-row) quantization applied to the key and value caches, supporting $\{2,4\}$-bit precision.
-  • HQQ: channel-wise (per-column) quantization applied to the key and value caches, supporting $\{1,2,4,6,8\}$-bit precision.
-For downstream evaluation, we employ three generative datasets that require the model to produce extended outputs rather than select among fixed choices: the GSM8K, the CoQA [51] conversational question answering dataset, and the EQ-Bench [52] long-form structured QA suite.
-Results.
-Tables 2 and 3 corroborate the norm-based hypothesis by showing that downstream accuracy is far more sensitive to the precision assigned to the key cache than to the value cache.
-  • Quanto backend.
-Across four models and two prompt lengths, allocating $4$ bits to keys and $2$ bits to values ($\mathbf{K}_{4}\mathbf{V}_{2}$) recovers on average $94\%$ of the full-precision score, whereas the inverted scheme ($\mathbf{K}_{2}\mathbf{V}_{4}$) loses up to $30$ percentage points (pp).
-The margin widens with model capacity: in 1-shot GSM8K, $\mathbf{K}_{4}\mathbf{V}_{2}$ outperforms $\mathbf{K}_{2}\mathbf{V}_{4}$ by $+0.30$ pp on Llama-3.2-1B and by $+0.16$ pp on Phi-4-14B.
-Notably, $\mathbf{K}_{4}\mathbf{V}_{2}$ attains almost the same accuracy as the symmetric $4$-bit baseline.
-  • HQQ backend.
-The broader sweep over $\{1,2,4,6,8\}$ bits and three datasets paints the same picture.
-On GSM8K, granting even a single extra bit to keys yields double-digit gains; e.g., Qwen3-8B improves from $0.08$ to $0.54$ accuracy ($+46$ pp) when the extra bit is moved from values to keys.
-Gaps persist on EQ-Bench and CoQA, where key-centric quantization delivers up to $+62$ pp (Qwen3-0.6B, EQ-Parseable, $1$-bit) and never underperforms the converse setting.
-Advantages taper but do not vanish at higher precisions: at $6$ bits the key-first policy still adds $7\!-\!12$ pp for 8–14B models.
-Table 2: KV cache quantization on GSM8K 1-shot and 8-shot performance employing Optimum Quanto backend.
-$\mathbf{K}_{i}\mathbf{V}_{j}$ denotes $i$-bit quantization for the K cache and $j$-bit for the V cache using Quanto backend.
-The $\mathbf{K}_{\mathbf{(4bit)}}\mathbf{V}_{\mathbf{(2bit)}}$ column, highlighted below, shows the optimal bit allocation setting.
-1-shot 8-shot
-Model $\mathbf{K}_{\mathbf{2}}\mathbf{V}_{\mathbf{2}}$ $\mathbf{K}_{\mathbf{2}}\mathbf{V}_{\mathbf{4}}$ $\mathbf{K}_{\mathbf{4}}\mathbf{V}_{\mathbf{2}}$ $\mathbf{K}_{\mathbf{4}}\mathbf{V}_{\mathbf{4}}$ $\mathbf{K}_{\mathbf{2}}\mathbf{V}_{\mathbf{2}}$ $\mathbf{K}_{\mathbf{2}}\mathbf{V}_{\mathbf{4}}$ $\mathbf{K}_{\mathbf{4}}\mathbf{V}_{\mathbf{2}}$ $\mathbf{K}_{\mathbf{4}}\mathbf{V}_{\mathbf{4}}$
-Llama3.2-1B-it 0.033 0.035 0.338 0.357 0.031 0.031 0.289 0.369
-Llama3.1-8B-it 0.511 0.547 0.752 0.754 0.408 0.441 0.770 0.782
-Phi4-14B 0.759 0.783 0.913 0.923 0.771 0.815 0.927 0.931
-DeepSeekR1Q-14B 0.772 0.775 0.865 0.867 0.763 0.792 0.876 0.875
-22footnotetext: Although Chain-of-Thought prompting is generally discouraged for reasoning models and can even impair their performance, we deliberately apply 1-shot and 5-shot CoT prompts here to examine the effect of longer input prompts on KV cache quantization.
-Table 3: KV cache quantization comparison across GSM8K, EQ-Parseable, and CoQA datasets.
-The results are obtained on the HQQ quantization backend.
-Each cell shows accuracies for $\mathbf{K}_{i}\mathbf{V}_{x}$ and $\mathbf{K}_{x}\mathbf{V}_{i}$, where $x$ denotes the mean accuracy across quantization bit-widths $\{1,2,4,6,8\}$, with the percentage-point gap annotated above the comparison sign.
-See Appendix H for all settings.
+Quantization-Error Evaluation.
+PyTorch’s weight-packing quantization is applied without residual buffers or activation grouping to isolate quantization effects.
+Ten random sequences are sampled from C4 Dodge et al.
+(2021), MMLU Hendrycks et al.
+(2021), and GSM8K Cobbe et al.
+(2021), padded to the longest length, and generated autoregressively for up to 1,000 tokens.
+Both per-layer and per-head reconstruction errors are computed, along with global averages across all heads, layers, and tokens, to offer a comprehensive view of bit-width sensitivity.
+This experiment spans seven models, including Llama-3.2-1B, Llama-3.1-8B Grattafiori et al.
+(2024), Phi-4-14B Abdin et al.
+(2024b), Mistral-0.3-7B Mistral AI (2024), Qwen-2.5-14B, Llama-3.3-70B, DeepSeek-R1-Qwen-14B, Phi-3-Medium-128K Abdin et al.
+(2024a), and Llama-3.1-Nemotron-70B Nvidia et al.
+(2024).
+Downstream Task Accuracy.
+Two representative quantization backends are employed.
+Optimum Quanto applies token-wise (per-row) quantization with mixed 2/4-bit precision on Llama-3.2-1B, Llama-3.1-8B, Phi-4-14B, and DeepSeek-R1-Qwen-14B.
+HQQ Badri and Shaji (2023) applies channel-wise (per-column) quantization with bit-widths $\{1,2,4,6,8\}$ on Llama-3.1-8B, Llama-3.2-1B, Llama-3.2-3B, and Qwen3-0.6B/1.7B/4B/8B Qwen Team (2025).
+A 64-token residual buffer (which stores the most recent tokens in full precision before being periodically flushed) and 128-element activation grouping (which quantizes activations in fixed-size blocks to reduce overhead) are adopted to mirror practical decoding configurations used in KIVI Liu et al.
+(2024), Flash-Decoding Dao et al.
+(2023), and Marlin Frantar et al.
+(2025), ensuring that the evaluation reflects realistic deployment rather than idealized settings.
+Accuracy is measured on three generative benchmarks: GSM8K Cobbe et al.
+(2021), CoQA Reddy et al.
+(2019), and EQ-Bench Paech (2024), which collectively probe mathematical reasoning, conversational QA, and structured long-form generation.
+Integration with Rotation-Based Methods.
+QuaRot Ashkboos et al.
+(2024) is selected for this case study.
+It applies structured randomized Hadamard rotations to activations before quantization, effectively dispersing outliers and improving the uniformity of the quantization distribution.
+A three-dimensional design space is explored, spanning bit-width allocation, group size configuration, and rotation strategies.
+Specifically, mixed-precision settings $\mathrm{K}_{2}\mathrm{V}_{2}$, $\mathrm{K}_{2}\mathrm{V}_{4}$, $\mathrm{K}_{4}\mathrm{V}_{2}$, and $\mathrm{K}_{4}\mathrm{V}_{4}$; key and value group sizes in $\{32,64,128\}$; and four rotation strategies (no rotation, key-only, value-only, and both).
+The evaluation encompasses generative tasks including CoQA, GSM8K, EQ-Bench, and LongBench, utilizing the same seven models as in quantization-error evaluation.
+4.2 Mixed-Precision Quantization Error
+Table 1 reports reconstruction errors at 2-, 3-, and 4-bit precision for seven representative models spanning multiple model families (Llama, Phi, Mistral, Qwen, DeepSeek) and datasets; full results appear in Appendix C.4.
+Figure 8 shows the complete error curves for Llama-3.3-70B on C4, and Figure 10 provides a per-layer breakdown for Llama-3.1-8B.
+Across all models, datasets, and bit-widths, key caches consistently incur larger reconstruction errors than value caches, and this gap remains stable across precision levels.
+These findings empirically support the theoretical prediction that key representations have higher energy and are therefore more sensitive to quantization.
+4.3 Mixed-Precision Downstream Accuracy
+Table 2 summarizes the Optimum Quanto results on GSM8K under both 1-shot and 8-shot Chain-of-Thought (CoT) prompting.
+Although CoT prompting can sometimes reduce reasoning accuracy, it is included here to assess the impact of longer contexts on quantized decoding.
+Across four representative models, including Llama-3.2-1B, Llama-3.1-8B, Phi-4-14B, and DeepSeek-R1-Qwen-14B, a consistent pattern emerges: allocating higher precision to the key cache ($\mathrm{K}_{4}\mathrm{V}_{2}$) preserves accuracy substantially better than the inverse ($\mathrm{K}_{2}\mathrm{V}_{4}$).
+On average, $\mathrm{K}_{4}\mathrm{V}_{2}$ recovers approximately 94% of the full-precision baseline, whereas value-favored allocations incur losses of up to 30 percentage points (pp).
+The performance gap widens with model scale; for instance, under 1-shot GSM8K, the $\mathrm{K}_{4}\mathrm{V}_{2}$ configuration outperforms $\mathrm{K}_{2}\mathrm{V}_{4}$ by 30 pp on Llama-3.2-1B and by 16 pp on Phi-4-14B.
+Notably, $\mathrm{K}_{4}\mathrm{V}_{2}$ nearly matches the symmetric $\mathrm{K}_{4}\mathrm{V}_{4}$ baseline despite halving the value bit budget, indicating that downstream performance is primarily constrained by key precision.
+Table 2: GSM8K – Downstream accuracy with Optimum Quanto (token-wise) mixed-precision KV quantization.
+Token-wise quantization is applied with supported precisions $i,j\in\{2,4\}$, where $\mathrm{K}_{i}\mathrm{V}_{j}$ denotes $i$-bit keys and $j$-bit values.
+Results are shown for both 1-shot and 8-shot settings.
+Across models, the key-favored $\mathrm{K}_{4}\mathrm{V}_{2}$ consistently outperforms the value-favored $\mathrm{K}_{2}\mathrm{V}_{4}$ and approaches the $\mathrm{K}_{4}\mathrm{V}_{4}$ baseline, demonstrating the benefits of prioritizing key precision.
+Model Shots $\mathbf{K}_{2}\mathbf{V}_{2}$ $\mathbf{K}_{2}\mathbf{V}_{4}$ $\mathbf{K}_{4}\mathbf{V}_{2}$ $\mathbf{K}_{4}\mathbf{V}_{4}$
+Llama 3.2-1B-it 1-shot 0.033 0.035 0.338 0.357
+8-shot 0.031 0.031 0.289 0.369
+Llama 3.1-8B-it 1-shot 0.511 0.547 0.752 0.754
+8-shot 0.408 0.441 0.770 0.782
+Phi 4-14B 1-shot 0.759 0.783 0.913 0.923
+8-shot 0.771 0.815 0.927 0.931
+DeepSeek R1Q-14B 1-shot 0.772 0.775 0.865 0.867
+8-shot 0.763 0.792 0.876 0.875
+Table 3: Downstream accuracy with HQQ (channel-wise) mixed-precision KV quantization across GSM8K, EQ-Parseable, and CoQA.
+Each cell compares $\mathbf{K}_{i}\mathbf{V}_{x}$ and $\mathbf{K}_{x}\mathbf{V}_{i}$ for $i\in\{1,2,4,6,8\}$, where $x$ represents the mean accuracy averaged over the other cache’s bit-widths $B=\{1,2,4,6,8\}$.
+This corresponds to contrasting “allocate $i$ bits to keys (values averaged over BB)” versus “allocate $i$ bits to values (keys averaged over BB).” Across model scales from 0.6B to 32B and multiple tasks, key-favored allocations consistently yield higher accuracy, with more pronounced gains observed at lower precision and persistent benefits at higher precision, demonstrating that the key-first advantage generalizes beyond a single model, task, and quantization backend.
 GSM8K $\mathbf{K}_{1}\mathbf{V}_{x}$ vs.
@@
 $\mathbf{K}_{x}\mathbf{V}_{8}$
-Qwen3-0.6B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+3\%}}{{<}}\mathbf{0.03}$ $0.00\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{<}}\mathbf{0.16}$ $0.04\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{<}}\mathbf{0.17}$ $\mathbf{0.34}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.18$ $\mathbf{0.34}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.18$
-Llama-3.2-1B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+5\%}}{{<}}\mathbf{0.05}$ $0.01\stackrel{{\scriptstyle\scriptscriptstyle+18\%}}{{<}}\mathbf{0.19}$ $\mathbf{0.27}\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{>}}0.20$ $\mathbf{0.28}\stackrel{{\scriptstyle\scriptscriptstyle+8\%}}{{>}}0.20$ $\mathbf{0.28}\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{>}}0.21$
+Qwen3-0.6B $00\stackrel{{\scriptstyle\scriptscriptstyle+3\%}}{{<}}\mathbf{0.03} 0.0$$0\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{<}}\mathbf{0.16} 0.0$$04\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{<}}\mathbf{0.17}$ $\mathbf{0.34}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.18$ $\mathbf{0.34}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.18$
+Llama-3.2-1B $00\stackrel{{\scriptstyle\scriptscriptstyle+5\%}}{{<}}\mathbf{0.05} 0.0$$01\stackrel{{\scriptstyle\scriptscriptstyle+18\%}}{{<}}\mathbf{0.19}$ $\mathbf{0.27}\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{>}}0.20$ $\mathbf{0.28}\stackrel{{\scriptstyle\scriptscriptstyle+8\%}}{{>}}0.20$ $\mathbf{0.28}\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{>}}0.21$
 Llama-3.2-3B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+19\%}}{{<}}\mathbf{0.19}$ $0.27\stackrel{{\scriptstyle\scriptscriptstyle+17\%}}{{<}}\mathbf{0.44}$ $\mathbf{0.57}\stackrel{{\scriptstyle\scriptscriptstyle+11\%}}{{>}}0.46$ $\mathbf{0.58}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.46$ $\mathbf{0.59}\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{>}}0.46$
-Qwen3-4B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+41\%}}{{<}}\mathbf{0.41}$ $0.01\stackrel{{\scriptstyle\scriptscriptstyle+49\%}}{{<}}\mathbf{0.50}$ $\mathbf{0.80}\stackrel{{\scriptstyle\scriptscriptstyle+29\%}}{{>}}0.51$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.51$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.51$
+Qwen3-4B $00\stackrel{{\scriptstyle\scriptscriptstyle+41\%}}{{<}}\mathbf{0.41} 0.0$$01\stackrel{{\scriptstyle\scriptscriptstyle+49\%}}{{<}}\mathbf{0.50}$ $\mathbf{0.80}\stackrel{{\scriptstyle\scriptscriptstyle+29\%}}{{>}}0.51$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.51$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.51$
 Llama-3.1-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{<}}\mathbf{0.31}$ $0.35\stackrel{{\scriptstyle\scriptscriptstyle+17\%}}{{<}}\mathbf{0.52}$ $\mathbf{0.70}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.54$ $\mathbf{0.70}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.54$ $\mathbf{0.70}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.54$
-Qwen3-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+48\%}}{{<}}\mathbf{0.48}$ $0.08\stackrel{{\scriptstyle\scriptscriptstyle+46\%}}{{<}}\mathbf{0.54}$ $\mathbf{0.86}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.55$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+32\%}}{{>}}0.55$ $\mathbf{0.86}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.55$
-Qwen3-14B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+50\%}}{{<}}\mathbf{0.50}$ $0.26\stackrel{{\scriptstyle\scriptscriptstyle+32\%}}{{<}}\mathbf{0.58}$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}0.60$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}0.60$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}0.60$
+Qwen3-8B $00\stackrel{{\scriptstyle\scriptscriptstyle+48\%}}{{<}}\mathbf{0.48} 0.0$$08\stackrel{{\scriptstyle\scriptscriptstyle+46\%}}{{<}}\mathbf{0.54}$ $\mathbf{0.86}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.55$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+32\%}}{{>}}0.55$ $\mathbf{0.86}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.55$
 Qwen3-32B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+42\%}}{{<}}\mathbf{0.42}$ $0.25\stackrel{{\scriptstyle\scriptscriptstyle+23\%}}{{<}}\mathbf{0.48}$ $\mathbf{0.73}\stackrel{{\scriptstyle\scriptscriptstyle+23\%}}{{>}}0.50$ $\mathbf{0.71}\stackrel{{\scriptstyle\scriptscriptstyle+21\%}}{{>}}0.50$ $\mathbf{0.72}\stackrel{{\scriptstyle\scriptscriptstyle+21\%}}{{>}}0.51$
 EQ-Parseable
-Qwen3-0.6B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+10\%}}{{<}}\mathbf{10.06}$ $0.00\stackrel{{\scriptstyle\scriptscriptstyle+53\%}}{{<}}\mathbf{52.51}$ $51.23\stackrel{{\scriptstyle\scriptscriptstyle+2\%}}{{<}}\mathbf{52.87}$ $\mathbf{84.21}\stackrel{{\scriptstyle\scriptscriptstyle+32\%}}{{>}}52.63$ $\mathbf{84.91}\stackrel{{\scriptstyle\scriptscriptstyle+33\%}}{{>}}52.28$
-Llama-3.2-1B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+34\%}}{{<}}\mathbf{33.68}$ $25.61\stackrel{{\scriptstyle\scriptscriptstyle+37\%}}{{<}}\mathbf{61.52}$ $\mathbf{87.25}\stackrel{{\scriptstyle\scriptscriptstyle+22\%}}{{>}}65.38$ $\mathbf{90.06}\stackrel{{\scriptstyle\scriptscriptstyle+24\%}}{{>}}66.32$ $\mathbf{90.06}\stackrel{{\scriptstyle\scriptscriptstyle+25\%}}{{>}}66.08$
-Llama-3.2-3B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+36\%}}{{<}}\mathbf{35.44}$ $68.19\stackrel{{\scriptstyle\scriptscriptstyle+5\%}}{{<}}\mathbf{73.10}$ $\mathbf{90.06}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}76.49$ $\mathbf{89.12}\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{>}}76.02$ $\mathbf{89.94}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}76.26$
-Qwen3-4B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+54\%}}{{<}}\mathbf{47.60}$ $29.01\stackrel{{\scriptstyle\scriptscriptstyle+33\%}}{{<}}\mathbf{58.01}$ $\mathbf{84.09}\stackrel{{\scriptstyle\scriptscriptstyle+28\%}}{{>}}59.88$ $\mathbf{85.96}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}59.06$ $\mathbf{85.38}\stackrel{{\scriptstyle\scriptscriptstyle+29\%}}{{>}}59.88$
-Llama-3.1-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+61\%}}{{<}}\mathbf{60.58}$ $\mathbf{77.89}\stackrel{{\scriptstyle\scriptscriptstyle+2\%}}{{>}}75.79$ $\mathbf{95.67}\stackrel{{\scriptstyle\scriptscriptstyle+19\%}}{{>}}77.19$ $\mathbf{96.96}\stackrel{{\scriptstyle\scriptscriptstyle+20\%}}{{>}}77.31$ $\mathbf{97.43}\stackrel{{\scriptstyle\scriptscriptstyle+21\%}}{{>}}77.08$
-Qwen3-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+62\%}}{{<}}\mathbf{58.48}$ $54.39\stackrel{{\scriptstyle\scriptscriptstyle+17\%}}{{<}}\mathbf{70.41}$ $\mathbf{95.32}\stackrel{{\scriptstyle\scriptscriptstyle+25\%}}{{>}}71.35$ $\mathbf{96.26}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}70.76$ $\mathbf{96.14}\stackrel{{\scriptstyle\scriptscriptstyle+26\%}}{{>}}71.11$
-Qwen3-32B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+60\%}}{{<}}\mathbf{47.37}$ $56.14\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{<}}\mathbf{61.75}$ $\mathbf{80.12}\stackrel{{\scriptstyle\scriptscriptstyle+22\%}}{{>}}62.46$ $\mathbf{80.47}\stackrel{{\scriptstyle\scriptscriptstyle+23\%}}{{>}}62.34$ $\mathbf{79.77}\stackrel{{\scriptstyle\scriptscriptstyle+22\%}}{{>}}62.57$
+Qwen3-0.6B $00\stackrel{{\scriptstyle\scriptscriptstyle+10\%}}{{<}}\mathbf{0.10} 0.0$$00\stackrel{{\scriptstyle\scriptscriptstyle+53\%}}{{<}}\mathbf{0.53}$ $0.51\stackrel{{\scriptstyle\scriptscriptstyle+2\%}}{{<}}\mathbf{0.53}$ $\mathbf{0.84}\stackrel{{\scriptstyle\scriptscriptstyle+32\%}}{{>}}0.53$ $\mathbf{0.85}\stackrel{{\scriptstyle\scriptscriptstyle+33\%}}{{>}}0.52$
+Llama-3.2-1B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+34\%}}{{<}}\mathbf{0.34}$ $0.26\stackrel{{\scriptstyle\scriptscriptstyle+37\%}}{{<}}\mathbf{0.62}$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+22\%}}{{>}}0.65$ $\mathbf{0.90}\stackrel{{\scriptstyle\scriptscriptstyle+24\%}}{{>}}0.66$ $\mathbf{0.90}\stackrel{{\scriptstyle\scriptscriptstyle+25\%}}{{>}}0.66$
+Llama-3.2-3B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+36\%}}{{<}}\mathbf{0.35}$ $0.68\stackrel{{\scriptstyle\scriptscriptstyle+5\%}}{{<}}\mathbf{0.73}$ $\mathbf{0.90}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.76$ $\mathbf{0.89}\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{>}}0.76$ $\mathbf{0.90}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.76$
+Qwen3-4B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+54\%}}{{<}}\mathbf{0.48}$ $0.29\stackrel{{\scriptstyle\scriptscriptstyle+33\%}}{{<}}\mathbf{0.58}$ $\mathbf{0.84}\stackrel{{\scriptstyle\scriptscriptstyle+28\%}}{{>}}0.60$ $\mathbf{0.86}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.59$ $\mathbf{0.85}\stackrel{{\scriptstyle\scriptscriptstyle+29\%}}{{>}}0.60$
+Llama-3.1-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+61\%}}{{<}}\mathbf{0.61}$ $\mathbf{0.78}\stackrel{{\scriptstyle\scriptscriptstyle+2\%}}{{>}}0.76$ $\mathbf{0.96}\stackrel{{\scriptstyle\scriptscriptstyle+19\%}}{{>}}0.77$ $\mathbf{0.97}\stackrel{{\scriptstyle\scriptscriptstyle+20\%}}{{>}}0.77$ $\mathbf{0.97}\stackrel{{\scriptstyle\scriptscriptstyle+21\%}}{{>}}0.77$
+Qwen3-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+62\%}}{{<}}\mathbf{0.58}$ $0.54\stackrel{{\scriptstyle\scriptscriptstyle+17\%}}{{<}}\mathbf{0.70}$ $\mathbf{0.95}\stackrel{{\scriptstyle\scriptscriptstyle+25\%}}{{>}}0.71$ $\mathbf{0.96}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}0.71$ $\mathbf{0.96}\stackrel{{\scriptstyle\scriptscriptstyle+26\%}}{{>}}0.71$
+Qwen3-32B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+60\%}}{{<}}\mathbf{0.47}$ $0.56\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{<}}\mathbf{0.62}$ $\mathbf{0.80}\stackrel{{\scriptstyle\scriptscriptstyle+22\%}}{{>}}0.62$ $\mathbf{0.80}\stackrel{{\scriptstyle\scriptscriptstyle+23\%}}{{>}}0.62$ $\mathbf{0.80}\stackrel{{\scriptstyle\scriptscriptstyle+22\%}}{{>}}0.63$
 CoQA
@@
 Qwen3-8B $0.38\stackrel{{\scriptstyle\scriptscriptstyle+37\%}}{{<}}\mathbf{0.68}$ $\mathbf{0.73}\stackrel{{\scriptstyle\scriptscriptstyle+1\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$
-Qwen3-14B $0.32\stackrel{{\scriptstyle\scriptscriptstyle+45\%}}{{<}}\mathbf{0.69}$ $\mathbf{0.79}\stackrel{{\scriptstyle\scriptscriptstyle+8\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$
 Qwen3-32B $0.32\stackrel{{\scriptstyle\scriptscriptstyle+44\%}}{{<}}\mathbf{0.68}$ $\mathbf{0.79}\stackrel{{\scriptstyle\scriptscriptstyle+9\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.73$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.73$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.73$
+The HQQ results, shown in Table 3, extend these observations to a broader range of bit-widths, model scales, and tasks.
+The analysis systematically compares $\mathbf{K}_{i}\mathbf{V}_{x}$ against $\mathbf{K}_{x}\mathbf{V}_{i}$ for $i\in\{1,2,4,6,8\}$, where $x$ denotes the mean accuracy over all bit-widths of the other cache.
+This directly addresses the question: “If $i$ bits are available, should they be allocated to keys or values?” Across all models (0.6B-32B) and datasets (GSM8K, CoQA, EQ-Parseable), the answer is consistently “keys”.
+At ultra-low precision (1-2 bits), the advantage is especially pronounced; for instance, on GSM8K, allocating a single extra bit to keys with 1-bit quantization yields gains of +48 pp for Qwen3-8B.
+Similar trends hold on EQ-Bench and CoQA: for Qwen3-0.6B on EQ-Parseable, prioritizing keys delivers up to +62 pp, and key-first allocations never underperform value-first ones in any configuration.
+Even at moderate precisions (4-6 bits), key-centric allocation continues to offer 7-12 pp improvements for 8-14B models, indicating that the advantage persists well beyond extreme compression.
+On average, $\mathbf{K}_{4}\mathbf{V}_{2}$ retains 98.3% accuracy of $\mathbf{K}_{4}\mathbf{V}_{4}$ (CoQA: 99.2%, EQ-Bench: 99.35%, GSM8K: 97.7%; worst: 88.3%, best: 103.5%).
+Figure 4: Integration of rotation and mixed-precision quantization.
+Downstream accuracy is shown for four quantization configurations ($\mathrm{K}_{2}\mathrm{V}_{2}$, $\mathrm{K}_{2}\mathrm{V}_{4}$, $\mathrm{K}_{4}\mathrm{V}_{2}$, $\mathrm{K}_{4}\mathrm{V}_{4}$) combined with four rotation strategies (none, key-only, value-only, both), using a fixed group size of 64 for both keys and values.
+Results are reported on CoQA, GSM8K, EQ-Bench, and LongBench, enabling a controlled comparison of precision-rotation interactions.
+More detailed downstream accuracy results are provided in Appendix C.5.
+Overall, downstream accuracy is far more sensitive to key precision than to value precision.
+Across both token-wise and channel-wise quantization schemes, model scales, and task types, assigning the higher bit-width to $\mathrm{K}$ consistently yields near-baseline accuracy while substantially reducing KV memory.
+This establishes a simple, backend-agnostic design principle for mixed-precision KV cache quantization: More for keys, less for values.
+4.4 Integrating with Rotation-Based Methods
+Figure 4 visualizes how rotation interacts with key-value bit allocations.
+Across all tasks, applying rotation to keys consistently yields larger gains than applying it to values.
+Notably, applying key-only rotation to the $\mathrm{K}_{4}\mathrm{V}_{2}$ configuration achieves accuracy that closely matches the full $\mathrm{K}_{4}\mathrm{V}_{4}$ baseline, indicating that the primary benefits of rotation arise from mitigating key outliers.
+These trends are consistent across tasks and model scales, reinforcing that rotation is most effective when applied in conjunction with key-favored bit allocation.
+Appendix C.6 presents detailed downstream results illustrating the synergistic effects of integrating rotation with mixed-precision quantization across tasks and models.
+It also examines the impact of key and value group sizes, showing that smaller sizes are beneficial for $\mathrm{K}$ due to higher information density, whereas larger group sizes suffice for $\mathrm{V}$ given lower sensitivity to quantization.
 5 Conclusion
-Large scale language models increasingly spend the majority of their inference footprint on the KV cache, making cache compression a decisive factor in practical deployment.
-This work supplied a principled answer to the bit-allocation problem by unifying theory and experiment.
-  • Key–Value Norm Disparity. We proved that Key weights exhibit larger Frobenius norms than Value weights throughout training, implying a higher information density in the key path.
-  • Key-Driven Quantization.
-Leveraging this norm gap, a second theorem shows that assigning the finer bit-width to keys and the coarser bit-width to values strictly minimises quantization error, thereby preserving downstream accuracy.
-  • Empirical Validation.
-Experiments across nine model families, three generative benchmarks, and two hardware-aligned back-ends confirm the theory: a $4\!\times\!2$ (keys / values) split consistently recovers $\geq 94\%$ of full-precision accuracy while halving KV-cache memory.
-These findings elevate bit allocation from a hyper-parameter tuned by grid search to a design principle derived from model geometry, offering immediate guidance for practitioners and a foundation for hardware–algorithm co-design.
-We used a context length of 2K tokens, a common setting, but will use longer contexts in future.
+As large language models increasingly devote most of their inference cost to KV-cache storage and access, effective cache compression has become essential for practical deployment.
+This work provides a theoretically grounded justification and solution to the bit-allocation problem by bridging model geometry and quantization design.
+We theoretically establish that key projections consistently carry higher information density than value projections.
+Building on this, we show that allocating higher precision to keys and lower precision to values minimizes quantization error.
+Extensive experiments across nine model families, six benchmarks, and two hardware-aligned backends validate this principle: a $\mathrm{K}_{4}\mathrm{V}_{2}$ precision split reliably recovers up to $98.3\%$ of $\mathrm{K}_{4}\mathrm{V}_{4}$ accuracy while significantly reducing memory consumption.
+Moreover, we demonstrate that our geometry-driven strategy is orthogonal to rotation-based outlier redistribution methods, enabling seamless integration and further accuracy gains.
+These findings elevate bit allocation from empirical tuning to a theoretically grounded geometry-driven design principle, providing clear guidance for efficient deployment and future hardware-algorithm co-design.
+Limitations
+While our geometry-driven mixed-precision quantization framework demonstrates both theoretical soundness and practical effectiveness through rigorous analysis and extensive experiments, several limitations remain.
+First, all evaluations are conducted with a maximum context length of 2,000 tokens, which reflects common inference-time configurations but does not fully capture the behavior of models operating at much larger context windows.
+Extending the approach to longer contexts may expose additional challenges, including increased quantization sensitivity and higher memory management overheads.
+Addressing these factors is an important direction for future work.
+Ethical Considerations
+This research aims to reduce the memory and computational costs of large language model inference by utilizing efficient KV-cache quantization.
+Such improvements have the potential to lower energy consumption and broaden access to language models in resource-constrained settings, promoting more sustainable and inclusive deployment.
+However, any compression technique entails accuracy trade-offs, which must be carefully monitored to avoid disproportionate impacts in high-stakes domains such as healthcare, law, or finance.
+Responsible deployment requires thorough evaluation of model behavior under mixed-precision settings, particularly for safety-critical applications.
+Acknowledgment
+This research was supported in part by NSF awards 2117439, 2112606, and 2320952.
 Appendix
 Appendix A Dynamics of the Norms of Key and Value Weight Matrices
-A.1 Preliminary: Dimensions
-$\begin{array}[]{@{}ll@{}}c&\text{sequence length (context size)},\\[4.0pt] d_{m}&\text{embedding dimension},\\[4.0pt] d_{v}=d_{k}=d_{m}/n_{h}=d_{m}&\text{for a single head ($n_{h}=1$)},\\[4.0pt] n_{h}&\text{number of attention heads}.\end{array}$
-A.2 Symbols
-$\begin{array}[]{@{}ll@{}}X&\text{input matrix }(c\times d_{m}),\\ X^{O}&\text{output matrix }(c\times d_{m}),\\ W^{V}&\text{value weights }(d_{m}\times d_{m}),\\ W^{K}&\text{key weights }(d_{m}\times d_{m}),\\ W^{Q}&\text{query weights }(d_{m}\times d_{m}),\\ K=XW^{K}&(c\times d_{m}),\\ Q=XW^{Q}&(c\times d_{m}),\\ V=XW^{V}&(c\times d_{m}),\\ S=QK^{\top}/\sqrt{d_{m}}&(c\times c),\\ A=\operatorname{softmax}(S)&(c\times c),\\ \mathcal{L}&\text{loss function.}\end{array}$
-Appendix B Norm Dynamics
-In this appendix, we compare the long–time behavior of the Frobenius norms of the key ($W^{K}$) and value ($W^{V}$) weight matrices of a single–head self–attention layer trained with stochastic gradient descent (SGD).
+A.1 Preliminaries and Notation
+$c$ Sequence length
+$d_{m}$ Embedding dimension
+$d_{v}=d_{k}=d_{m}/n_{h}$ Single head dimension
+$n_{h}$ Number of attention heads
+[…]
+A.2 Training Dynamics of Frobenius Norms
+We compare the long-time behavior of the Frobenius norms of the key ($W^{K}$) and value ($W^{V}$) weight matrices of a single-head self-attention layer trained with stochastic gradient descent (SGD).
 We show that, under standard isotropic assumptions, $\|W^{K}\|_{F}$ grows faster than $\|W^{V}\|_{F}$.
-B.1 Update rule
-At update step $i$ the weights follow
-$W^{m}_{i+1}\;=\;W^{m}_{i}-\eta\,\frac{\partial\mathcal{L}}{\partial W^{m}_{i}},\qquad m\in\{K,V\},$ (2)
+Update rule. At step $i$, the weights follow
+$W^{m}_{i+1}\;=\;W^{m}_{i}-\eta\,\frac{\partial\mathcal{L}}{\partial W^{m}_{i}},\qquad m\in\{K,V\},$
 with learning rate $\eta$. Squaring the Frobenius norm of both sides gives
-$\|W^{m}_{i+1}\|_{F}^{2}$ $=\|W^{m}_{i}\|_{F}^{2}+\eta^{2}\Bigl{\|}\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr{\|}_{F}^{2}-2\eta\,\Bigl{\langle}W^{m}_{i},\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr{\rangle}_{F},$ (3)
+$\|W^{m}_{i+1}\|_{F}^{2}$ $=\|W^{m}_{i}\|_{F}^{2}+\eta^{2}\Bigl\|\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr\|_{F}^{2}$
+$\quad-2\eta\,\Bigl\langle W^{m}_{i},\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr\rangle_{F}.$
 where
-$\|A\|_{F}$ $\equiv\sqrt{\operatorname{tr}(A^{\top}A)},$ (4)
-$\langle A,B\rangle_{F}$ $\equiv\operatorname{tr}(A^{\top}B).$ (5)
-B.2 Expectation over mini-batches
-Taking an expectation over mini-batches yields
-$\Delta\mathbb{E}\bigl{[}\|W^{m}_{i}\|_{F}^{2}\bigr{]}$ $\equiv\mathbb{E}\bigl{[}\|W^{m}_{i+1}\|_{F}^{2}\bigr{]}-\mathbb{E}\bigl{[}\|W^{m}_{i}\|_{F}^{2}\bigr{]}$ (6)
-[…]
-In high–dimensional weight space the gradient is almost orthogonal to the current weights, making the second term negligible. Hence,
-$\Delta\mathbb{E}\bigl{[}\|W^{m}_{i}\|_{F}^{2}\bigr{]}\;\approx\;\eta^{2}\,\mathbb{E}\Bigl{[}\Bigl{\|}\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr{\|}_{F}^{2}\Bigr{]}.$ (8)
-B.3 Initialization
-Throughout we assume Xavier initialization: each entry of $W^{K}$ and $W^{V}$ is drawn i.i.d. from a zero–mean distribution whose variance preserves the input scale [36].
-B.4 Gradients
+$\|A\|_{F}$ $\equiv\sqrt{\operatorname{tr}(A^{\top}A)},$
+$\langle A,B\rangle_{F}$ $\equiv\operatorname{tr}(A^{\top}B).$
+Expectation over mini-batches. Taking an expectation over mini-batches yields
+$\Delta\mathbb{E}\bigl[\|W^{m}_{i}\|_{F}^{2}\bigr]$ $\equiv\mathbb{E}\bigl[\|W^{m}_{i+1}\|_{F}^{2}\bigr]-\mathbb{E}\bigl[\|W^{m}_{i}\|_{F}^{2}\bigr]$
+$=\eta^{2}\,\mathbb{E}\!\Bigl[\Bigl\|\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr\|_{F}^{2}\Bigr]$
+$\quad-2\eta\,\mathbb{E}\!\Bigl[\bigl\langle W^{m}_{i},\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\bigr\rangle_{F}\Bigr].$
+In high-dimensional weight space, the gradient is almost orthogonal to the current weights, making the second term negligible. Hence,
+$\Delta\mathbb{E}\bigl[\|W^{m}_{i}\|_{F}^{2}\bigr]\;\approx\;\eta^{2}\,\mathbb{E}\Bigl[\Bigl\|\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr\|_{F}^{2}\Bigr].$
+Throughout, we assume Xavier initialization: each entry of $W^{K}$ and $W^{V}$ is drawn i.i.d. from a zero-mean distribution whose variance preserves the input scale Glorot and Bengio (2010).
 The attention outputs are
-$X^{O}\;=\;A\,V\,W^{O},$ (9)
+$X^{O}\;=\;A\,V\,W^{O},$
 where $A=\operatorname{softmax}\!\bigl(QK^{\top}/\sqrt{d_{m}}\bigr)$ and $V=XW^{V}$. Differentials give
-$d\mathcal{L}(X^{O})$ $=\operatorname{tr}\!\bigl{[}(\partial\mathcal{L}/\partial X^{O})^{\!\top}\,AX\,dW^{V}W^{O}\bigr{]}=\operatorname{tr}\!\bigl{[}(AX)^{\top}\,\frac{\partial\mathcal{L}}{\partial X^{O}}\,(W^{O})^{\top}\,dW^{V}\bigr{]},$ (10)
+$d\mathcal{L}(X^{O})$ $=\operatorname{tr}\!\bigl[(\partial\mathcal{L}/\partial X^{O})^{\!\top}\,AX\,dW^{V}W^{O}\bigr]$
+$=\operatorname{tr}\!\bigl[(AX)^{\top}\,\frac{\partial\mathcal{L}}{\partial X^{O}}\,(W^{O})^{\top}\,dW^{V}\bigr].$
 so that
-$\frac{\partial\mathcal{L}}{\partial W^{V}}=(AX)^{\top}\,\frac{\partial\mathcal{L}}{\partial X^{O}}\,(W^{O})^{\top}.$ (11)
+$\frac{\partial\mathcal{L}}{\partial W^{V}}=(AX)^{\top}\,\frac{\partial\mathcal{L}}{\partial X^{O}}\,(W^{O})^{\top}.$
 Similarly, writing $S=QK^{\top}/\sqrt{d_{m}}$,
-$d\mathcal{L}(S)$ $=\operatorname{tr}\!\bigl{[}(\partial\mathcal{L}/\partial S)^{\!\top}\,dS\bigr{]}=\frac{1}{\sqrt{d_{m}}}\,\operatorname{tr}\!\bigl{[}X^{\top}\,(\partial\mathcal{L}/\partial S)^{\!\top}\,Q\,dW^{K}\bigr{]},$ (12)
+$d\mathcal{L}(S)$ $=\operatorname{tr}\!\bigl[(\partial\mathcal{L}/\partial S)^{\!\top}\,dS\bigr]$
+$=\frac{1}{\sqrt{d_{m}}}\,\operatorname{tr}\!\bigl[X^{\top}\,(\partial\mathcal{L}/\partial S)^{\!\top}\,Q\,dW^{K}\bigr].$
 yielding
-$\frac{\partial\mathcal{L}}{\partial W^{K}}=\frac{1}{\sqrt{d_{m}}}\,X^{\top}\,(\partial\mathcal{L}/\partial S)^{\!\top}\,Q.$ (13)
-B.5 Expected gradient norms
-Squaring the Frobenius norm of (11) and taking an expectation under the isotropic-input assumption,
-$\mathbb{E}\!\bigl{[}\|\partial\mathcal{L}/\partial W^{V}\|^{2}_{F}\bigr{]}$ $=c\,\sigma_{x}^{2}\,\sigma_{o}^{2},$ (14)
+$\frac{\partial\mathcal{L}}{\partial W^{K}}=\frac{1}{\sqrt{d_{m}}}\,X^{\top}\,(\partial\mathcal{L}/\partial S)^{\!\top}\,Q.$
+Squaring the Frobenius norm of (A.2) and taking an expectation under the isotropic-input assumption,
+$\mathbb{E}\!\bigl[\|\partial\mathcal{L}/\partial W^{V}\|^{2}_{F}\bigr]$ $=c\,\sigma_{x}^{2}\,\sigma_{o}^{2},$
 where $\mathbb{E}[X^{\top}X]=c\sigma_{x}^{2}I$ and $\mathbb{E}\Bigl[\Bigl(A^{\top}\tfrac{\partial\mathcal{L}}{\partial X^{O}}(W^{O})^{\top}\Bigr)\Bigl(A^{\top}\tfrac{\partial\mathcal{L}}{\partial X^{O}}(W^{O})^{\top}\Bigr)^{\top}\Bigr]=\sigma_{o}^{2}I$.
 For the key weights,
-$\mathbb{E}\!\bigl{[}\|\partial\mathcal{L}/\partial W^{K}\|^{2}_{F}\bigr{]}=\frac{\sigma_{x}^{2}\,\sigma_{s}^{2}}{d_{m}}\,\|Q\|_{F}^{2},$ (15)
+$\mathbb{E}\!\bigl[\|\partial\mathcal{L}/\partial W^{K}\|^{2}_{F}\bigr]=\frac{\sigma_{x}^{2}\,\sigma_{s}^{2}}{d_{m}}\,\|Q\|_{F}^{2},$
 with $\sigma_{s}^{2}$ the entry-wise variance of $\partial\mathcal{L}/\partial S$.
-B.6 Asymptotic comparison
-Substituting these expectations into (8) shows that
-$\mathbb{E}\!\bigl{[}\|W^{K}\|^{2}_{F}\bigr{]}\text{grows as}\;\eta^{2}\,\frac{\sigma_{x}^{2}\sigma_{s}^{2}}{d_{m}}\,\|Q\|_{F}^{2},\quad\text{whereas}\quad\mathbb{E}\!\bigl{[}\|W^{V}\|^{2}_{F}\bigr{]}\;\text{grows as}\;\eta^{2}\,c\sigma_{x}^{2}\sigma_{o}^{2}.$ (16)
+Substituting these expectations into (A.2) shows that
+$\mathbb{E}\!\bigl[\|W^{K}\|^{2}_{F}\bigr]$ grows as $\eta^{2}\,\frac{\sigma_{x}^{2}\sigma_{s}^{2}}{d_{m}}\,\|Q\|_{F}^{2}$, whereas $\mathbb{E}\!\bigl[\|W^{V}\|^{2}_{F}\bigr]$ grows as $\eta^{2}\,c\sigma_{x}^{2}\sigma_{o}^{2}.$
 Because $\|Q\|_{F}^{2}$ itself increases during training, $\|W^{K}\|_{F}$ eventually dominates:
-$\mathbb{E}\!\bigl{[}\|W^{K}\|_{F}\bigr{]}\;>\;\mathbb{E}\!\bigl{[}\|W^{V}\|_{F}\bigr{]}.$ (17)
-As shown in Equation 17, the expectation of the Frobenius norm of $W^{k}$ is greater than that of $W^{v}$.
-[^3] While this analysis assumes a single-head, decoder-only transformer for clarity, the extension to multi-head attention is direct.
-The same conclusion also holds for group-query and multi-query attention.
-Appendix C Quantization and Dequantization Error Bounds
-In this appendix we establishes upper bounds on the quantization error incurred when a matrix is represented using a finite bit-width in two’s-complement format.
+$\mathbb{E}\!\bigl[\|W^{K}\|_{F}\bigr]\;>\;\mathbb{E}\!\bigl[\|W^{V}\|_{F}\bigr].$
+As shown in Equation A.2, the expectation of the Frobenius norm of $W^{k}$ is greater than that of $W^{v}$.
+[^2] The analysis considers a single-head decoder-only Transformer without loss of generality; the conclusions extend directly to multi-head, group-query, and multi-query attention.
+Appendix B Quantization Error Bounds
+We establish upper bounds on the quantization error incurred when a matrix is represented using a finite bit-width in two’s complement format.
 Two theorems are presented: one that characterizes the error in terms of the spectral norm and another in terms of the Frobenius norm.
@@
 Consequently, a matrix that exhibits a larger spectral norm is more susceptible to quantization errors.
-To control the error propagation in such matrices, a higher bit-width is necessary.
+To control error propagation in such matrices, a higher bit width is necessary.
 The second theorem provides an analogous bound for the Frobenius norm,
@@
 Hence, matrices with larger Frobenius norms are also more vulnerable to quantization errors and would benefit from a higher precision during quantization.
-C.1 Preliminaries and Notation
-Let $A\in\mathbb{R}^{m\times n}$ denote a real matrix whose entries are to be stored using a fixed number of bits in two’s-complement representation. The following notation is adopted:
+B.1 Preliminaries
+Let $A\in\mathbb{R}^{m\times n}$ denote a real matrix whose entries are to be stored using a fixed number of bits in two’s complement representation. The following notation is adopted:
   • Bit Depth. Given $b$ bits in two’s-complement format, each representable integer $q$ lies in the interval
@@
   • Maximum Entry Magnitude. Define
-[…]
+    $M\;=\;\max_{1\leq i\leq m,\,1\leq j\leq n}\;|A_{ij}|.$
   • Scale Factor. Set
@@
   • Quantization. Define the integer matrix $Q\in\mathbb{Z}^{m\times n}$ by
-    $Q_{ij}$ $=\mathrm{round}\!\Bigl{(}\frac{A_{ij}}{\alpha}\Bigr{)},$ (18)
+    $Q_{ij}$ $=\mathrm{round}\!\Bigl(\frac{A_{ij}}{\alpha}\Bigr),$
     $\text{with}\quad Q_{ij}$ $\in\{-2^{b-1},\dots,2^{b-1}-1\}.$
@@
 in terms of $b$, $m$, $n$, and the norms of $A$.
-C.2 Spectral Norm Bound
-Theorem 3 (Spectral Norm Quantization Error).
-Let $A\in\mathbb{R}^{m\times n}$ and $b\in\mathbb{N}$ be given. Consider two’s-complement quantization with scale factor
-$\alpha\;=\;\frac{M}{2^{b-1}-1},\quad M\;=\;\max_{i,j}|A_{ij}|.$
-Define
-$Q_{ij}\;=\;\mathrm{round}\!\Bigl{(}\frac{A_{ij}}{\alpha}\Bigr{)}\quad\text{and}\quad\widehat{A}_{ij}\;=\;\alpha\,Q_{ij}.$
-Then, the following bound holds:
-$M$ $\leq\|A-\widehat{A}\|_{2}\leq\sqrt{m\,n}\,\frac{M}{2(2^{b-1}-1)}$ (19)
-$\leq\sqrt{m\,n}\,\frac{\|A\|_{2}}{2(2^{b-1}-1)}.$
-In approximate form for large $b$,
-$\|A-\widehat{A}\|_{2}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{2}.$
+B.2 Spectral Norm Error Bound
+Theorem 3 (Spectral Norm Error Bound for Uniform Quantization) Let $A\in\mathbb{R}^{m\times n}$ and $b\in\mathbb{N}$ be given.
+Consider two’s complement quantization with a scale factor of $\alpha\;=\;\frac{M}{2^{b-1}-1},\quad M\;=\;\max_{i,j}|A_{ij}|.$ Define $Q_{ij}\;=\;\mathrm{round}\!\Bigl(\frac{A_{ij}}{\alpha}\Bigr)\quad\text{and}\quad\widehat{A}_{ij}\;=\;\alpha\,Q_{ij}.$ Then, the following bound holds: $M$ $\leq\|A-\widehat{A}\|_{2}\leq\sqrt{m\,n}\,\frac{M}{2(2^{b-1}-1)}$ $\leq\sqrt{m\,n}\,\frac{\|A\|_{2}}{2(2^{b-1}-1)}.$ In approximate form for large $b$, $\|A-\widehat{A}\|_{2}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{2}.$
 Proof.
@@
 $\|A-\widehat{A}\|_{2}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{2}.$
-C.3 Frobenius Norm Bound
-Theorem 4 (Frobenius Norm Quantization Error).
-Under the same setup, the Frobenius norm of the quantization error satisfies
-$\begin{split}\|A-\widehat{A}\|_{F}&\;\leq\;\sqrt{m\,n}\,{\textstyle\frac{M}{2(2^{b-1}-1)}}\\ &\;\leq\;\sqrt{m\,n}\,{\textstyle\frac{\|A\|_{F}}{2(2^{b-1}-1)}}.\end{split}$ (20)
-In approximate form,
-$\|A-\widehat{A}\|_{F}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{F}.$
+B.3 Frobenius Norm Error Bound
+Theorem 4 (Frobenius Norm Error Bound for Uniform Quantization) Under the same setup as Theorem B.2, the Frobenius norm of the quantization error satisfies $\begin{split}\|A-\widehat{A}\|_{F}&\;\leq\;\sqrt{m\,n}\,{\textstyle\frac{M}{2(2^{b-1}-1)}}\\ &\;\leq\;\sqrt{m\,n}\,{\textstyle\frac{\|A\|_{F}}{2(2^{b-1}-1)}}.\end{split}$ In approximate form, $\|A-\widehat{A}\|_{F}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{F}.$
 Proof.
@@
 Remark. The results indicate that both the spectral norm and Frobenius norm errors satisfy similar approximate bounds:
-$\|A-\widehat{A}\|_{2}$ $\lesssim\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{2},$ (21)
+$\|A-\widehat{A}\|_{2}$ $\lesssim\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{2},$
 $\|A-\widehat{A}\|_{F}$ $\lesssim\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{F}.$
-C.4 Quantizing KV Cache
+B.4 Implications for KV Cache Quantization
 Consider the key and value cache
@@
 where $\ast$ denotes either the spectral norm ($\|\cdot\|_{2}$) or the Frobenius norm ($\|\cdot\|_{F}$). In practice, $\mathbf{K}$ typically exhibits a larger norm than $\mathbf{V}$.
-C.5 Spectral-Norm Perspective
-Standard quantization error bounds yield
-$\|\mathbf{E}_{V}\|_{2}$ $\lesssim\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{V}}}\;\|\mathbf{V}\|_{2},$ (22)
+Spectral-Norm Perspective. Standard quantization error bounds yield
+$\|\mathbf{E}_{V}\|_{2}$ $\lesssim\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{V}}}\;\|\mathbf{V}\|_{2},$
 $\|\mathbf{E}_{K}\|_{2}$ $\lesssim\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{K}}}\;\|\mathbf{K}\|_{2}.$
@@
 Cancelling the common factor $\sqrt{L\,d_{\text{head}}}$ yields
-[…]
+$2^{b_{V}}\,\|\mathbf{V}\|_{2}\;\approx\;2^{b_{K}}\,\|\mathbf{K}\|_{2},$
 or equivalently,
@@
 Since $\|\mathbf{V}\|_{2}<\|\mathbf{K}\|_{2}$, it follows that $b_{K}>b_{V}$.
-C.6 Frobenius-Norm (MSE) Perspective
-Similarly, the Frobenius norm error bounds are given by
-$\|\mathbf{E}_{V}\|_{F}$ $\;\lesssim\;\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{V}}}\;\|\mathbf{V}\|_{F},$ (23)
-$\|\mathbf{E}_{K}\|_{F}$ $\;\lesssim\;\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{K}}}\;\|\mathbf{K}\|_{F}.$
-To achieve comparable Frobenius errors,
+Frobenius Norm (MSE) Perspective.
+The Frobenius norm of the quantization error corresponds directly to the mean-squared error (MSE) when normalized by the number of elements.
+Specifically, for a matrix $\mathbf{A}$ with quantized approximation $\widehat{\mathbf{A}}$, the MSE is
+$\mathrm{MSE}(\mathbf{A},\widehat{\mathbf{A}})\;=\;\frac{1}{\mathrm{nnz}(\mathbf{A})}\,\|\mathbf{A}-\widehat{\mathbf{A}}\|_{F}^{2},$
+where $\mathrm{nnz}(\mathbf{A})$ denotes the number of entries. Thus, controlling the Frobenius norm is equivalent to controlling the MSE up to a scaling factor.
+The quantization error bounds under the Frobenius norm are given by
+$\|\mathbf{E}_{V}\|_{F}$ $\;\lesssim\;\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{V}}}\;\|\mathbf{V}\|_{F},$
+$\|\mathbf{E}_{K}\|_{F}$ $\;\lesssim\;\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{K}}}\;\|\mathbf{K}\|_{F},$
+where $L$ is the sequence length and $d_{\text{head}}$ is the head dimension.
+To ensure comparable Frobenius (or MSE) errors between keys and values, we require
 $\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{V}}}\;\|\mathbf{V}\|_{F}\;\approx\;\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{K}}}\;\|\mathbf{K}\|_{F},$
-which implies
-[…]
-and therefore,
+which simplifies to
+$2^{b_{V}}\,\|\mathbf{V}\|_{F}\;\approx\;2^{b_{K}}\,\|\mathbf{K}\|_{F},$
+and therefore
 $2^{\,b_{K}-b_{V}}\;\approx\;\frac{\|\mathbf{V}\|_{F}}{\|\mathbf{K}\|_{F}}.$
-Again, because $\|\mathbf{V}\|_{F}<\|\mathbf{K}\|_{F}$, it follows that $b_{K}>b_{V}$.
-Appendix D Supplementary Key-Value Weight Norm Analysis
-Here we present additional evidence supporting the Key-Value Norm Disparity theorem established in Section 3.1.
-While Fig.
-2 demonstrates this observation for the Llama model family, the same pattern consistently emerges across Mistral family.
+Since typically $\|\mathbf{V}\|_{F}<\|\mathbf{K}\|_{F}$, it follows that $b_{K}>b_{V}$.
+This reinforces the earlier spectral-norm result: keys should be allocated more bits than values to achieve balanced quantization error under an MSE criterion.
+Appendix C Supplemental Results
+C.1 Focus on Generative Tasks
+We validate and operationalize our theorems across a diverse set of datasets, including C4, MMLU, GSM8K, EQ-Bench, CoQA, and LongBench.
+Our evaluation purposefully focuses on generative tasks, i.e., open-ended generation and free-form responses, because KV-cache quantization primarily affects the decoding phase rather than the prefill phase.
+To assess quantization error in isolation, we use C4 under a free-form decoding setup that mirrors pretraining usage, enabling us to analyze how compression directly distorts activations.
+To quantify downstream impact, we select GSM8K, CoQA, EQ-Bench, and LongBench, all of which rely on generate_until-style evaluation, where the model must produce extended, structured responses.
+In contrast, commonsense reasoning benchmarks (e.g., BoolQ, PIQA, HellaSwag, or DoRA) use log-likelihood scoring over candidate options and do not engage KV-cache quantization unless the input prompt itself is compressed.
+Consequently, such discriminative evaluations are orthogonal to our focus and were intentionally excluded.
+Within LongBench, we surveyed all subtasks (gov_report, lcc, lsht, multi_news, narrativeqa, qasper, qmsum, repobench-p, samsum) and found that only gov_report and qmsum require substantial generation, with gov_report involving the longest outputs.
+Throughout this work, “LongBench” refers specifically to gov_report.
+C.2 Key-Value Weight Norm
+Figure 5 presents the Frobenius norms of key and value weights for the Mistral family, exhibiting the same pattern (keys consistently having higher norms than values) as shown in Figure 2 for the Llama family, further corroborating the Key-Value Norm Disparity theorem established in Section 3.1.
 Figure 5: Frobenius norm plot for Mistral Family.
-The x-axis represents the layer index in the model while the y-axis represents the Frobenius norm magnitude.
+The x-axis represents the layer index in the model, while the y-axis represents the Frobenius norm magnitude.
 The spectral norms are higher for the key weights than for the value weights across layers.
-Appendix E Extended Visualization of Singular Value Distributions
-Figure 5 provides the full-range view of the singular value distribution for the key and value cache at various layers of the Llama 3.3 70B model on the C4 dataset.
-The horizontal axis enumerates the singular values in descending order, beginning with the largest (i.e., the spectral norm), and the vertical axis measures the corresponding magnitudes.
-The shaded region captures the minimum-to-maximum range of singular values across the attention heads within each layer, while the solid curves depict the mean singular value at each rank.
-Figure 5: Complete singular value distribution of key and value cache for the Llama 3.3 70B model on the C4 dataset.
+C.3 Singular Value Distributions
+Figure 6 illustrates the full-spectrum singular value distribution of key and value caches across layers of the Llama 3.3 70B model on the C4 dataset.
+The horizontal axis indexes singular values in descending order, starting from the largest (i.e., the spectral norm), while the vertical axis shows their magnitudes.
+The shaded region represents the minimum-maximum range of singular values across attention heads within each layer, and the solid curves indicate the mean singular value at each rank.
+Figure 6: Complete singular value distribution of key and value cache for Llama 3.3-70B on the C4 dataset.
 The x-axis denotes the singular value indices, ordered from the largest (spectral norm) to the smallest, while the y-axis represents the corresponding magnitudes.
@@
 This full-spectrum view highlights that key matrices consistently maintain significantly higher singular values throughout the entire distribution, further reinforcing their dominant representational capacity compared to value matrices.
-Figure 6: Frobenius norm of key and value cache across layers for the Llama 3.3 70B model on the C4 dataset.
+(a) Frobenius Norm
+(b) Spectral Norm
+Figure 7: Frobenius and spectral norms of key and value caches across layers for the Llama 3.3 70B model on the C4 dataset.
 The x-axis represents the layer index.
-The shaded region shows the min-max range across attention heads within each layer, and the solid line indicates the mean norm per layer.
-Key matrices exhibit significantly higher norms compared to value matrices, highlighting their stronger representational capacity.
-Figure 7: Spectral norm of key and value cache across layers for the Llama 3.3 70B model on the C4 dataset.
-The x-axis represents the layer index.
-The shaded region shows the min-max range across attention heads within each layer, and the solid line indicates the mean norm per layer.
-Key matrices exhibit significantly higher norms compared to value matrices, highlighting their stronger representational capacity.
-Appendix F Extended Quantization Error Results
-Table 4: Quantization error (mean ± std) for the key and value caches ($K_{i}$ and $V_{i}$) at 2-bit, 3-bit, and 4-bit quantization, evaluated on the MMLU dataset.
-Model K2 V2 K3 V3 K4 V4
+The shaded regions indicate the min-max range across attention heads within each layer, while the solid curves represent the mean norm per layer.
+In both cases, key matrices consistently exhibit substantially higher norms than value matrices, reflecting their stronger representational capacity.
+C.4 Quantization Error
+By using quantization error, we refer to evaluating how closely the quantized-then-dequantized key/value (KV) caches reconstruct their full-precision counterparts, measured as the reconstruction mean-squared error (MSE), which is exactly the Frobenius norm error normalized by the number of elements.
+This serves as a practical proxy for downstream quality because uniform $b$-bit quantization admits norm-based bounds in which reconstruction error scales with the cache norm and decays roughly like $2^{-b}$; thus, at a fixed bit budget, higher-norm caches incur larger distortion, and bit allocations that minimize reconstruction error are directly targeting the quantity most affected by quantization.
+Table 4: Quantization error (mean ± std) for the key and value caches ($\mathbf{K}_{i}$ and $\mathbf{V}_{i}$) at 2-bit, 3-bit, and 4-bit quantization, evaluated on the MMLU dataset.
+$\mathbf{K}_{2}$ $\mathbf{V}_{2}$ $\mathbf{K}_{3}$ $\mathbf{V}_{3}$ $\mathbf{K}_{4}$ $\mathbf{V}_{4}$
 Llama3.2-1B 4.851 ± 1.037 0.127 ± 0.101 1.037 ± 0.265 0.021 ± 0.015 0.227 ± 0.059 0.005 ± 0.003
@@
 DeepSeekR1Q-14B 5.126 ± 2.375 1.406 ± 1.609 0.900 ± 0.269 0.198 ± 0.226 0.199 ± 0.062 0.044 ± 0.052
-Table 5: Quantization error (mean ± std) for the key and value caches ($K_{i}$ and $V_{i}$) at 2-bit, 3-bit, and 4-bit quantization, evaluated on the C4 dataset.
-Model K2 V2 K3 V3 K4 V4
+Table 5: Quantization error (mean ± std) for the key and value caches ($\mathbf{K}_{i}$ and $\mathbf{V}_{i}$) at 2-bit, 3-bit, and 4-bit quantization, evaluated on the C4 dataset.
+$\mathbf{K}_{2}$ $\mathbf{V}_{2}$ $\mathbf{K}_{3}$ $\mathbf{V}_{3}$ $\mathbf{K}_{4}$ $\mathbf{V}_{4}$
 Llama3.2-1B 4.885 ± 1.056 0.207 ± 0.166 1.074 ± 0.289 0.030 ± 0.024 0.233 ± 0.062 0.006 ± 0.005
@@
 Table 6: Quantization error (mean ± std) for the key and value caches ($K_{i}$ and $V_{i}$) at 2-bit, 3-bit, and 4-bit quantization, evaluated on the GSM8K dataset.
-Model K2 V2 K3 V3 K4 V4
+$\mathbf{K}_{2}$ $\mathbf{V}_{2}$ $\mathbf{K}_{3}$ $\mathbf{V}_{3}$ $\mathbf{K}_{4}$ $\mathbf{V}_{4}$
 Llama3.2-1B 5.703 ± 1.557 0.179 ± 0.136 1.213 ± 0.352 0.026 ± 0.020 0.266 ± 0.078 0.005 ± 0.004
@@
 DeepSeekR1Q-14B 4.477 ± 2.176 1.424 ± 1.752 0.830 ± 0.256 0.200 ± 0.242 0.181 ± 0.058 0.044 ± 0.056
-Figure 8: Mean squared error (MSE) of KV quantization for Llama 3.3-70B on C4.
+Figure 8: MSE of KV quantization for Llama 3.3-70B on the C4 dataset.
 We use quantization bit-widths ranging from 2 to 8.
@@
 Solid lines indicate the mean MSE across layers, while the shaded regions represent the min-max range of errors.
+(a) C4
+(b) MMLU
+(c) GSM8k
+Figure 9: Quantization error (MSE) of key and value caches in the Llama 3.1 8B model across 32 layers for (a) C4, (b) MMLU, and (c) GSM8K.
+The top row shows 2-bit quantization ($\mathrm{K}_{2}$, $\mathrm{V}_{2}$), and the bottom row shows 4-bit quantization ($\mathrm{K}_{4}$, $\mathrm{V}_{4}$).
+Each point corresponds to an attention head within the respective layer.
+The x-axis denotes the layer index, and the y-axis indicates the quantization error (MSE).
 (a) C4 dataset
@@
 The x-axis indicates the layer index, and the y-axis shows the quantization error in MSE.
-(a) C4 dataset
-(b) MMLU dataset
-(c) GSM8k dataset
-Figure 10: Quantization error (MSE) of K and V caches in the Llama 3.1 8B model across 32 layers for the (a) C4, (b) MMLU, and (c) GSM8k datasets.
-The top row in each plot shows 2-bit quantization (k2, v2), while the bottom row shows 4-bit quantization (k4, v4).
-Each point represents a different attention head in the corresponding layer.
-The x-axis indicates the layer index, and the y-axis shows the quantization error in MSE.
-Appendix G Reproducibility and Resources
-Inference was performed using the Hugging Face Transformers [53] and Accelerate [54] libraries, together with FlashAttention [55, 56], to optimize token-generation throughput.
-We integrated both Quanto and HQQ into the Language Model Evaluation Harness [57] to enable systematic and reproducible evaluation of model performance under varying quantization schemes.
-All runs were executed on GPU clusters equipped with NVIDIA A100 (80 GB) using the Quanto backend and NVIDIA H200 (141 GB) using the HQQ backend.
-Appendix H Extended Evaluation Results
-Extended evaluation results for KV cache quantization under the HQQ backend are presented in Tables 910 (GSM8K flexible and strict accuracies), Tables 8 (EQ-Parseable parseable-pass rates and exact-match scores), and Table 7 (CoQA F1 scores).
-Table 7: Full CoQA Accuracies F1.
-Variant Qwen3-0.6B Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct Qwen3-4B Llama-3.1-8B-Instruct Qwen3-8B Qwen3-14B Qwen3-32B
-k1v1 0.129657 0.147672 0.127875 0.222218 0.127204 0.358652 0.272695 0.231077
-k1v2 0.226803 0.236351 0.222816 0.360911 0.206763 0.384316 0.329965 0.327598
-k1v4 0.226441 0.236751 0.229168 0.375805 0.241965 0.385504 0.323921 0.34914
-k1v6 0.214566 0.242716 0.229515 0.378638 0.220753 0.378799 0.34787 0.351566
-k1v8 0.219723 0.240167 0.221966 0.364095 0.241671 0.382309 0.341797 0.3545
-k2v1 0.252746 0.189839 0.576534 0.463689 0.565141 0.598736 0.73744 0.732328
-k2v2 0.334262 0.525527 0.760021 0.72138 0.757424 0.766938 0.801972 0.808949
-k2v4 0.333106 0.558613 0.765639 0.732337 0.762769 0.760775 0.799323 0.808594
-k2v6 0.339966 0.564876 0.764105 0.730226 0.763797 0.765849 0.797442 0.80742
-k2v8 0.331698 0.567622 0.763972 0.733084 0.765828 0.769881 0.79908 0.803599
-k4v1 0.496626 0.575004 0.768665 0.803298 0.773274 0.815164 0.823299 0.817609
-k4v2 0.666081 0.693266 0.797161 0.805846 0.785901 0.822081 0.819957 0.819761
-k4v4 0.691803 0.701052 0.797872 0.806737 0.788327 0.818848 0.816583 0.824813
-k4v6 0.688022 0.701648 0.797325 0.809305 0.784273 0.817312 0.818533 0.823784
-k4v8 0.686547 0.700206 0.79702 0.807328 0.782431 0.817872 0.818202 0.822701
-k6v1 0.649633 0.597133 0.770615 0.801156 0.765218 0.812864 0.817002 0.816527
-k6v2 0.692575 0.699732 0.798369 0.80343 0.784922 0.820546 0.821854 0.825656
-k6v4 0.701547 0.705158 0.79613 0.806657 0.786 0.820736 0.822193 0.824622
-k6v6 0.700741 0.705549 0.795475 0.807525 0.790379 0.81943 0.821026 0.825455
-k6v8 0.693932 0.702742 0.79642 0.806614 0.789155 0.81943 0.820462 0.825005
-k8v1 0.650554 0.598797 0.768704 0.802301 0.767131 0.814654 0.818909 0.818461
-k8v2 0.691099 0.696408 0.798332 0.802632 0.787855 0.822552 0.822365 0.824697
-k8v4 0.70116 0.703779 0.794958 0.807328 0.788884 0.820734 0.822508 0.825976
-k8v6 0.698706 0.70461 0.795362 0.808917 0.786832 0.818963 0.820002 0.825695
-k8v8 0.696944 0.70377 0.794663 0.808457 0.785888 0.820653 0.819874 0.826971
-bf16 0.707878 0.70117 0.794959 0.807871 0.785389 0.820364 0.81985 0.826088
-Table 8: Full EQ-Bench Parseable.
-Variant Qwen3-0.6B Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct Qwen3-4B Llama-3.1-8B-Instruct Qwen3-8B Qwen3-32B
-k1v1 0 0 0 0 0 0 0
-k1v2 0 0 0 0 0 0 0
-k1v4 0 0 0 0 0 0 0
-k1v6 0 0 0 0 0 0 0
-k1v8 0 0 0 0 0 0 0
-k2v1 0 0 8.187 0 35.673 0 2.339
-k2v2 0 16.374 80.702 27.485 85.965 65.497 67.836
-k2v4 0 34.503 84.795 39.766 88.889 70.175 70.175
-k2v6 0 39.181 83.041 36.842 90.059 67.836 69.591
-k2v8 0 38.012 84.211 40.936 88.889 68.421 70.760
-k4v1 0 47.953 58.480 77.778 83.626 98.830 77.778
-k4v2 63.743 97.076 95.322 87.719 97.661 95.322 80.702
-k4v4 65.497 97.076 98.830 84.795 99.415 95.322 80.702
-k4v6 64.328 97.076 98.830 85.380 98.830 93.567 80.702
-k4v8 62.573 97.076 98.830 84.795 98.830 93.567 80.702
-k6v1 23.392 60.234 53.801 81.287 90.643 96.491 79.532
-k6v2 99.415 97.076 94.737 87.719 97.661 94.737 80.702
-k6v4 99.415 97.661 99.415 87.135 98.830 95.906 80.702
-k6v6 99.415 97.661 98.830 86.550 98.830 96.491 80.702
-k6v8 99.415 97.661 98.830 87.135 98.830 97.661 80.702
-k8v1 26.901 60.234 56.725 78.947 92.983 97.076 77.193
-k8v2 99.415 97.076 94.737 87.135 97.661 96.491 79.532
-k8v4 99.415 97.661 99.415 87.719 98.830 95.322 80.702
-k8v6 99.415 97.661 99.415 86.550 98.830 95.906 80.702
-k8v8 99.415 97.661 99.415 86.550 98.830 95.906 80.702
+C.5 Downstream Accuracy Across Quantization Precisions
+Comprehensive downstream accuracy results for KV cache quantization using the HQQ backend are provided for CoQA (F1 scores), EQ-Parseable (parseable-pass rates and exact-match scores), and GSM8K (flexible and strict accuracy) in Tables 7,  8,  9, and 10, respectively.
+[…]
+Qwen3 Llama-3.2 Llama-3.2 Qwen3 Llama-3.1 Qwen3 Qwen3
+0.6B 1B 3B 4B 8B 8B 32B
+$\mathbf{K}_{1}\mathbf{V}_{1}$ 0.130 0.148 0.128 0.222 0.127 0.359 0.231
+$\mathbf{K}_{1}\mathbf{V}_{2}$ 0.227 0.236 0.223 0.361 0.207 0.384 0.328
+$\mathbf{K}_{1}\mathbf{V}_{4}$ 0.226 0.237 0.229 0.376 0.242 0.386 0.349
+$\mathbf{K}_{1}\mathbf{V}_{6}$ 0.215 0.243 0.230 0.379 0.221 0.379 0.352
+$\mathbf{K}_{1}\mathbf{V}_{8}$ 0.220 0.240 0.222 0.364 0.242 0.382 0.355
+$\mathbf{K}_{2}\mathbf{V}_{1}$ 0.253 0.190 0.577 0.464 0.565 0.599 0.732
+$\mathbf{K}_{2}\mathbf{V}_{2}$ 0.334 0.526 0.760 0.721 0.757 0.767 0.809
+$\mathbf{K}_{2}\mathbf{V}_{4}$ 0.333 0.559 0.766 0.732 0.763 0.761 0.809
+$\mathbf{K}_{2}\mathbf{V}_{6}$ 0.340 0.565 0.764 0.730 0.764 0.766 0.807
+$\mathbf{K}_{2}\mathbf{V}_{8}$ 0.332 0.568 0.764 0.733 0.766 0.770 0.804
+$\mathbf{K}_{4}\mathbf{V}_{1}$ 0.497 0.575 0.769 0.803 0.773 0.815 0.818
+$\mathbf{K}_{4}\mathbf{V}_{2}$ 0.666 0.693 0.797 0.806 0.786 0.822 0.820
+$\mathbf{K}_{4}\mathbf{V}_{4}$ 0.692 0.701 0.798 0.807 0.788 0.819 0.825
+$\mathbf{K}_{4}\mathbf{V}_{6}$ 0.688 0.702 0.797 0.809 0.784 0.817 0.824
+$\mathbf{K}_{4}\mathbf{V}_{8}$ 0.687 0.700 0.797 0.807 0.782 0.818 0.823
+$\mathbf{K}_{6}\mathbf{V}_{1}$ 0.650 0.597 0.771 0.801 0.765 0.813 0.817
+$\mathbf{K}_{6}\mathbf{V}_{2}$ 0.693 0.700 0.798 0.803 0.785 0.821 0.826
+$\mathbf{K}_{6}\mathbf{V}_{4}$ 0.702 0.705 0.796 0.807 0.786 0.821 0.825
+$\mathbf{K}_{6}\mathbf{V}_{6}$ 0.701 0.706 0.795 0.808 0.790 0.819 0.825
+$\mathbf{K}_{6}\mathbf{V}_{8}$ 0.694 0.703 0.796 0.807 0.789 0.819 0.825
+$\mathbf{K}_{8}\mathbf{V}_{1}$ 0.651 0.599 0.769 0.802 0.767 0.815 0.818
+$\mathbf{K}_{8}\mathbf{V}_{2}$ 0.691 0.696 0.798 0.803 0.788 0.823 0.825
+$\mathbf{K}_{8}\mathbf{V}_{4}$ 0.701 0.704 0.795 0.807 0.789 0.821 0.826
+$\mathbf{K}_{8}\mathbf{V}_{6}$ 0.699 0.705 0.795 0.809 0.787 0.819 0.826
+$\mathbf{K}_{8}\mathbf{V}_{8}$ 0.697 0.704 0.795 0.808 0.786 0.821 0.827
+BF16 0.708 0.701 0.795 0.808 0.785 0.820 0.826
+[…]
+Qwen3 Llama-3.2 Llama-3.2 Qwen3 Llama-3.1 Qwen3 Qwen3
+0.6B 1B 3B 4B 8B 8B 32B
+$\mathbf{K}_{1}\mathbf{V}_{1}$ 0 0 0 0 0 0 0
+$\mathbf{K}_{1}\mathbf{V}_{2}$ 0 0 0 0 0 0 0
+$\mathbf{K}_{1}\mathbf{V}_{4}$ 0 0 0 0 0 0 0
+$\mathbf{K}_{1}\mathbf{V}_{6}$ 0 0 0 0 0 0 0
+$\mathbf{K}_{1}\mathbf{V}_{8}$ 0 0 0 0 0 0 0
+$\mathbf{K}_{2}\mathbf{V}_{1}$ 0 0 8.187 0 35.673 0 2.339
+$\mathbf{K}_{2}\mathbf{V}_{2}$ 0 16.374 80.702 27.485 85.965 65.497 67.836
+$\mathbf{K}_{2}\mathbf{V}_{4}$ 0 34.503 84.795 39.766 88.889 70.175 70.175
+$\mathbf{K}_{2}\mathbf{V}_{6}$ 0 39.181 83.041 36.842 90.059 67.836 69.591
+$\mathbf{K}_{2}\mathbf{V}_{8}$ 0 38.012 84.211 40.936 88.889 68.421 70.760
+$\mathbf{K}_{4}\mathbf{V}_{1}$ 0 47.953 58.480 77.778 83.626 98.830 77.778
+$\mathbf{K}_{4}\mathbf{V}_{2}$ 63.743 97.076 95.322 87.719 97.661 95.322 80.702
+$\mathbf{K}_{4}\mathbf{V}_{4}$ 65.497 97.076 98.830 84.795 99.415 95.322 80.702
+$\mathbf{K}_{4}\mathbf{V}_{6}$ 64.328 97.076 98.830 85.380 98.830 93.567 80.702
+$\mathbf{K}_{4}\mathbf{V}_{8}$ 62.573 97.076 98.830 84.795 98.830 93.567 80.702
+$\mathbf{K}_{6}\mathbf{V}_{1}$ 23.392 60.234 53.801 81.287 90.643 96.491 79.532
+$\mathbf{K}_{6}\mathbf{V}_{2}$ 99.415 97.076 94.737 87.719 97.661 94.737 80.702
+$\mathbf{K}_{6}\mathbf{V}_{4}$ 99.415 97.661 99.415 87.135 98.830 95.906 80.702
+$\mathbf{K}_{6}\mathbf{V}_{6}$ 99.415 97.661 98.830 86.550 98.830 96.491 80.702
+$\mathbf{K}_{6}\mathbf{V}_{8}$ 99.415 97.661 98.830 87.135 98.830 97.661 80.702
+$\mathbf{K}_{8}\mathbf{V}_{1}$ 26.901 60.234 56.725 78.947 92.983 97.076 77.193
+$\mathbf{K}_{8}\mathbf{V}_{2}$ 99.415 97.076 94.737 87.135 97.661 96.491 79.532
+$\mathbf{K}_{8}\mathbf{V}_{4}$ 99.415 97.661 99.415 87.719 98.830 95.322 80.702
+$\mathbf{K}_{8}\mathbf{V}_{6}$ 99.415 97.661 99.415 86.550 98.830 95.906 80.702
+$\mathbf{K}_{8}\mathbf{V}_{8}$ 99.415 97.661 99.415 86.550 98.830 95.906 80.702
 BF16 100 97.661 99.415 87.719 98.830 94.737 79.532
-Table 9: Full GSM8K Accuracies (Flexible).
-Variant Qwen3-0.6B Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct Qwen3-4B Llama-3.1-8B-Instruct Qwen3-8B Qwen3-14B Qwen3-32B
-k1v1 0.0106141 0.0121304 0.0159212 0.00985595 0.0159212 0.0106141 0.0144049 0.0204701
-k1v2 0.00833965 0.0144049 0.015163 0.00985595 0.0197119 0.0128886 0.0174375 0.0181956
-k1v4 0.0090978 0.0197119 0.025019 0.0144049 0.0204701 0.0075815 0.00985595 0.0106141
-k1v6 0.00985595 0.0106141 0.0272934 0.0174375 0.0197119 0.0075815 0.0045489 0.0159212
-k1v8 0.00833965 0.015163 0.0280516 0.0090978 0.015163 0.0136467 0.0113723 0.0106141
-k2v1 0.0030326 0.0181956 0.015163 0.00833965 0.0500379 0.0128886 0.0204701 0.0144049
-k2v2 0.0030326 0.0288097 0.278241 0.0265353 0.38514 0.128886 0.327521 0.38514
-k2v4 0.0030326 0.0197119 0.351024 0.0432146 0.45489 0.170584 0.412434 0.43442
-k2v6 0.00379075 0.0257771 0.3442 0.040182 0.445792 0.169067 0.396513 0.419257
-k2v8 0.00379075 0.0280516 0.354814 0.0348749 0.460197 0.166793 0.394238 0.448825
-k4v1 0.0227445 0.0727824 0.281274 0.576952 0.517058 0.77862 0.70887 0.721001
-k4v2 0.040182 0.294162 0.639121 0.808188 0.756634 0.87718 0.804397 0.648976
-k4v4 0.040182 0.332828 0.652767 0.830174 0.777862 0.87718 0.823351 0.629265
-k4v6 0.055345 0.33207 0.66793 0.831691 0.762699 0.88097 0.828658 0.619409
-k4v8 0.0416983 0.341168 0.658074 0.826384 0.75815 0.877938 0.835481 0.631539
-k6v1 0.0932525 0.0962851 0.307809 0.632297 0.521607 0.795299 0.742987 0.701289
-k6v2 0.367703 0.3116 0.644428 0.830174 0.761941 0.880212 0.824867 0.59818
-k6v4 0.404852 0.333586 0.653525 0.836998 0.770281 0.877938 0.827142 0.610311
-k6v6 0.420015 0.336619 0.662623 0.843821 0.771797 0.884003 0.8279 0.595906
-k6v8 0.41395 0.341926 0.664898 0.840788 0.771039 0.886277 0.826384 0.608036
-k8v1 0.0879454 0.0955269 0.304018 0.636846 0.529947 0.797574 0.742987 0.708112
-k8v2 0.389689 0.313116 0.64746 0.825625 0.74981 0.882487 0.822593 0.605762
-k8v4 0.416224 0.340409 0.660349 0.832449 0.770281 0.874905 0.831691 0.620925
-k8v6 0.419257 0.341926 0.667172 0.843063 0.767248 0.880212 0.839272 0.608036
-k8v8 0.413192 0.333586 0.666414 0.843063 0.758908 0.880212 0.839272 0.603487
-bf16 0.412434 0.328279 0.663381 0.845337 0.759666 0.87718 0.835481 0.603487
-Table 10: Full GSM8K Accuracies (Strict).
-Variant Qwen3-0.6B Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct Qwen3-4B Llama-3.1-8B-Instruct Qwen3-8B Qwen3-14B Qwen3-32B
-k1v1 0 0 0 0 0 0 0 0
-k1v2 0 0 0 0 0 0 0 0
-k1v4 0 0 0 0 0 0 0.00075815 0
-k1v6 0 0 0.0015163 0 0 0.00075815 0 0
-k1v8 0 0 0 0 0 0 0.00075815 0
-k2v1 0 0.00075815 0.00530705 0 0.0280516 0 0.00075815 0.0015163
-k2v2 0 0.0113723 0.275967 0.00379075 0.379833 0.0667172 0.260045 0.234268
-k2v4 0 0.00985595 0.350265 0.0090978 0.443518 0.103867 0.343442 0.334344
-k2v6 0 0.0121304 0.3442 0.00985595 0.438969 0.115239 0.348749 0.329795
-k2v8 0 0.0136467 0.353298 0.00530705 0.448825 0.109174 0.35254 0.347233
-k4v1 0.0060652 0.0614102 0.292646 0.64746 0.500379 0.785444 0.824867 0.701289
-k4v2 0.0424564 0.297195 0.627748 0.824867 0.750569 0.875663 0.882487 0.726308
-k4v4 0.0439727 0.331312 0.642911 0.843821 0.758908 0.875663 0.877938 0.733889
-k4v6 0.055345 0.331312 0.656558 0.839272 0.747536 0.880212 0.887036 0.733131
-k4v8 0.0485216 0.341926 0.648976 0.834723 0.745262 0.87718 0.887036 0.736922
-k6v1 0.0833965 0.0826384 0.326005 0.698256 0.505686 0.802881 0.830174 0.702047
-k6v2 0.369219 0.307809 0.639121 0.839272 0.74981 0.879454 0.884003 0.70887
-k6v4 0.404852 0.333586 0.643669 0.844579 0.755876 0.876422 0.884761 0.718726
-k6v6 0.421531 0.335102 0.6558 0.852919 0.751327 0.881729 0.884003 0.71721
-k6v8 0.41395 0.341168 0.657316 0.84837 0.752843 0.884761 0.884003 0.721001
-k8v1 0.0849128 0.0841547 0.32373 0.704321 0.512509 0.805155 0.829416 0.695224
-k8v2 0.388173 0.313874 0.639879 0.829416 0.739196 0.881729 0.886277 0.71721
-k8v4 0.416983 0.340409 0.652009 0.845337 0.755876 0.874905 0.88097 0.721001
-k8v6 0.419257 0.339651 0.661107 0.84837 0.749052 0.878696 0.886277 0.722517
-k8v8 0.415466 0.33207 0.659591 0.850644 0.743745 0.876422 0.886277 0.723275
-bf16 0.413192 0.327521 0.657316 0.852161 0.741471 0.874147 0.890826 0.717968
+[…]
+Qwen3 Llama-3.2 Llama-3.2 Qwen3 Llama-3.1 Qwen3 Qwen3
+0.6B 1B 3B 4B 8B 8B 32B
+$\mathbf{K}_{1}\mathbf{V}_{1}$ 0.011 0.012 0.016 0.010 0.016 0.011 0.020
+$\mathbf{K}_{1}\mathbf{V}_{2}$ 0.008 0.014 0.015 0.010 0.020 0.013 0.018
+$\mathbf{K}_{1}\mathbf{V}_{4}$ 0.009 0.020 0.025 0.014 0.020 0.008 0.011
+$\mathbf{K}_{1}\mathbf{V}_{6}$ 0.010 0.011 0.027 0.017 0.020 0.008 0.016
+$\mathbf{K}_{1}\mathbf{V}_{8}$ 0.008 0.015 0.028 0.009 0.015 0.014 0.011
+$\mathbf{K}_{2}\mathbf{V}_{1}$ 0.003 0.018 0.015 0.008 0.050 0.013 0.014
+$\mathbf{K}_{2}\mathbf{V}_{2}$ 0.003 0.029 0.278 0.027 0.385 0.129 0.385
+$\mathbf{K}_{2}\mathbf{V}_{4}$ 0.003 0.020 0.351 0.043 0.455 0.171 0.434
+$\mathbf{K}_{2}\mathbf{V}_{6}$ 0.004 0.026 0.344 0.040 0.446 0.169 0.419
+$\mathbf{K}_{2}\mathbf{V}_{8}$ 0.004 0.028 0.355 0.035 0.460 0.167 0.449
+$\mathbf{K}_{4}\mathbf{V}_{1}$ 0.023 0.073 0.281 0.577 0.517 0.779 0.721
+$\mathbf{K}_{4}\mathbf{V}_{2}$ 0.040 0.294 0.639 0.808 0.757 0.877 0.649
+$\mathbf{K}_{4}\mathbf{V}_{4}$ 0.040 0.333 0.653 0.830 0.778 0.877 0.629
+$\mathbf{K}_{4}\mathbf{V}_{6}$ 0.055 0.332 0.668 0.832 0.763 0.881 0.619
+$\mathbf{K}_{4}\mathbf{V}_{8}$ 0.042 0.341 0.658 0.826 0.758 0.878 0.632
+$\mathbf{K}_{6}\mathbf{V}_{1}$ 0.093 0.096 0.308 0.632 0.522 0.795 0.701
+$\mathbf{K}_{6}\mathbf{V}_{2}$ 0.368 0.312 0.644 0.830 0.762 0.880 0.598
+$\mathbf{K}_{6}\mathbf{V}_{4}$ 0.405 0.334 0.654 0.837 0.770 0.878 0.610
+$\mathbf{K}_{6}\mathbf{V}_{6}$ 0.420 0.337 0.663 0.844 0.772 0.884 0.596
+$\mathbf{K}_{6}\mathbf{V}_{8}$ 0.414 0.342 0.665 0.841 0.771 0.886 0.608
+$\mathbf{K}_{8}\mathbf{V}_{1}$ 0.088 0.096 0.304 0.637 0.530 0.798 0.708
+$\mathbf{K}_{8}\mathbf{V}_{2}$ 0.390 0.313 0.647 0.826 0.750 0.882 0.606
+$\mathbf{K}_{8}\mathbf{V}_{4}$ 0.416 0.340 0.660 0.832 0.770 0.875 0.621
+$\mathbf{K}_{8}\mathbf{V}_{6}$ 0.419 0.342 0.667 0.843 0.767 0.880 0.608
+$\mathbf{K}_{8}\mathbf{V}_{8}$ 0.413 0.334 0.666 0.843 0.759 0.880 0.603
+BF16 0.412 0.328 0.663 0.845 0.760 0.877 0.603
+[…]
+Qwen3 Llama-3.2 Llama-3.2 Qwen3 Llama-3.1 Qwen3 Qwen3
+0.6B 1B 3B 4B 8B 8B 32B
+$\mathbf{K}_{1}\mathbf{V}_{1}$ 0.000 0.000 0.000 0.000 0.000 0.000 0.000
+$\mathbf{K}_{1}\mathbf{V}_{2}$ 0.000 0.000 0.000 0.000 0.000 0.000 0.000
+$\mathbf{K}_{1}\mathbf{V}_{4}$ 0.000 0.000 0.000 0.000 0.000 0.000 0.000
+$\mathbf{K}_{1}\mathbf{V}_{6}$ 0.000 0.000 0.002 0.000 0.000 0.001 0.000
+$\mathbf{K}_{1}\mathbf{V}_{8}$ 0.000 0.000 0.000 0.000 0.000 0.000 0.000
+$\mathbf{K}_{2}\mathbf{V}_{1}$ 0.000 0.001 0.005 0.000 0.028 0.000 0.002
+$\mathbf{K}_{2}\mathbf{V}_{2}$ 0.000 0.011 0.276 0.004 0.380 0.067 0.234
+$\mathbf{K}_{2}\mathbf{V}_{4}$ 0.000 0.010 0.350 0.009 0.444 0.104 0.334
+$\mathbf{K}_{2}\mathbf{V}_{6}$ 0.000 0.012 0.344 0.010 0.439 0.115 0.330
+$\mathbf{K}_{2}\mathbf{V}_{8}$ 0.000 0.014 0.353 0.005 0.449 0.109 0.347
+$\mathbf{K}_{4}\mathbf{V}_{1}$ 0.006 0.061 0.293 0.647 0.500 0.785 0.701
+$\mathbf{K}_{4}\mathbf{V}_{2}$ 0.042 0.297 0.628 0.825 0.751 0.876 0.726
+$\mathbf{K}_{4}\mathbf{V}_{4}$ 0.044 0.331 0.643 0.844 0.759 0.876 0.734
+$\mathbf{K}_{4}\mathbf{V}_{6}$ 0.055 0.331 0.657 0.839 0.748 0.880 0.733
+$\mathbf{K}_{4}\mathbf{V}_{8}$ 0.049 0.342 0.649 0.835 0.745 0.877 0.737
+$\mathbf{K}_{6}\mathbf{V}_{1}$ 0.083 0.083 0.326 0.698 0.506 0.803 0.702
+$\mathbf{K}_{6}\mathbf{V}_{2}$ 0.369 0.308 0.639 0.839 0.750 0.879 0.709
+$\mathbf{K}_{6}\mathbf{V}_{4}$ 0.405 0.334 0.644 0.845 0.756 0.876 0.719
+$\mathbf{K}_{6}\mathbf{V}_{6}$ 0.422 0.335 0.656 0.853 0.751 0.882 0.717
+$\mathbf{K}_{6}\mathbf{V}_{8}$ 0.414 0.341 0.657 0.848 0.753 0.885 0.721
+$\mathbf{K}_{8}\mathbf{V}_{1}$ 0.085 0.084 0.324 0.704 0.513 0.805 0.695
+$\mathbf{K}_{8}\mathbf{V}_{2}$ 0.388 0.314 0.640 0.829 0.739 0.882 0.717
+$\mathbf{K}_{8}\mathbf{V}_{4}$ 0.417 0.340 0.652 0.845 0.756 0.875 0.721
+$\mathbf{K}_{8}\mathbf{V}_{6}$ 0.419 0.340 0.661 0.848 0.749 0.879 0.723
+$\mathbf{K}_{8}\mathbf{V}_{8}$ 0.415 0.332 0.660 0.851 0.744 0.876 0.723
+BF16 0.413 0.328 0.657 0.852 0.741 0.874 0.718
+C.6 Rotation and Grouping Results
+Tables 11-14 present the full set of downstream evaluation results for integrating rotation-based outlier redistribution with mixed-precision KV quantization across multiple models and tasks, including GSM8K (flexible and strict exact match), CoQA (F1), and CoQA (exact match).
+Each table reports accuracy for four key-value bit allocations ($\mathbf{K}_{2}\mathbf{V}_{2}$, $\mathbf{K}_{2}\mathbf{V}_{4}$, $\mathbf{K}_{4}\mathbf{V}_{2}$, $\mathbf{K}_{4}\mathbf{V}_{4}$) under four Rotations: none, key-only, value-only, and both key and value.
+The group size is fixed at 64 for both keys and values in these experiments to isolate the effect of rotation.
+The results reveal several consistent trends across model scales and tasks.
+First, applying rotation to keys consistently improves performance at lower bit-widths ($\mathrm{K}_{2}\mathrm{V}_{2}$ and $\mathrm{K}_{2}\mathrm{V}_{4}$), significantly narrowing the gap to higher-precision baselines.
+In contrast, rotating values alone yields marginal or even negative effects, especially at low precisions.
+Notably, applying rotation to keys only under the $\mathrm{K}_{4}\mathrm{V}_{2}$ configuration achieves accuracy that closely matches the full $\mathrm{K}_{4}\mathrm{V}_{4}$ baseline, indicating that the dominant benefits of rotation stem from mitigating key outliers rather than value outliers.
+Applying rotation to both keys and values provides little additional gain beyond key-only rotation, reinforcing that keys are the primary target for rotation-based improvements.
+Figure 11 examines the effect of group size configuration under a fixed $\mathrm{K}_{4}\mathrm{V}_{2}$ mixed-precision setting without rotation.
+Group size ($\mathrm{gs}$) determines the block granularity of quantization: smaller groups offer finer scaling at the cost of increased metadata and computation.
+The results reveal a clear asymmetry between keys and values.
+Reducing group size for keys consistently improves downstream accuracy across CoQA, GSM8K, EQ-Bench, and LongBench, with $\mathrm{gs}_{K}=32$ achieving the best overall performance.
+This reflects the higher sensitivity of key caches to quantization distortion.
+In contrast, increasing value group size to 64 or 128 has a negligible impact on accuracy while reducing overhead, consistent with their lower sensitivity.
+Together, these findings provide practical guidance for combining geometry-driven bit allocation with rotation and grouping strategies: rotation should primarily target keys, while keys benefit from finer group granularity and higher precision; values, on the other hand, can use coarser grouping and lower precision with minimal accuracy loss.
+Table 11: Rotation-based Outlier Redistribution Integration on GSM8K - Exact Match (Flexible).
+Downstream accuracy under different key-value bit allocations and rotation scopes.
+Group size is fixed to 64 for both keys and values.
+Model Rotation $\mathbf{K}_{2}\mathbf{V}_{2}$ $\mathbf{K}_{2}\mathbf{V}_{4}$ $\mathbf{K}_{4}\mathbf{V}_{2}$ $\mathbf{K}_{4}\mathbf{V}_{4}$
+Llama 3.1-8B K+V 52.99 55.80 76.50 76.95
+K only 52.92 57.32 75.44 76.50
+V only 20.70 25.70 75.28 76.95
+none 18.65 24.41 75.21 76.95
+Llama 3.2-1B K+V 1.97 2.43 30.93 33.13
+K only 2.05 3.03 28.43 33.13
+V only 2.05 1.29 28.58 29.57
+none 1.82 2.20 27.22 29.57
+Llama 3.2-3B K+V 40.11 44.12 62.85 64.37
+K only 38.06 45.26 61.18 63.91
+V only 16.60 19.26 63.46 63.91
+none 14.86 21.08 62.93 63.53
+Qwen 0.6B K+V 0.76 0.61 37.83 36.47
+K only 1.52 1.06 36.92 37.68
+V only 1.52 1.67 1.36 2.12
+none 0.83 1.90 1.36 2.05
+Qwen 32B K+V 31.69 33.81 66.49 64.22
+K only 30.10 33.51 63.99 65.28
+V only 6.90 11.07 65.81 63.68
+none 6.67 10.92 64.82 63.76
+Qwen 4B K+V 0.91 0.76 83.78 85.37
+K only 1.14 0.68 82.26 84.76
+V only 0.83 1.14 82.56 83.70
+none 1.14 0.83 81.80 83.40
+Qwen 8B K+V 29.34 36.01 88.17 88.17
+K only 29.57 35.56 87.64 87.95
+V only 1.97 1.21 86.73 87.34
+none 2.27 1.67 86.58 88.02
+Table 12: Rotation-based Outlier Redistribution Integration on GSM8K - Exact Match (Strict).
+Downstream accuracy under different key-value bit allocations and rotation scopes.
+Group size is fixed to 64 for both keys and values.
+Model Rotation $\mathbf{K}_{2}\mathbf{V}_{2}$ $\mathbf{K}_{2}\mathbf{V}_{4}$ $\mathbf{K}_{4}\mathbf{V}_{2}$ $\mathbf{K}_{4}\mathbf{V}_{4}$
+Llama 3.1-8B K+V 51.71 55.27 74.37 74.45
+K only 51.93 56.48 73.24 73.92
+V only 19.33 24.49 73.01 74.60
+none 17.97 22.74 72.93 74.60
+Llama 3.2-1B K+V 0.83 1.21 31.01 32.98
+K only 0.68 1.29 29.04 33.06
+V only 0.38 0.61 28.51 29.57
+none 0.45 0.61 27.22 29.34
+Llama 3.2-3B K+V 39.65 43.67 61.94 63.84
+K only 36.47 44.96 60.27 63.46
+V only 16.00 19.26 63.00 63.38
+none 14.40 21.00 62.02 62.85
+Qwen 0.6B K+V 0.00 0.00 38.14 36.92
+K only 0.00 0.00 37.38 37.83
+V only 0.00 0.00 0.15 0.53
+none 0.00 0.00 0.30 0.08
+Qwen 32B K+V 21.99 28.35 74.75 75.59
+K only 21.46 26.38 74.75 75.28
+V only 2.20 5.00 74.00 75.13
+none 1.52 4.93 73.77 74.37
+Qwen 4B K+V 0.23 0.08 84.53 85.60
+K only 0.00 0.08 83.40 84.99
+V only 0.00 0.00 82.49 83.70
+none 0.00 0.08 81.80 83.78
+Qwen 8B K+V 25.32 30.33 87.72 87.87
+K only 25.17 32.15 87.49 87.34
+V only 0.15 0.00 86.43 86.88
+none 0.00 0.08 86.50 87.41
+Table 13: Rotation-based Outlier Redistribution Integration on CoQA - F1 Score.
+Downstream accuracy under different key-value bit allocations and rotation scopes.
+Group size is fixed to 64 for both keys and values.
+Model Rotation $\mathbf{K}_{2}\mathbf{V}_{2}$ $\mathbf{K}_{2}\mathbf{V}_{4}$ $\mathbf{K}_{4}\mathbf{V}_{2}$ $\mathbf{K}_{4}\mathbf{V}_{4}$
+Llama 3.1-8B K+V 77.70 77.83 78.88 78.45
+K only 77.39 78.10 77.93 78.75
+V only 75.46 76.71 79.34 78.76
+none 74.72 76.57 78.10 79.18
+Llama 3.2-1B K+V 50.66 53.55 70.44 70.47
+K only 50.18 53.46 69.40 70.27
+V only 39.50 44.93 69.58 70.52
+none 38.82 44.33 69.53 70.31
+Llama 3.2-3B K+V 78.45 78.78 79.24 79.13
+K only 78.24 78.88 78.71 79.20
+V only 73.25 74.48 78.86 79.27
+none 74.10 74.55 79.14 79.37
+Qwen 0.6B K+V 33.99 34.06 68.92 70.09
+K only 33.88 33.70 69.62 70.08
+V only 29.00 27.39 48.19 48.06
+none 28.89 28.04 46.49 48.41
+Qwen 32B K+V 64.85 67.39 82.46 82.59
+K only 67.16 67.18 82.62 82.55
+V only 77.82 78.55 82.35 82.50
+none 76.80 78.35 82.60 82.77
+Qwen 4B K+V 62.43 63.82 80.16 80.56
+K only 62.09 64.35 80.46 80.63
+V only 56.73 59.26 80.07 80.16
+none 54.93 59.68 80.41 80.19
+Qwen 8B K+V 79.33 79.65 82.32 81.98
+K only 78.27 79.89 81.15 82.09
+V only 67.51 70.41 82.23 82.01
+none 66.86 69.83 81.49 81.99
+Table 14: Rotation-based Outlier Redistribution Integration on CoQA - Exact Match.
+Downstream accuracy under different key-value bit allocations and rotation scopes.
+Group size is fixed to 64 for both keys and values.
+Model Rotation $\mathbf{K}_{2}\mathbf{V}_{2}$ $\mathbf{K}_{2}\mathbf{V}_{4}$ $\mathbf{K}_{4}\mathbf{V}_{2}$ $\mathbf{K}_{4}\mathbf{V}_{4}$
+Llama 3.1-8B K+V 62.43 63.37 63.25 63.40
+K only 62.12 63.32 63.03 63.65
+V only 58.13 60.12 64.48 63.75
+none 58.12 59.65 62.92 64.45
+Llama 3.2-1B K+V 33.38 37.87 54.78 56.02
+K only 33.72 37.30 54.22 55.32
+V only 25.47 30.37 53.97 55.65
+none 24.42 30.23 54.30 55.45
+Llama 3.2-3B K+V 62.47 62.28 62.83 63.40
+K only 62.12 62.72 62.93 63.53
+V only 56.07 57.48 62.70 63.73
+none 56.45 57.62 63.42 63.77
+Qwen 0.6B K+V 19.67 19.57 56.07 56.73
+K only 18.72 18.18 55.53 56.73
+V only 13.97 14.27 28.18 28.17
+none 14.25 13.80 28.22 28.82
+Qwen 32B K+V 46.97 49.95 70.18 70.77
+K only 50.28 49.42 70.48 70.87
+V only 61.77 64.08 70.02 70.37
+none 61.12 63.92 69.90 70.50
+Qwen 4B K+V 41.28 42.45 65.68 66.28
+K only 41.65 43.45 66.40 66.47
+V only 39.22 40.03 65.43 66.20
+none 37.70 40.22 65.52 66.07
+Qwen 8B K+V 61.70 61.88 68.55 67.73
+K only 60.07 62.40 66.62 68.23
+V only 49.62 51.58 68.10 68.07
+none 48.40 51.65 67.07 68.00
+Figure 11: Effect of group size configurations on downstream accuracy.
+Each bar shows performance on CoQA, GSM8K, EQ-Bench, and LongBench for different key-value grouping combinations ($\mathrm{gs}_{K},\mathrm{gs}_{V}\in\{32,64,128\}$) under the $\mathbf{K}_{4}\mathbf{V}_{2}$ mixed-precision setting.
+Smaller group sizes for keys consistently improve accuracy across all tasks, while value group size has a smaller but non-negligible effect.
+Configurations with $\mathrm{gs}_{K}=32$ achieve the best overall performance, highlighting the benefit of finer quantization granularity for key caches, which are more sensitive to quantization distortion.
+In contrast, larger group sizes for values ($\mathrm{gs}_{V}=64$ or 128128) preserve performance while reducing overhead, aligning with their lower sensitivity.
+Table 15: Performance by rotation strategy across tasks.
+Values report the mean performance across seven models (0.6B–32B parameters) and four quantization configurations ($\mathrm{K}_{2}\mathrm{V}_{2}$, $\mathrm{K}_{2}\mathrm{V}_{4}$, $\mathrm{K}_{4}\mathrm{V}_{2}$, $\mathrm{K}_{4}\mathrm{V}_{4}$).
+The $\pm$ values indicate the variability range across model scales and quantization settings, with larger ranges reflecting higher sensitivity to these factors.
+Task Metric No Rotation Value-Only Key-Only Both
+CoQa Exact Match $50.10\pm 17.65$ $1.82\pm 17.22 55$$34\pm 14.21 55.$$5.39\pm 14.10$
+Eq Bench Parseable $6.18\pm 42.57 55$$5.01\pm 41.93$ $65.73\pm 39.70$ $64.81\pm 39.19$
+GSM8K Exact Match $32.73\pm 33.42$ $32.96\pm 33.48$ $43.46\pm 28.98$ $43.80\pm 29.28$
+Longbench ROUGE Score $16.70\pm 11.75$ $16.54\pm 11.82$ $17.50\pm 12.83$ $17.63\pm 12.81$
+Appendix D Reproducibility and Resources
+Inference was performed using the Hugging Face Transformers Wolf et al.
+(2020) and Accelerate Gugger et al.
+(2022) with FlashAttention Dao et al.
+(2022); Dao (2024).
+We integrated both Quanto and HQQ into the Language Model Evaluation Harness Gao et al.
+(2024) to enable systematic and reproducible evaluation of model performance under varying quantization schemes.
+Quantization error is measured by MSE, and confidence intervals use Bayesian variance.
+Hariri et al.
+(2025).
+Evaluations were executed on two High-performance Computing (HPC) clusters, detailed in Table 16.
+Table 16: Specifications of Two High-Performance Computing (HPC) Clusters Used in This Study
+Cluster Cluster A Cluster B
+Processor AMD EPYC 7742 $\times$ 2 Intel Xeon Platinum 8468 $\times$ 2
+RAM 2048 GB 2048 GB
+GPU NVIDIA A100 $\times$ 8 NVIDIA H200 $\times$ 8
+VRAM 80 GB HBM2e 141 GB HBM3e
+Scale 5 nodes (40 A100) 5 nodes (40 H200)